### v1

# SymbolicAI: A framework for logic-based approaches combining generative models and solvers

Abstract
We introduce SymbolicAI, a versatile and modular framework employing a logic-based approach to concept learning and flow management in generative processes. SymbolicAI enables the seamless integration of generative models with a diverse range of solvers by treating large language models (LLMs) as semantic parsers that execute tasks based on both natural and formal language instructions, thus bridging the gap between symbolic reasoning and generative AI. We leverage probabilistic programming principles to tackle complex tasks, and utilize differentiable and classical programming paradigms with their respective strengths. The framework introduces a set of polymorphic, compositional, and self-referential operations for data stream manipulation, aligning LLM outputs with user objectives. As a result, we can transition between the capabilities of various foundation models endowed with zero- and few-shot learning capabilities and specialized, fine-tuned models or solvers proficient in addressing specific problems. In turn, the framework facilitates the creation and evaluation of explainable computational graphs. We conclude by introducing a quality measure and its empirical score for evaluating these computational graphs, and propose a benchmark that compares various state-of-the-art LLMs across a set of complex workflows. We refer to the empirical score as the ”Vector Embedding for Relational Trajectory Evaluation through Cross-similarity”, or VERTEX score for short. The framework codebase111 SymbolicAI framework released on January 20th, 2023, on GitHub: https://github.com/ExtensityAI/symbolicai. and benchmark222 Evaluation benchmark released on February 1st, 2024, on GitHub: https://github.com/ExtensityAI/benchmark. are linked below.
Figure 1: Our neuro-symbolic framework enables a seamless transition between symbolic and differentiable programming, each with distinct dynamics and strengths. Differentiable programming provides access to foundational and specialist models. Classical programming, on the other hand, shifts between abstraction and implementation, focusing on high-level concepts before delving into the details of implementation.
1 Introduction
The recent surge in generative AI, particularly involving large language models (LLMs), has demonstrated their wide-ranging applicability across various domains (Badita, 2022; Degrave, 2022). These models have enhanced the functionality of tools for search-based interactions (Microsoft, 2023; YouWrite, 2022; Writesonic, 2022), program synthesis (Jain et al., 2021; Romera-Paredes et al., 2023; Key et al., 2023), chat-based interactions (ReplikaAI, 2016; OpenAI, 2022; Google, 2023), and many more. Moreover, language-based approaches have facilitated connections between different modalities, enabling text-to-image (Ramesh et al., 2021; Saharia et al., 2022), text-to-video (Singer et al., 2022), text-to-3D (Poole et al., 2022), text-to-audio (Oord et al., 2016; Wang et al., 2017), and text-to-code (Wang et al., 2021b; Lu et al., 2021; Li et al., 2022b) transformations, to name a few. Therefore, by training on vast quantities of unlabelled textual data, LLMs have been shown to not only store factual knowledge (Petroni et al., 2019; Kassner et al., 2020) and approximate users’ intentions to some extent (Andreas, 2022), but also to unlock deep specialist capabilities through innovative prompting techniques (Nori et al., 2023). Yet, these applications merely scratch the surface of the transformation that language-based interactions are expected to bring to human-computer interactions in both the near and distant future.
In part, instruction-based fine-tuning of LLMs through reinforcement learning from human feedback (Ouyang et al., 2022; Li et al., 2023) or direct preference optimization (Rafailov et al., 2023) has shown promising results dealing with value misalignment issues (Bradley Knox & Stone, 2008; MacGlashan et al., 2017; Christiano et al., 2017; Ibarz et al., 2018; Goyal et al., 2022), unlocking new possibilities for chain of thoughts (Wei et al., 2022b), tree of thoughts (Yao et al., 2023a), and graph of thoughts interactions (Besta et al., 2023). However, recent research also highlights the limitations of LLMs in functional linguistic competence despite their proficiency in formal linguistic competence (Mahowald et al., 2023). Whereas formal linguistic competence encompasses the ability to understand and generate language, functional linguistic competence pertains to the application of language in real-world contexts, such as conveying sensory input or recalling information from memory. Examples of functional linguistic competence include implicatures (Ruis et al., 2022) and contextual language comprehension beyond the statistical manifestation of data distributions (Bransford & Johnson, 1972). Consequently, operating LLMs through a purely inference-based approach confines their capabilities within their provided context window, severely limiting their horizon. This results in deficiencies for situational modeling, non-adaptability through contextual changes, and short-term problem-solving, amongst other capabilities. However, simply increasing the context length may not yield greater capabilities, as demonstrated by the observed U-shaped performance curve (Liu et al., 2023) where LLMs excel when using information at the beginning or end of the input context, but struggle with information located in the middle, especially as context increases. These challenges are actively being researched, with novel approaches such as Hyena (Poli et al., 2023), RWKV (Bo, 2021), GateLoop (Katsch, 2023), and Mamba (Gu & Dao, 2023) surfacing. Meanwhile, the re-emergence of interest in retrieval-augmented generative approaches (Li et al., 2022a) offers an alternative by circumventing the autoregressive nature of the widely-used Transformer architecture (Vaswani et al., 2017), enabling context enrichment with lateral information. In parallel, efforts have focused on developing tool-based approaches (Schick et al., 2023) or template frameworks (Chase, 2023) to extend large LLMs’ capabilities and enable a broader spectrum of applications. However, these efforts only partially capture the vast potential inherent in leveraging LLMs as semantic parsers.
In light of these considerations, we introduce SymbolicAI, a compositional neuro-symbolic (NeSy) framework able to represent and manipulate compositional, multi-modal, and self-referential structures (Schmidhuber, 2007; Fernando et al., 2023). SymbolicAI augments the generative process of LLMs with functional zero- and few-shot learning operations and enables the creation of versatile applications through in-context learning (Wei et al., 2022a). These operations guide the generative process and facilitate a modular design with a wide range of existing solvers, including formal language engines for mathematical expression evaluation, theorem provers, knowledge bases, and search engines for information retrieval. It exposes these solvers as building blocks for constructing compositional functions as computational graphs, and facilitates the development of an extensible toolkit that bridges classical and differentiable programming paradigms, aiming to create domain-invariant problem solvers. In designing the architecture of SymbolicAI, we drew inspiration from the body of evidence that suggests the human brain possesses a selective language processing module (Deniz et al., 2019; Hu et al., 2022; Menenti et al., 2011; Scott et al., 2016; Macsweeney, 2002; Fedorenko et al., 2010; Regev et al., 2013), prior research on cognitive architectures (Newell & Simon, 1956; Newell et al., 1957; Newell & Simon, 1972; Newell, 1990; Laird, 2022), and the significance of language on the structure of semantic maps in the human brain (Huth et al., 2016). We consider language as a central processing module, distinct from other cognitive processes such as reasoning or memory (Paischer et al., 2022, 2023), that defines a stepping stone towards broad AI systems (see Section B).
Lastly, alongside the framework, we introduce a benchmark and derive a quality measure and its empirical score to address the evaluation of multi-step NeSy generative processes. See Section 6 for more details.
In summary, we highlight the following key contributions of this work:
  • We introduce SymbolicAI, a logic-based framework for concept learning and flow management in generative processes, enabling seamless integration with a wide range of foundation models and solvers.
  • We combine the broad applicability of LLMs as semantic parsers with symbolic expressions by leveraging a modular probabilistic programming paradigm, facilitating the creation of complex computational graphs.
  • We introduce a quality measure and its empirical score alongside a benchmark for comparing state-of-the-art LLMs across a wide range of tasks.
2 Related Work
Symbolic Methods
The field of symbolic AI has its foundations in the works of the Logic Theorist (LT) (Newell & Simon, 1956) and the General Problem Solver (GPS) (Newell et al., 1957). These programs represented the first steps towards automated reasoning and problem-solving using symbolic representations. Despite their advancements, both faced challenges in dealing with the complexity of real-world problems, particularly due to the combinatorial nature of the solution space. To address these limitations, the Soar (Laird et al., 1987) cognitive architecture was developed, advancing the notion that intelligent behavior results from goal-oriented search through a problem space (Newell & Simon, 1972; McCarthy et al., 2006), with each step consisting of selecting and applying operators. Soar introduced components like reinforcement learning, impasses, substates, and chunking to enhance its problem-solving capabilities. It also demonstrated the importance of learning from experiences to adapt and improve performance over time. However, Santoro et al. (2022) emphasizes the subjectivity of symbols and suggests that human-like symbolic fluency could develop in machines through learning algorithms immersed in socio-cultural contexts. This perspective, anchored in the notion that symbols are triadic and their meaning emerges from consensus, seeks to move away from traditional symbolic AI methodologies towards AI that adaptively learns meaning and behaviors from human-like experiences. The goal is to cultivate machines that demonstrate symbolic behaviors across a spectrum of competencies, potentially mirroring the evolutionary and social learning processes observed in humans. Lastly, symbolic AI struggles with real-world data’s unpredictability and variability. These challenges have led to the employment of statistical learning methodologies, like deep learning (Alom et al., 2018), which are more adept at managing noise and uncertain information through vector-valued representations.
Sub-Symbolic Methods
The sub-symbolic framework, rooted in neural network paradigms, began with pioneering works such as the perceptron (McCulloch & Pitts, 1943), with the first hardware implementation quickly following (Rosenblatt, 1958). The foundational notion of distributed processing (Rumelhart et al., 1986) was later bolstered and further expanded by demonstrating that multilayer feedforward networks with a single hidden layer can serve as universal approximators for any Borel measurable function, given sufficient hidden units (Hornik et al., 1989). Fast-forward, contemporary frameworks achieve a significant leap with the introduction of the Transformer architecture (Vaswani et al., 2017), which underpins most of today’s LLMs. These LLMs demonstrate exceptional capabilities in in-context learning, a method popularized by the likes of GPT-3 (Brown et al., 2020), where models improve task performance through natural language instruction and examples provided directly in the input prompt. While in-context learning bypasses the need for explicit retraining, it demands meticulous prompt design to steer models towards desired behaviors. Despite their versatility, current LLMs face challenges such as fallacious reasoning and the generation of erroneous content, commonly referred to as hallucinations (Jones & Steinhardt, 2022). These limitations highlight the importance of integrating complementary symbolic methods to validate and guide the generative processes of LLMs, ensuring more accurate and reliable outputs.
Neuro-Symbolic Methods
To overcome the limitations of each individual method, NeSy approaches meld the statistical inference strengths of deep neural architectures with the generalization and explainability of symbolic systems (Besold et al., 2017; Yu et al., 2023; Hamilton et al., 2022; Garcez et al., 2015; d’Avila Garcez et al., 2019; d’Avila Garcez & Lamb, 2020; Lamb et al., 2020). Some approaches focus on different strategies for integrating learning and reasoning processes (Yu et al., 2023; Fang et al., 2024). Firstly, learning for reasoning methods treat the learning aspect as an accelerator for reasoning, in which deep neural networks are employed to reduce the search space for symbolic systems (Qu & Tang, 2019; Silver et al., 2016, 2017b, 2017a; Schrittwieser et al., 2020). Secondly, reasoning for learning views reasoning as a way to regularize learning, in which symbolic knowledge acts as a guiding constraint that oversees machine learning tasks (Hu et al., 2016; Xu et al., 2018). Thirdly, the learning-reasoning category enables a symbiotic relationship between learning and reasoning. Here, both elements interact and share information to boost problem-solving capabilities (Donadello et al., 2017; Manhaeve et al., 2018; Mao et al., 2019; Ellis, 2023). This synergy further extends when considering graph-based methods, which closely align with the objectives of our proposed framework. Research in this area, such as CycleGT (Guo et al., 2020) and Paper2vec (Ganguly & Pudi, 2017), explored unsupervised techniques for bridging graph and text representations. Subsequently, graph embeddings, when utilized within symbolic frameworks, can enhance knowledge graph reasoning tasks (Zhang et al., 2021), or more generally, provide the bedrock for learning domain-invariant representations (Park et al., 2023). Lastly, building upon the insights from Sun et al. (2022), the integration of NeSy techniques in scientific workflows promises significant acceleration in scientific discovery. While previous work has effectively identified opportunities and challenges, we have taken a more ambitious approach by developing a comprehensive framework from the ground up to facilitate a wide range of NeSy integrations.
3 Problem Definition
Conventional approaches employing foundation models for inference, such as LLMs, are predominantly confined to single-step or few-step executions and primarily reliant on hand-crafted in-context learning prompt instructions. This restricted scope limits the utilization to single modalities, lacks refinement or verification, and exhibits limited tool proficiency. We posit that the integration of NeSy engines as core computation units, realized through logic-based methodologies coupled with sub-symbolic foundation models, offers a more general, robust, and verifiable perspective. This approach has several advantages. Firstly, it facilitates the integration of pre-existing engineered solutions (e.g. various classical algorithms), offloading computational complexity and bridging various modalities. Secondly, it enables sub-symbolic generalization to focus on evidence-based decision-making. Thirdly, it provides an interpretable language-based control layer for explainable, autonomous systems. Central to our solution is a method to define and measure the orchestration of interactions between symbolic and sub-symbolic systems, and the level at which instructions are formulated for effective control and task execution.
4 Design Principles
In the following subsections, we elaborate on the key design principles underlying SymbolicAI and how we guide the generative processes of NeSy engines.
Symbols and Expressions
As posited by Newell & Simon (1976), symbols are elemental carriers of meaning within a computational context333 We base our framework’s name on the aspirational work of Newell and Simon.. These symbols define physical patterns capable of composing complex structures, and are central to the design and interpretation of logic and knowledge representations (Augusto, 2022). Thus, SymbolicAI conceptualizes the notion that symbols, and the expressions they form, are reflections of the information inherent in a system, and serve as surrogate for the interaction between the system and the problem space. Moreover, we argue that real patterns, as Dennett (1991) speaks of, can be effectively realized through the use of symbols because these symbols act as versatile abstractions that capture and represent the underlying structures and dynamics of these patterns, facilitating their interpretation and manipulation in computational models.
Furthermore, we attribute task-specific mappings to a language-centric strategy, leveraging their inherent semantics and abstraction to describe the states and properties of the problem at hand. These mappings are universal and may be used to define scene descriptions, long-horizon planning, acoustic properties, emotional states, physical conditions, etc. Therefore, we adhere to the analogy of language representing the convex hull of the knowledge of our society, utilizing it as a fundamental tool to define symbols. This approach allows us to map the complexities of the world onto language, where language itself serves as a comprehensive, yet abstract, framework encapsulating the diversity of these symbols and their meanings. This perspective resonates with our innate human tendencies to attribute existing physical objects with abstract concepts, as exemplified by our natural inclination to link tangible objects to colors and emotions, such as blending the color ”red” with ”heart”, ”warm”, and ”passion”.
However, this language-centric model does not inherently encompass all forms of representation, such as sensory inputs and non-discrete elements, requiring the establishment of additional mappings to fully capture the breadth of the world. This limitation is manageable, since we care to engage in operations within this abstract conceptual space, and then define corresponding mappings back to the original problem space. These are typically applied through function approximation, as in typical modality-to-language and language-to-modality use cases, where modality is a placeholder for various skill sets such as text, image, video, audio, motion, etc.
Ultimately, this approach also anchors our work in the field of formal language theory, as we require a structured method to construct mappings from the world to language. This grounding sets the foundation for employing formal language structures, such as grammars, to systematically define our language-centric approach to problem-solving and the associated translation of real-world complexities into linguistic terms.
Formal Languages
In formal language theory and linguistics, languages are structured following the Chomsky hierarchy, which classifies languages by the complexity of their grammatical structure (Chomsky, 1956). This hierarchy, comprising four types of grammars (Type-3 to Type-0), delineates formal languages by their grammatical complexity. A grammar in this context consists of terminal and non-terminal symbols, production rules, and a designated start symbol, facilitating the generation of valid strings within a language. In developing SymbolicAI, we posit that all symbols can be represented as strings, augmented with conditional instructions and types derived from a domain-specific language (DSL) tailored for directing NeSy computation engines, like LLMs (see Figure 2).
A key advancement of LLMs over previous systems lies in their ability to generalize from formal languages (Wang et al., 2023a) and knowledge systems, primarily due to their world-knowledge and proficiency in understanding context-based analogies. While there is currently no universal consensus among experts regarding the precise classification of natural language within the Chomsky hierarchy, we have seen in our empirical evaluations promising results utilizing LLMs as semantic parsers. This approach can be viewed as employing a form of flexible, context-sensitive grammar, which enables the processing of instructions and analogies with a nuanced understanding of language’s inherent variability and complexity. The intersection between formal and natural languages becomes evident when considering how language patterns, through prompts like ”You are a helpful assistant…”, elicit structured responses, indicating a potential underlying formal mechanism at play. This observation underlines the utility of such a grammar in our framework, particularly within in-context learning, where it serves as an explicit schema guiding the structure of examples used in few-shot learning scenarios. For instance, equating ”3.1415…” with ”$\pi$” or ”August 4, 1961” with ”1961-08-04” in a given context demonstrates this probabilistic, context-dependent interpretation of symbols. Such a system doesn’t rigidly adhere to standard grammatical rules but instead adjusts and interprets based on the presented context, effectively creating a dynamic and situation-specific grammar.
In-Context Learning
Recently, several in-context learning methodologies evolved to enable tool usage through LLMs (Schick et al., 2023), or refine the generative outcome of LLMs (Yang et al., 2023). This includes chain-of-thought (CoT) prompting, a method that conditions the model to reveal its step-by-step reasoning process (Singhal et al., 2023; Wei et al., 2022b). CoT prompting breaks down complex tasks into simpler, sequential steps, and helps with interpreting LLM’s output. Self-generated CoT, where models are encouraged to generate their own reasoning chains based on training examples, surpasses even expertly crafted CoT (Fernando et al., 2023). This observation echoes other reports that GPT-4 has an emergent self-improving capability through introspection, such as self-verification (Weng et al., 2023) or self-consistency (Wang et al., 2023b). Tree of Thoughts (ToT) enables LLMs to solve complex problems by exploring multiple reasoning paths through a search tree of coherent text units, demonstrating significant problem-solving enhancements in tasks requiring strategic planning and search (Yao et al., 2023a). Ensemble techniques further enhance the robustness and accuracy of model predictions by combining several strategies to establish a consensus (Nori et al., 2023). Conceptually, we designed our framework to enable all these techniques and combine them into dedicated components and sub-processes. Figure 2 outlines a typical expression evaluation in our NeSy pipeline.
Figure 2: Illustration for NeSy pipeline, showcasing conceptual usage of in-context learning methodologies, domain-specific language (DSL) structures, and the expression evaluations through NeSy engines. The expression showcases the left shift operator $\ll$ and how the information of the symbol B is included in the symbol AC. The violet placeholder in the DSL Prompt represents an instruction, such as ”Insert the right-hand side value into the left-hand value in a chronological order.” The positions below that represent task-specific few-shot examples.
Domain-Invariant Associations
In-context learning enabled LLMs to become versatile task solvers by interpolating within the training distribution, to the extent that even potentially unseen tasks are addressable (Brown et al., 2020). We attribute this to associations formed within the input space and the capacity of Transformer architectures for defining domain-invariant feature sub-spaces. This phenomenon has striking parallels with advancements in few-shot learning approaches such as SubGD (Gauch et al., 2022), a method based on identifying and utilizing a low-dimensional subspace, learned from various tasks, that effectively acts to regularize the learning process, representing features that are invariant across different learning tasks. Furthermore, SubGD reflects the potential of in-context learning when combined with task-specific fine-tuning by showing that fine-tuning within a learned subspace significantly outperforms traditional fine-tuning methods. We believe that the extent of in-context learning is not yet exhausted, holding considerable promise when use alongside with task-specific fine-tuning and solvers. To develop learning and reasoning systems capable of general problem-solving, we adopt a hybrid methodology. This approach leverages the in-context generalization capability of LLMs, constructing symbolic associations that aim to preserve and propagate situational context, and validating solutions with established solvers.
Function Composition
In SymbolicAI, function composition is relevant for constructing complex hierarchies and behaviors from more basic, fundamental elements. It enables our framework to model interconnected processes, where the output of one function seamlessly transitions into the input of another, thus creating a cohesive sequence of operations. Through function composition, we construct computational graphs, in which intermediate symbols represent the nodes or states within these graphs. Formally, function composition is denoted by $\circ$, where combining functions $f$ and $g$ yields a new function $h=g\circ f$, defined as $h(x)=g(f(x))$ For functions $f:X\rightarrow Y$ and $g:Y\rightarrow Z$, their composition results in a function mapping elements from domain $X$ to codomain $Z$ through $g(f(x))$.
Although traditionally the codomain of the inner function $f$ aligns with the domain of the outer function $g$, SymbolicAI relaxes this constraint by allowing for any subset relationship between these domains and codomains, enhancing data flow flexibility. For example, this relaxed constraint in domain and codomain alignment is particularly beneficial for in-context learning. By leveraging functional few-shot learning, where few-shot examples act as dynamic elements of the function’s domain, SymbolicAI enhances its ability to interpret and respond to diverse input contexts. For instance, a function can classify a user request and select an appropriate engine to process the request. The target modality may vary based on the respective engine. Therefore, in SymbolicAI, the use of function composition leads to the creation of richer and more nuanced symbol-based generative flows, where each functional unit is capable of representing either logical or data-driven behaviors. Importantly, function composition is not confined to strictly symbolic representations; it also connects with sub-symbolic processes. This enables SymbolicAI to handle complex data streams, establishing function composition as a central tenet in bridging multiple modalities and coordinating a variety of tasks.
5 Framework
In this section, we discuss the design choices pertaining to implementation and the specifics of the framework we’ve employed. For an extended overview, see our supplementary materials Section C.
Types and Representations
Analogous to the Python object type, the base type of SymbolicAI is a symbol represented through its name equivalent base type Symbol. A Symbol object marks a non-reducible atomic unit. All other subtypes, such as Expression and its derivatives, are analogous to their mathematical namesakes, representing expressions or units that can be further evaluated and simplified. These subtypes inherit from Symbol the base attributes, primitive operators, and helper methods. Furthermore, each Symbol object contains valued and vector-valued representations, obtained through value and embedding attributes. The latter, in particular, serve as a means to attribute a symbol’s current context, akin to embedding text and storing it as a PyTorch tensor (Paszke et al., 2019) or NumPy array (Harris et al., 2020). While for an LLM, the numerical tensors may lack inherent meaning, vector-valued representations play a strategic role when 1) composite symbols coalesce into more complex expressions, and 2) these embedded tensors become amenable to updates through gradient-based optimization. This design choice is critical in the dynamic landscape of symbolic interactions, particularly for promising applications, such as developing automated and self-evolving NeSy systems.
To enable the processing of symbols by LLMs, we assume that each Symbol object is represented through Python’s native string functionality, where the __str__ method acts an interpretable string representation. Worth noting is that encoding a complex object into a string sometimes precludes the object reconstitution. However, this concern does not substantially impede our methodology: we can employ approximations or proxy representations stored by the vector-valued property to effectively re-map objects. These representations are obtained through respective embedding models. Therefore, we can theoretically assert that any Python object is parsable by an LLM by design. For more details, see our supplementary materials Section E.
Polymorphic Context
Polymorphism is a central concept in programming language theory and prominently featured in SymbolicAI, with significant implications for the design and functionality of our NeSy architecture. Polymorphism refers to the ability of different objects to be accessed through the same interface, or of a single identifier to represent different types based on the context of execution. The provision of a single interface to entities of different types allows operations to be performed in ways specific to their derived types. We employ a polymorphic structure for instruction composition. We designed the Symbol object to contain a global context, which is composed of static and dynamic context parts. The static context is class dependent and defined at design time. The dynamic context is runtime adaptable and can be changed to adhere to runtime specific logic and changes. Moreover, Symbol associated operations resolve in a polymorphic manner before being evaluated by the NeSy engines. SymbolicAI’s engine implementation contains a prepare method to resolve and compile the engine specific representation by evaluating the Symbol-specific operations and context. For example, when utilizing GPT-4 vision as a NeSy engine backend, we compose the system and user level prompts in the prepare statement and resolve image or video related URL queries by parsing meta tags such as the <<vision: ... :>> tag using regular expressions.
Operators and Methods
In SymbolicAI, operators are overloaded to facilitate transformations of Symbol objects. These operator primitives employ dynamic casting to assure type compatibility, simplifying declarations. Consequently, Symbol objects can be easily manipulated through type specific attributions or symbolically evaluated by the NeSy engine. For example, a central operation for boolean logic is measuring equality between symbols. To evaluate the equality of symbols, we primarily adhere to the type specific implementation, because we prioritize strict comparisons over probabilistic evaluations. If the evaluation was unsuccessful, we then consider semantic equality through the NeSy engine. SymbolicAI leverages decorators for compose operators and custom class methods. For more details, see supplementary materials Section C. Upon invoking an operator or method, the respective primitive function evaluates the symbol’s specific type and its respective attributes, and if further necessary, resolves a nested decorated function that then references the NeSy engine for an evaluation. Should the evaluation fail, a predefined fallback implementation executes. Absent a fallback, or if both evaluations fail, an error state is raised. The processing of an operator or custom method involves a pipeline consisting of pre- and post-processing steps, as well as constraint enforcement. Constraints cover aspects like return types, value ranges, and structural integrity (e.g. JSON formatting through grammar-based verification). In Figure 3 we give an overview of the entire prompt composition based on the user input, the Symbol object structure, and the NeSy engine evaluation pipeline.
Figure 3: The illustrated pipeline utilizes GPT-4 as a NeSy engine and maps all user input and object instances to prompts. Description from left to right; yellow marks user data, blue marks templating and transformations, red marks constraints, and violet the model prediction placeholder. The user arguments and Symbol instance values are evaluated and transformed according to the pre- and post-processing steps. A custom method is constructed by using a reference implementation of decorators and passes its user arguments and instance attributes to the respective engine. A pre-processor manipulates the representation of user arguments and Symbol instance attributes before forwarding them to the engine. For instance, in case the engine is an LLM (i.e. NeSy engine), the static and dynamic context of a Symbol instance are mapped to the system and user prompts for that respective LLM using the prepare method of the engine. Other attributes, such as examples or templates, can help pre-define structures for the desired LLM output format. After the result is computed by the engine, it is passed to the post-processors and validated against the constraints.
Self-Referential Structures
SymbolicAI augments the generative process by enabling systems to introspect and modify their behavior dynamically. We leverage LLMs to execute tasks based on both natural and formal language instructions, adhering to the specified user objectives and with innate self-referential structures. We derive subtypes from Expression and enclose their functionalities in task-specific components, which we then expose again through templating and the model-driven design of the NeSy engine. This design choice allows a system to create and utilize its own sub-process definitions, analogous to concepts discussed in Schmidhuber (2007, 2009). Concretely, we utilize generalization properties from LLMs to interpret and formulate a set of operations that incorporate self-instructions (Wang et al., 2022). Consequently, the operations hold the flexibility to adapt to the context, and derive sub-processes that self-instruct LLMs to engage in situational modeling and context-sensitive problem-solving. Ultimately, this enables the construction of hierarchical computational graphs for self-referential meta-reasoning systems without the need to explicitly training a meta-learner (Kirsch & Schmidhuber, 2022).
6 Evaluation
In this section, we conclude our paper with an empirical evaluation and potential areas of application using state-of-the-art LLMs as NeSy engine backends. For our evaluation we focus on the GPT family of models GPT-3.5 Turbo (revision 1106) and GPT-4 Turbo (revision 1106) (Brown et al., 2020) as they are the most proficient models to this date; Gemini-Pro as the best performing model available through API from Google; LlaMA 2 13B as it defines a good reference implementation for available open-source LLMs from Meta; Mistral 7B and Zephyr 7B as good baselines for revised and fine-tuned open-source contestants respectively. The selected open-source models Mistral, Zephyr, and LlaMA 2 are expected to have roughly equivalent parameter counts compared to GPT-3.5 Turbo and Gemini-Pro. All our experiments are expected to require a context size smaller or equal to $4096$ to enable the comparisons among the in-context capabilities across model architectures. For LlaMA 2 we use the chat version since it better follows instructions. Our evaluation focuses on three main parts. First, we evaluate basic capabilities of the models to verify their consistency and proficiency with in-context learning tasks. Second, we evaluate the logic capability of models by defining logical expressions in mathematical and natural language form and measure the capability of the models to translate and evaluate logical statements across domains. Lastly, we evaluate the capability of models in planning, constructing, maintaining, and executing hierarchical computational graphs. The models are instructed to extract an execution plan, which they then must schedule and execute. A significant challenge is the evaluation of the semantics of a multi-step generative process, since the naïve assessment of task succession would score all models to zero and render them as unusable. Therefore, we derive a quality measure to address these issues.
Performance Measure
One of the challenges in model evaluation is handling irrelevant predictions from models. Even if models follow instructions and produce parts of the expected solution, we regularly observe that they — especially open-source models — append a continuation of task irrelevant predictions. Such predictions result in failure modes when applying conditions and validations, and halt any multi-step procedure. Our solution is an evaluation protocol that refines the performance measurement, allowing for more nuanced diagnostics and the possibility of continuing the evaluation despite intermediate failures. To derive our quality measure, we borrow ideas from the utilization of the Fréchet distance for generative processes (Heusel et al., 2017).
We generate trajectories through a NeSy sequential process that creates a trajectory of distributions $\mathbb{P}$ over multiple iterations of generative nodes. Each node in the process can be aligned to a reference distribution, which marks the desired behavior. To quantify the validity of the generated trajectories, we measure the total distance between the generated and reference data distribution along the path trajectory. We therefore adopt a cumulative measure capable of taking into account the entire generative trajectory. In theory, this process would entail calculating the path integral over the latent space representations for models, cumulating the Fréchet distances (Dowson & Landau, 1982) traversed along these distributional trajectories:
[…]
where $\mathcal{D}(\mathbb{P}_{\text{gen}},\mathbb{P}_{\text{ref}})$ denotes the integral of the Fréchet distances between two data distributions along the generative path trajectory from an initial time $t_{0}$ to a final time $t_{f}$, $d(\mathcal{N}(m_{t},C_{t}),\mathcal{N}(m_{w,t},C_{w,t}))$ is the Fréchet distance calculated at each time $t$ between the generated multivariate normal data distribution with mean $m_{t}$ and covariance $C_{t}$, and the reference multivariate normal data distribution with mean $m_{w,t}$ and covariance $C_{w,t}$. The resulting measure follows properties of normal distributions and is consistent with increasing disturbances.
However, this approach is computationally intractable for large-scale problems, and requires access to latent representations, which — especially in the context of LLMs — is not always given. For computational feasibility, we therefore introduce an approximation that measures the embedding distances over the path trajectories utilizing an auxiliary embedding model, based on prior work on distribution regression (Szabó et al., 2016). The embedding model maps the symbolic representations into a RKHS, such that we can apply a kernel mean embedding function to measure their respective distances (You et al., 2019; Dinu et al., 2023). We assess the distance through the mean embeddings w.r.t. to a kernel function $K(\cdot,\cdot)$ of the samples $\mathbf{e}^{t}_{x}\sim\nu^{t}_{\text{gen}}\in\mathbb{P}_{\text{gen}}$ and $\mathbf{e}^{t}_{y}\sim\nu^{t}_{\text{ref}}\in\mathbb{P}_{\text{ref}}$ produced by the generated data distribution and a reference data distribution respectively. We denote by $\mu_{\mathbf{e}^{t}_{x}}$, $\mu_{\mathbf{e}^{t}_{y}}$ the mean embeddings associated to the respective samples, i.e. $\mu_{\mathbf{e}^{t}_{x}}(z)=\frac{1}{n}\sum_{i=1}^{n}K(x_{i}^{t},z)$ in case $\mathbf{e}^{t}_{x}=(x_{i}^{t})_{i=1}^{n}$ is a sample of size $n$ of the respective mean embeddings. To compute the similarity between the embeddings of the generated and reference distributions, we evaluate the associated maximum mean discrepancy $\text{MMD}^{2}(\mu_{\mathbf{e}^{t}_{x}},\mu_{\mathbf{e}^{t}_{y}})$ (Gretton et al., 2012) and then, as before for the Fréchet distances, we integrate over $t$:
[…]
In empirical evaluations, however, we care about normalized values for ease of interpretation. We therefore analyze the properties of the MMD and derive a similarity score, which follows the same statistical principles as the MMD, and is bound between $[0,1]$. We concluded that we can use only the MMD cross terms to evaluate the similarities. See our supplementary materials Section A for more details. For our comparisons as referenced in Figure 5 we therefore denote the similarities rather than distances. We then come to the following formulation and refer to our empirical measure as the ”Vector Embedding for Relational Trajectory Evaluation through Cross-similarity”, or VERTEX score for short:
[…]
We approximate the integral across time steps through Monte Carlo approximation. The introduced normalization constants denote the similarities to a random sequence $z_{\text{rand}}$, which functions as a baseline subtraction to recenter our results, and a given reference score to rescale w.r.t. to scores obtained from comparing related solutions $z$. The use of the min-max scaling ensures the final measure is bounded between $[0,1]$. This process reflects properties such as Hölder continuity that bounds the kernel function within certain limits, ensuring that the similarity measure is appropriately scaled. To compute the embeddings, we utilize the embedding model all-mpnet-base-v2 (Song et al., 2020), due to its widespread availability, and its balance between speed and quality. As a similarity measure, we select a Gaussian kernel following our derivation from the supplementary materials Section A. In our supplementary implementations, we also explore other kernels, including some preliminary experiments with cosine similarity. We also note that one can integrate Bernoulli distributed trials into our score, with 00 values representing failure modes and values of $1$ being successes. Furthermore, if we relax our definition, we can integrate other similarity measures which are bound between $[0,1]$, which then reflect on domain-specific attributions, i.e. including a similarity measure tailored towards capturing the nuances between two sub-structures of abstract syntax tree.
In our evaluation of Figure 5 we conclude with the cumulative score, for the following base performance criteria.
Basic Capabilities
We analyze the consistency and stability of models and establish baselines by addressing the following performance criteria:
  • Associative Prediction: We evaluate the success rate of models to follow simple and complex instructions and associations with zero- and few-shot examples. We therefore address the proficient use of our operators between Symbol types.
  • Multi-modal Binding: We perform data transformations between multiple modalities by binding through language-based representations, and evaluate their proficiency in tool utilization, classification and routing of requests to relevant modules.
  • Program Synthesis: We evaluate executable code with and without including concepts from retrieval augmented generation, model-driven development, such as templating to direct the generative flow, and experiment with self-generated instructions by creating self-referential expressions. We not only address similarity between produced solutions, but also include the measurement of abstract syntax tree equivalence and the success of executions.
Logical Components
To evaluation the logic capabilities of models, we conditioned them to create a sequence of expressions as self-contained components, and refer to higher-order predicate logic for their assessment. Based on the underlying type theory originating from Whitehead & Russell (1925–1927), we evaluate a model’s capability in the form of there exists x such that x fulfills y. Such quantifiers define the standard semantics over expressions, where the meaning is given by a semantic function. A semantic function maps a term from an abstract definition to a point in a domain, which is an interpretation of the term’s type and value. Therefore, these functions operate on types and values of expressions, and relations thereof. Subsequently, NeSy engines can formulate and evaluate at inference time logic-based instructions through Lisp, Prolog, or Mathematica (McCarthy, 1959; Colmerauer & Roussel, 1993; Chen et al., 1993; Inc., 2022), or leverage SAT solvers such as Z3 (Moura & Bjørner, 2008). Therefore, the evaluation of a natural language statement can be interpreted by any expert system which defines the corresponding semantic functions and process them either in a symbolic (Feigenbaum et al., 1965; Gamble et al., 1994), differentiable (Veličković & Blundell, 2021; Ibarz et al., 2022), or hybrid manner (Kuncicky et al., 1991).
We evaluate how proficient models interpret custom DSLs and define expression statements. The DSL is designed to express logical relations and operations in a structured format, and supports human-readable and machine-interpretable formulations. The following example illustrates such relationships by translating a natural language statement into an expression statement, as follows:
Marvins has four paws and likes to meow when I pet its fur. Is Marvins a cat?
A DSL may enforce the usage of $\text{HAS}(\cdot)$, $\text{IS}(\cdot)$, etc. and may condition a LLM to produce the following expressions:
  • $\text{HasFourPaws}(x)$: $x$ has four paws.
[…]
  • $\text{IsCat}(x)$: $x$ is a cat.
These are then used to define the following logical expression:
$\forall x\big{(}\text{HasFourPaws}(x)\land\text{LikesToMeowWhenPetted}(x)\Rightarrow\text{IsCat}(x)\big{)}.$
An automated theorem prover can now evaluate this statement for all $x$ values and assess the validity of the original query. Lastly, our evaluation uses symbolic mathematics to manipulate algebraic expressions. This involves defining symbols and performing operations like factorization, simplification, and algebraic manipulation. The symbols are placeholders for any value, enabling the definition of general expressions without specifying their values upfront.
Hierarchical Computational Graphs
We evaluate the capabilities of models to orchestrate a multi-step generative process and evaluate a set of tasks. Models need to direct sub-processes and associate computational results from and to Symbol nodes, and maintain relationships between these nodes. Given that the field is currently at an early stage in developing even sequential schedulers for LLM-based planning systems, our initial evaluations will be confined to sequential execution only. In Algorithm 1, we detail our evaluation protocol, which is designed not only to analyze and score a series of instructions but also to provide a structured basis for recording these processes. We also note that our evaluation protocol is generally formulated, which allows the application of non-sequential plan schedulers. In Figure 4 we illustrate a step-wise evaluation of the contextual computational graph, in which the NeSy engine is processing conditioned on the current execution context and producing the next symbol prediction.
Figure 4: We showcase a multi-step hierarchical computational graph, with each node in the graph represented by a symbol. The edges are relations between symbols. The left-hand side illustrates how a new node (Symbol 3) is obtained by evaluating an operation with its respective context on a NeSy engine. The right-hand side illustrates the context information window (yellow rectangle) and relationship of the resulting graph with its respective nodes.
Algorithm 1 VERTEX Protocol
1:NeSy engine: $\mathcal{V}:\mathcal{S}\rightarrow\mathcal{S}$, where $\mathcal{S}=\bigcup_{n\geq 0}\mathbb{L}^{n}$ is the set of all possible strings formed by concatenating characters from a character set $\mathbb{L}$, embedding engine $\mathcal{E}:\mathcal{S}\rightarrow\mathcal{H}\subset\mathbb{R}^{d}$, symbols $\{x_{0},x^{*},y^{*}\}\subset\mathcal{S}$, with $x_{0}$ as the initial instruction, $x^{*}$ as the payload resulted from executing $\mathcal{V}$, $y^{*}$ as the reference, and *** acting as a placeholder for $\mathcal{P},\mathcal{T},\mathcal{C}$, capabilities $\mathcal{C}=\{\mathcal{F}_{1},\mathcal{F}_{2},\mathcal{F}_{3},\ldots\}$, where each $\mathcal{F}_{i}$ represents a specific functional role within the system, plan $\mathcal{P}\subset\mathcal{S}$, task $\mathcal{T}\in\mathcal{P}$, memory buffer $\mathcal{M}\subset\mathcal{S}$, a scoring function $\tilde{s}:\mathcal{H}\times\mathcal{H}\rightarrow[0,1]$, a scheduler $\mathcal{Q}$, an aggregator $\mathcal{A}$, and score variables $\{s\}\in[0,1]$.
2:
3:$\mathcal{V},\mathcal{E},\mathcal{Q},\mathcal{C},y^{\mathcal{P}}\leftarrow$ Init($\cdot$) $\triangleright$ Initialize the engines, the scheduler, the capabilities, and expected plan.
4:$\mathcal{M}\leftarrow\emptyset,\mathcal{A}\leftarrow\emptyset$triangleright Initialize the memory buffer and aggregator.
5:$x^{\mathcal{P}}\leftarrow$ GeneratePlan($x_{0},\mathcal{V}$) $\triangleright$ Generate a plan starting from the initial instruction.
6:Evaluate($x^{\mathcal{P}},y^{\mathcal{P}},\mathcal{E},\mathcal{A},\tilde{s}$) $\triangleright$ Embed, score, and aggregate plan similarity.
[…]
8:while $\mathcal{P}\neq\emptyset$ do $\triangleright$ Run until all tasks have been dequed.
9:     $\mathcal{T},y^{\mathcal{C}},y^{\mathcal{T}}\leftarrow$ Schedule($\mathcal{M},\mathcal{Q}$) $\triangleright$ Select the next task according to the scheduler and progression.
10:     $\mathcal{F}_{i}\leftarrow$ Identify($\mathcal{T},\mathcal{C},\mathcal{V}$) $\triangleright$ Identify the task-related capability $\mathcal{F}_{i}$ using $\mathcal{V}$.
11:     $x^{\mathcal{C}},x^{\mathcal{T}}\leftarrow$ Execute($\mathcal{T},\mathcal{F}_{i}$) $\triangleright$ Execute $\mathcal{T}$ using capability $\mathcal{F}_{i}$ and assign the results $x^{\mathcal{C}},x^{\mathcal{T}}$.
12:     Evaluate($x^{\mathcal{C}},y^{\mathcal{C}},x^{\mathcal{T}},y^{\mathcal{T}},\mathcal{E},\mathcal{A},\tilde{s}$) $\triangleright$ Embed, score, and aggregate capability similarity.
13:     $\mathcal{P},\mathcal{M}\leftarrow$ Update($\mathcal{T},\mathcal{P},\mathcal{M},\mathcal{A}$) $\triangleright$ Update the plan and task progression.
14:end while
15:$s$leftarrow Finalize($\mathcal{A}$) $\triangleright$ Finalize aggregation of scores.
16:return $s$triangleright Return the aggregated score of the plan execution.
Algorithm 1 This algorithm demonstrates the implementation of our Vector Embedding for Relational Trajectory Evaluation through Cross-similarity (VERTEX) protocol. We start by initializing the NeSy engine $\mathcal{V}$, the embedding engine $\mathcal{E}$, the scheduler $\mathcal{Q}$, and a set of capabilities $\mathcal{C}$. The initial instruction $x_{0}$ is used to generate a plan $x^{\mathcal{P}}$ through $\mathcal{V}$. The plan and its expected outcome $y^{\mathcal{P}}$ are embedded, and their similarity is scored and aggregated. The plan is then unfolded into actionable tasks. Each task $\mathcal{T}$ is selected and executed using the appropriate capability $\mathcal{C}$, resulting in the capability and task results $x^{\mathcal{C}},x^{\mathcal{T}}$, and expected outcomes $y^{\mathcal{C}},y^{\mathcal{T}}$ updated in the memory buffer $\mathcal{M}$. The process continues, with each task’s result being embedded, scored, and aggregated until the plan is complete. The final aggregated score $s$ is returned, reflecting the overall effectiveness of the plan execution.
Benchmarks GPT-4 GPT-3.5 Gemini-Pro LlaMA 2 Mistral Zephyr Random Associations 0.93 0.61 1.00 0.41 0.49 0.28 0.01 Modality 0.83 0.77 0.75 0.47 0.39 0.43 0.07 Code 0.78 0.58 0.53 0.10 0.19 0.13 0.00 Logic 0.41 0.44 0.39 0.44 0.46 0.43 0.00 Graphs 0.36 0.31 0.21 0.05 0.06 0.08 0.00 Total 0.66 0.54 0.58 0.30 0.32 0.27 0.02
Figure 5: We evaluate GPT-3.5, GPT4, Gemini-Pro, LLaMA2-Chat 13B, Mistral 7B and Zephyr 7B on five benchmark categories: 1) Associative Prediction (Association) 2) Multi-modal Binding (Modality) 3) Program Synthesis (Code) 4) Functional Logic Components (Logic) and 5) Hierarchical Computational Graphs (Graphs). We denote the similarity score such that higher is better. The similarity score is measured according to a reference baseline and normalized to exclude noise.
7 Discussion
In this section, we address the limitations of SymbolicAI and the future directions we are focusing on. Some of the limitations stem from the inherent constraints of current technologies and dependencies on third-party systems. Additionally, the nuanced complexities of working with generative models presents further challenges.
7.1 Limitation
Embedding Measure
Our empirical measure is limited by the expressiveness of the embedding model and how well it captures the nuances in similarities between two representations. Furthermore, the obtained similarity scores are highly non-linear and difficult to interpret. For instance, two representations may address the same topic, such as the problem description and its respective solution, however, when measuring their similarity we obtain similarity scores of $\sim 70\%$. We normalize this by subtracting an inherent baseline and randomness effect, however, to ensure a more holistic and robust measurement we would need a significantly larger amount of baselines and experiments. Since we were very limited in the availability of development resources, and some presented models are only addressable through costly API walls. We are actively seeking investors and supporters to scale our solution and offer a more compelling benchmark suite in the future.
Model Capabilities
An obvious limitation revolves around the fixed context window size of the underlying language models. Despite the expansion of the context window in newer models such as GPT-4, the finite context still restricts the amount of data that can be processed in a single pass. All information outside the context needs to be added through information retrieval approaches, which come with their own challenges and limitations (Gao et al., 2023). This leads to side effects, including hallucination, given the model does not contain the necessary information to answer the prompted instruction, which makes it difficult to maintain long-term statefulness for complex reasoning tasks and computational graphs.
Error Handling
The complexity of error handling when evaluating complex expressions through function compositionality, especially between multiple modalities and different solvers, is another notable challenge. While SymbolicAI introduces mechanisms for error analysis and automated correction, these approaches are not infallible. They are often limited by the quality and expressiveness of the models, and the model’s capacity to understand deeply nested logical constructs. We also note that for our evaluations, we disabled any remedy protocol, such as truncating prompts or retry schema.
Generalization
This research is also limited by current LLM’s capacity for reasoning and generalization. Although progress has been made, models are still prone to hallucinations and reasoning errors, especially when dealing with abstract, novel, or highly complex problem statements (Marcus, 2020). Furthermore, our framework’s reliance on the model’s ability to grasp the semantics of operations can be influenced by the training data and the model’s innate biases and conceptual understanding (Mahowald et al., 2023). We also point out that the initial development of SymbolicAI started with the GPT family of models, and we may encounter innate biases in prompt design and expressiveness when using other reference models. However, we also point out that prompt engineering instruction-based statements is not a reliable direction for improvement. We instead advocate for enhancing the resilience of models through fault tolerance, focusing on their ability to better follow semantic instructions, not syntactic idiosyncrasies. Another concern is how to assess the disentanglement of evaluations of models on downstream tasks, to avoid testing on training samples, especially for closed-source solutions like GPT.
Interpretability and Transparency
Finally, the issue of explainability and transparency in AI systems remains challenging. While SymbolicAI makes steps towards making computational processes more explicit and explainable through symbolic manipulations, understanding the internal logic and decision-making of LLMs remains an open problem. This can hinder trust and adoption in sensitive applications where interpretability of predictions is important.
7.2 Future Work
The goal for Algorithm 1 is to be utilized by an advanced learning agent. This agent, employing reinforcement learning methodologies (Ouyang et al., 2022; Li et al., 2023; Rafailov et al., 2023), could leverage our evaluation measure in existing benchmarks (Milani et al., 2020; Swazinna et al., 2022; Schweighofer et al., 2022) as a means to obtain reward signals to addresses a central problem in reinforcement learning, namely credit assignment (Sutton, 1984; Arjona-Medina et al., 2019; Holzleitner et al., 2020; Patil et al., 2020; Widrich et al., 2021; Dinu et al., 2022). Over time, it aims to develop the ability to autonomously generate its own plans, efficiently schedule tasks and subtasks, and carefully select the most suitable tools for each task. Our protocol lays the groundwork for this agent to learn and expand its base set of capabilities (Amaro et al., 2023), moving towards more sophisticated, self-referential orchestration of multi-step tasks. We’ve already noticed that research is shifting towards this type of methodology (Yuan et al., 2024). Furthermore, in Section 6 we’ve only considered a sequential scheduler. However, our objective is to ultimately assess a non-sequential task execution model, allowing for dynamic insertion and out-of-sequence task execution.
7.3 Broader Impact
With LLMs becoming more and more accessible, progress recently made possible by the vast open source contributions from Köpf et al. (2023); Touvron et al. (2023); Taori et al. (2023); Xu et al. (2023); Geng et al. (2023); Biderman et al. (2023), embedded accelerators for LLMs – or more generally NeSY engines – will be ubiquitous in future computation platforms, such as wearables, smartphones, tablets, consoles, or notebooks. Although current execution cycles are slow and error-prone, we expect to see further performance gains through improved operating system level optimizations, dedicated GPU-centric hardware refinement, and improved software interoperability. We believe that modern programming paradigms should natively support probabilistic concepts and provide a boilerplate-free set of features for constructing and evaluating generative computational graphs. This includes but is not limited to compositional, parallelizable, and simulation-based executions with polymorphic and self-referential structures. Current programming languages often have disjointed or makeshift solutions for these concepts in the context of generative processes. We believe integral probabilistic support for these concepts into modern software and hardware will unlock new programming paradigms that can fully take advantage of generative architectures. We hope the community will consider these ideas as essential components of contemporary computing.
We also expect to see significant progress by processing central language concepts through system-on-a-chip (SoC) solutions of pre-trained models, with linear probing layers for hot-swappable weight exchange of task-specific projections and executions. A wide range of functionalities can be then offloaded to probabilistic programming languages to operate on dedicated symbols and streamline the vector-valued mappings between the concept space and underlying problem space, avoiding defining boilerplate code to load and unload network weights.
Furthermore, we believe that many gains in representational stability and consistency may be obtained through multi-modal data training and improved alignment based on operator learning oriented functionalities and workflow-related scoring functionalities, analogous to our introduced quality measure. Gains in representational stability also benefit self-instruction and self-referential sub-process evaluations, which enable the dynamic creation and evaluation of complex hierarchical computational graphs. This will enable online learning models to perform, in real-time, skill acquisition of complex concepts with only one or few examples at inference time. We believe this will enable the creation of autonomously self-evolving cognitive architectures (Langley et al., 2009; Dawid & LeCun, 2023; Sumers et al., 2023). We therefore see an inherent connection to generative design as an analogy for creating coherent and stable ”thought” computational graphs, and believe this paves the path toward broad AI systems (see Section B) and is a requirement for developing artificial general intelligent agents.
Finally, we also wish to express our concern about recent economic trends in the deep-tech industry, where we observe AI-related concentration of data and resources, coupled with a tendency towards closed-source practices. We strongly advocate for increased transparency and exchange of ideas to ensure a diverse and collective growth in our socio-economic landscape. Therefore, we push towards a democratic and open-source initiative.
7.4 Conclusion
In this work, we introduced SymbolicAI, a framework that unifies generative models with an array of solvers, blending the strengths of symbolic and sub-symbolic AI paradigms within a cohesive NeSy framework. SymbolicAI equips researchers and practitioners with a comprehensive toolkit to develop contextualized and explainable NeSy AI systems capable of addressing diverse challenges effectively. We also introduce a quality measure and a benchmark for comparing and evaluating a wide range of computational tasks. SymbolicAI provides a basis for further research in advanced program synthesis, hierarchical computational graphs, the development of self-referential systems, and the integration of probabilistic methods with AI design for creating autonomous agents. Furthermore, our commitment to open-source ideals attempts to promote a culture of collaborative growth and innovation.
Acknowledgement
The ELLIS Unit Linz, the LIT AI Lab, the Institute for Machine Learning, are supported by the Federal State Upper Austria. We thank the projects Medical Cognitive Computing Center (MC3), INCONTROL-RL (FFG-881064), PRIMAL (FFG-873979), S3AI (FFG-872172), DL for GranularFlow (FFG-871302), EPILEPSIA (FFG-892171), AIRI FG 9-N (FWF-36284, FWF-36235), AI4GreenHeatingGrids (FFG- 899943), INTEGRATE (FFG-892418), ELISE (H2020-ICT-2019-3 ID: 951847), Stars4Waters (HORIZON-CL6-2021-CLIMATE-01-01). We thank Audi.JKU Deep Learning Center, TGW LOGISTICS GROUP GMBH, Silicon Austria Labs (SAL), FILL Gesellschaft mbH, Anyline GmbH, Google, ZF Friedrichshafen AG, Robert Bosch GmbH, UCB Biopharma SRL, Merck Healthcare KGaA, Verbund AG, GLS (Univ. Waterloo), Software Competence Center Hagenberg GmbH, Borealis AG, TÜV Austria, Frauscher Sensonic, TRUMPF, the NVIDIA Corporation and Atlas.
We extend our appreciation to Andreas Windisch and Clemens Wasner of AI Austria for their unwavering support. Their valuable feedback, connections, and facilitation of introductions within their expansive network have been instrumental to the progress of ExtensityAI.
Our gratitude also goes to Sergei Pereverzyev, whose enlightened guidance and thoughtful ideas have been a beacon for our research endeavors. Our thanks are equally extended to Gary Marcus, whose stimulating discussions sparked numerous innovative ideas incorporated into our framework.
We are equally grateful to Markus Hofmarcher, a friend and colleague whose informed counsel and stimulating discussions have significantly sharpened various facets of our study. Additionally, our thanks are due to Fabian Paischer and Kajetan Schweighofer, whose preliminary work and assistance have been of enormous benefit.
We are also grateful to Tim Scarfe, a friend whose community has been a hub for exhilarating discussions. His online presence and engagement have enriched the AI research landscape and broadened our perspectives.
Moreover, we wish to honor the memories of the cherished family members we lost in 2023. Their influence in our lives extended beyond personal bonds, and the principles they instilled in us continue to shape our journey. It is with great respect and affection that we acknowledge the indelible impact they have made, enabling us to persist in our scientific pursuits with determination and integrity.
Appendix
Appendix A Connection between Fréchet Distance and Maximum Mean Discrepancy
Let us consider a Gaussian kernel defined by the expression
$K(x,y)=\exp\left(-\frac{\|x-y\|^{2}}{2\sigma^{2}}\right),$ (4)
[…]
More precisely, the MMD between two probability distributions $P$ and $Q$ is encoded in the RKHS through mean embeddings, which can be expressed as
$\text{MMD}^{2}(P,Q)=\left\|\mathbb{E}_{x\sim P}[\phi(x)]-\mathbb{E}_{y\sim Q}[\phi(y)]\right\|^{2}_{\text{RKHS}},$ (5)
where $\phi(\cdot)$ represents the feature mapping to the RKHS corresponding to the kernel $K$.
On the other hand, for multivariate Gaussian distributions, we can use the Fréchet distance as a measure of similarity, which is nothing but the associated Wasserstein-2 distance, for which an explicit formula is available in the Gaussian case. The resulting expression is as follows (Dowson & Landau, 1982):
$d^{2}(X_{1},X_{2})=\|\mu_{1}-\mu_{2}\|^{2}_{2}+\text{Tr}\left(C_{1}+C_{2}-2\left(C_{1}C_{2}\right)^{\frac{1}{2}}\right),$ (6)
where $X_{1}\sim\mathcal{N}(\mu_{1},C_{1})$ and $X_{2}\sim\mathcal{N}(\mu_{2},C_{2})$, and $\text{Tr}(\cdot)$ indicates the trace of a matrix.
To establish an approximation of the Fréchet distance using the Gaussian kernel, we take $C_{1}=\sigma^{2}I$ and $C_{2}=\sigma^{2}I$ as identity covariance matrices scaled by $\sigma^{2}$. This assumption allows us to focus solely on the disparities in mean vectors:
$d^{2}(X_{1},X_{2})\approx\|\mu_{1}-\mu_{2}\|^{2}_{2},$ (7)
setting aside the effect of different covariance structures.
Given these conditions, we attempt to argue that the Fréchet distance behaves similarly as MMD:
$d^{2}(X_{1},X_{2})\approx\|\mu_{1}-\mu_{2}\|^{2}_{2}\approx\text{MMD}^{2}(P,Q),$ (8)
Heuristically, at least for small $\|\mu_{1}-\mu_{2}\|$, also the associated kernel evaluations $K(X_{1},X_{2})$ tend to be small (see also Hochreiter & Schmidhuber (1997)), which leads to a small MMD, if we ignore the terms associated to $K(X_{1},X_{1})$, $K(X_{2},X_{2})$ (which cancel out due to same covariance structure).
In the next section we want to further elaobarate on the MMD and a possible score, that can be derived from it.
A.1 Extended Simplification of the MMD Calculation
To understand the simplification of the MMD, we are formally expressing the MMD in terms of kernel sums over pairs of samples within and across two samples $X$ and $Y$:
$\text{MMD}^{2}(X,Y)=\frac{1}{m(m-1)}\sum_{i}\sum_{j\neq i}k(x_{i},x_{j})-\frac{2}{mn}\sum_{i=1}^{m}\sum_{j=1}^{n}k(x_{i},y_{j})+\frac{1}{n(n-1)}\sum_{i}\sum_{j\neq i}k(y_{i},y_{j}),$ (9)
where $m$ and $n$ are the sizes of samples $X$ and $Y$, respectively.
Empirical observations have led to the conclusion that the within-sample terms $\sum_{i}\sum_{j\neq i}k(x_{i},x_{j})$ and $\sum_{i}\sum_{j\neq i}k(y_{i},y_{j})$ cancel out the cross terms $\sum_{i=1}^{m}\sum_{j=1}^{n}k(x_{i},y_{j})$ under certain conditions. This can be due to the following:
  • In high-dimensional embedding spaces, distributions of embedding vectors are often closely related and normally distributed.
  • If the samples $X$ and $Y$ are drawn from distributions $P$ and $Q$ where their mean embeddings are nearly orthogonal in the RKHS, it is the dissimilarity across samples, rather than that within, that is most relevant.
Therefore, under these specific conditions, it becomes justifiable to focus on the cross-terms, yielding the following proposal for a distance measure:
$\widetilde{\text{MMD}^{2}}(X,Y)\approx\frac{2}{mn}\sum_{i=1}^{m}\sum_{j=1}^{n}k(x_{i},y_{j}).$ (10)
Appendix B Broad AI and Neuro-Symbolic Systems
Our work focuses on broad artificial intelligence (AI) (Hochreiter, 2022) (see Figure 6) through the integration of symbolic and sub-symbolic AI methodologies. Broad AI extends beyond restricted focus on single-task performance of narrow AI. In broad AI, systems are engineered to handle a range of tasks with a high degree of autonomy, utilizing sensory input, accumulated experiences, and previously developed skills.
Figure 6: Integration of symbolic and sub-symbolic AI in broad AI Systems, showcasing central concepts of broad AI systems, including knowledge, adaptability, abstraction, interaction, interpretability, efficiency, robustness, and reasoning (Hochreiter, 2022).
Our primary objective is to combine the strengths of symbolic and sub-symbolic approaches to overcome individual limitations. Symbolic AI is characterized by its emphasis on knowledge representation, the ability to abstract and formulate mathematical concepts, and the capacity for interactions with users or other systems in a human-understandable manner. These attributes ensure that we develop reasoning-based, interpretable AI systems with innate robustness and trustworthiness (Winter et al., 2021).
Both methodologies form the basis for developing new cognitive architectures, as illustrated in Figure 7. This hybridization produces computational graphs capable of context-aware learning and reasoning, allowing AI to execute complex tasks with human-like flexibility.
Figure 7: Evolution of software paradigms: From Software 1.0’s rigid specification in classical programming to Software 2.0’s data-driven and objective function-focused differentiable programming, leading to Software 3.0’s NeSy systems that emphasize human-centric, interaction-based programming with computational graphs. This progression represents a shift from explicit task-specific programming to abstract, adaptive systems that cater to dynamic user preferences.
Borrowing nomenclature from Karpathy (2017); Dilhara et al. (2021), we refer to the next generation of software as Software 3.0, which consists of applications that are neither pre-determined at design time, nor learned through statistical inference, but triggered by an interaction which stimulates the realization of a computational graph analogous to neuromorphic circuits (Indiveri et al., 2011), however, purely established at inference time in the ”thought” process of a NeSy system.
To enable such systems, we require a more native integration (see illustration in Figure 8) of probabilistic programming paradigms into our contemporary programming stack, and make their utilization a commodity for practitioners and researchers alike.
Figure 8: The illustration demonstrates the integration of Neuro-Symbolic computation within the contemporary programming stack. Probabilistic programming paradigms are embedded natively alongside traditional languages and environments, facilitated by interfaces to various data sources, compute engines, and human interaction tools, streamlining their adoption in practical and research applications.
Appendix C Framework Structure
Primitives
In the SymbolicAI framework, at the core lies the concept of Primitives and the dynamic type creation of Symbol objects, which are central to inherit types of behaviors. Primitives are pre-defined operations that act on Symbol objects, encapsulating basic operations, such as arithmetic, logic, or casting operations, to name a few. These operations are crucial to the framework’s versatility and form the foundation for more complex interactions within computational graphs. Essentially, they can be viewed as contextualized functions that accept a Symbol object, send it to the NeSy engine for evaluation, and return one or more new objects, primarily new symbols. One of the key features of operations is their polymorphism, which allows for them to be applied to various data types, such as strings, integers, floats, lists, and more, with different behaviors depending on the specific object instance. To execute operations, we utilize the Symbol object’s value attribute containing the original data type. This will be then sent as a string representation to the engines to execute the needed operations. Consequently, all values are cast to a string representation. Remember, this was our implicit assumption (see Section 4). For custom objects, it is essential to define a suitable __str__ method to cast the object to a string representation while preserving the object’s semantics.
Symbol Objects Creation and Dynamic Typing
A Symbol object is a versatile entity that can encapsulate a variety of data types and behaviors. The creation of Symbol objects is facilitated through a metaclass, which enables the dynamic typing of these objects to inherit behaviors from a collection of primitives. This dynamic typing system is important for extending the functionality of Symbol objects beyond simple data containers; they contain specific behaviors appropriate for the operations they will perform. For instance, a Symbol object may possess the behaviors of arithmetic computations, string manipulations, or even logical comparisons, depending on the defined primitives.
Type Inheritance and Expression Creation
Type inheritance in SymbolicAI is leveraged to create new expressions, which are specialized forms of Symbol objects designed to represent parts of a computational graph. Expressions extend the capabilities of Symbol objects by providing a structured way to define complex functionalities that can later be evaluated to produce new Symbol objects or modify existing ones. In SymbolicAI, an Expression object inherits the properties of Symbol objects while also being able to define its own unique behavior through the use of a forward method, which is analogous to a computational graph node’s evaluation function. Figure 9 gives an overview of an exemplary inheritance branch. Each Expression must feature a forward method, which must be overwritten to define its behavior. The inherited __call__ method invokes the forward method, evaluating the expression and returning the result. This design pattern facilitates lazy evaluation of expressions, allowing for complex composition of expressions.
Inherited from the Symbol class, the _sym_return_type and static_context properties establish the context in which the current Expression operates. The static_context impacts all operations of the current Expression subclass, while the _sym_return_type guarantees the acquisition of the desired return object type post-evaluation. Typically, this returns the current type but can be configured to return a different type. A more in-depth examination of both notions will be provided in the following section.
Figure 9: Class diagram showing the inheritance and composition relationships among Symbol, Expression, and other inherited classes. Symbol serves as a base class for Expression where all the other types are derived from. Other types may contain or associate with zero or more Symbol types. For example, we illustrate how the Sequence derives from Expression and the multiplicity ’0..*’ indicates that a Sequence can contain any number of Expression instances.
Utilizing Decorators for Operation Definition
Decorators serve as a bridge between the declarative nature of symbolic operations and the imperative execution model of programming languages. By augmenting function definitions with decorators, the framework can dynamically assign operations to Symbol or Expression objects, which are then interpreted by the underlying NeSy engine or traditional solvers.
For example, the @core.logic decorator can be used to augment a Symbol object with the capability to perform logical and, or, or not operations contextually. Similarly, the @core.combine decorator allows the framework to define the semantics of combining or adding two symbolic values, regardless of their underlying data representations.
1# Example of using decorators to define logical operations
2@core.logic(operator=’and’)
3def _some_logic(self, other):
4  # implementation logic here
5  pass
Aspect-Oriented Programming
The aspect-oriented programming paradigm offers a functional approach for extending or modifying the behavior of functions or methods without altering their code directly. This adheres to the principles of modularity and separation of concerns, as it allows for the isolation of specific functionalities while maintaining the original function’s core purpose. By wrapping the original function, decorators provide an efficient and reusable way of adding or modifying behaviors. For instance, SymbolicAI integrates the zero- and few-shot learning with default fallback functionalities of pre-existing code.
The use of decorators brings several advantages (Beazley, 2009; Martelli et al., 2005; Summerfield, 2010; Lutz, 2013):
  • Reusability: Decorators promote code modularity, enhancing code reusability and contributing to software maintainability. This advantage is particularly salient when managing a variety of operations, reducing redundancy and simplifying the integration of new functionalities.
  • Composition: Decorators support function composition, allowing developers to construct complex functionalities from pre-existing code blocks without the need to expand the codebase or rely on complex inheritance hierarchies.
  • Adaptability: Using decorators we can easily modify or extend the behavior of operations without changing their core implementation. This flexibility facilitates the generation of adaptive workflows and reliable fallback mechanisms when experimental implementations do not fulfill required constraints.
Symbol Class and Computational Graph Elements
A computational graph in the SymbolicAI framework is an assembly of interconnected Symbol objects, each encapsulating a unit of data and the operations that can be performed on it. The exchange between these symbols forms a highly modular and interpretable system, capable of representing complex workflows.
The Symbol class is an abstraction representing data and context. It holds not only the value itself, but metadata that guides its transformation and interpretation. Through inheritance and compositionality, the Symbol can be extended into more complex expressions, and becoming nodes in a computational graph. Each Symbol instance can optionally contain a reference to its parent and children, naturally forming a directed graph structure where the nodes are symbols and edges represent relationships between a symbol and its derivative computations.
The Linker class, is a metadata subclass, and tracks relationships and results, effectively annotating the graph with execution details. It keeps records of nodes’ keys, allowing quick retrieval of related computational outcomes within the graph, and aids in tasks such as error tracing and debugging.
A central concept in this structure is the notion of root, which points to the origin of the computational sequence. Accessing the root allows backtracking through the graph, making it possible to aggregate results and inspect the flow of computation that led to the current node.
The computational graph’s structure is further enriched by properties like nodes, edges, and graph itself, which collectively enable the comprehensive query of the computation’s topology. These properties are used internally to enable graph visualizations, which are useful for debugging and analysis.
Expression of a Computational Graph
In practice, consider the Expression class, which extends the functionality of ‘Symbol‘. When composing a Sequence of Expression objects, we are effectively composing operations in a predetermined order.
For instance, an expression such as:
1Sequence(
2  Clean(),
3  Translate(),
4  Outline(),
5  Compose(),
6)
represents a procedure that first cleans data, then translates it, outlines the essential information, and composes it into a finalized form. When this sequence is executed, the operations unfold in the exact order specified, with each step receiving the output of its predecessor as input and passing its result to the successor.
Generating entire workflows
SymbolicAI allows the creation of entire workflows, such as writing scientific papers. The following example defines a Paper expression that takes in a sequence of expressions which are executed in sequentially. The Method expression contains a Source expression, which addresses the code base of the actual method of the paper. The RelatedWork expression contains a sequence of Cite expressions which are executed in parallel and are used to define the context of the related work section. Each expression in the sequence of expressions from Paper takes in the context of its predecessors. All expressions also link their results to a global linker object, which is used after the execution to retrieve individual results from distinct expression nodes. In Figure 10 we show the resulting computational graph of the Paper expression.
1# define the computational graph
2expression = Paper(
3  Method(
4    # link to original code base where the main method is defined
5    Source(file_link=’/path/to/.../file’),
6  ),
7  # gather resources and write the related work
8  RelatedWork(
9    Cite(file_link=’/path/to/.../file’),
10    Cite(file_link=’/path/to/.../file’),
11    ...
12  ),
13  # write the abstract and title
14  Abstract(),
15  Title(),
16)
17# run the graph
18paper_result = expression(’Write a scientific paper’)
19# access linker to retreive the results from the method expression
20method = expr.linker.find(’Method’)
21# print result of the method expression
22print(method)
Figure 10: We illustrate the computational graph for the Paper expression. Each node represents an instance of an expression with distinct properties and behaviors. The edges denote the reference relationship between expressions. The blue highlighted nodes mark the main sequence of expressions, used to create the paper, such as Method, RelatedWork, Abstract, etc.
Appendix D Installation
The installation of the SymbolicAI framework is straightforward and can be done through the Python package manager pip. To install SymbolicAI, open a terminal and execute the following command in your current python environment:
1pip install symbolicai
This command will install the latest version of SymbolicAI along with its core dependencies, enabling the integration of the framework into Python applications. If you intend to use the framework with local engines444 The local engines are experimental and are run on your local machine. For more details, refer to the ”Local Neuro-Symbolic Engine” section in the documentation., or with engines powered by external APIs such as OpenAI’s API, additional installation steps are required.
D.1 Engine Configuration
Before the first run, it is necessary to configure the required modules and optionally set necessary API keys to activate the respective engines. This can be done in multiple ways, but we recommend doing it through the configuration wizard by running this command in the terminal:
1symwzd
This step is essential to register the engines internally for subsequent runs.
For instance, SymbolicAI includes OpenAI’s GPT models as NeSy engine. To only set or change OpenAI API keys, the following command is used before starting a SymbolicAI instance:
1# Linux / MacOS
2export OPENAI_API_KEY="<OPENAI_API_KEY>"
After setting up the API keys, the SymbolicAI library is imported in Python using:
1import symai
For more low-level changes, we store everything under the \$HOME/.symai folder, such as the symai.config.json, which stores every key, both registered and not registered.
D.2 Optional Installations
The SymbolicAI framework is designed to leverage multiple engines for a variety of operations. To fully utilize these capabilities, you may install additional dependencies or set up the optional API keys for specific engines like WolframAlpha, SerpApi, and others. In Figure 11 we conceptually outline the connection between the utilization of an LLM and its interact with other tools and solvers. Instructions and operations can be initiated by any user, pre-scripted knowledge base or learned meta agent.
Figure 11: The SymbolicAI framework integrates a Large Language Model (LLM) with diverse tools and solvers through a conceptual interaction stack. The framework enables operations initiated by users, knowledge bases, or meta-learners to be processed by the LLM, which interfaces with specialized engines such as WolframAlpha and Whisper via conditions and constraints, enhancing the AI’s problem-solving capabilities.
For instructions on additional installations, including the support of optional engines, refer to the documentation provided with the framework. This documentation will give detailed steps on installing optional dependencies and configuring additional API keys.
D.3 Benchmark
To run the benchmark evaluation you first need to install the ExtensityAI/benchmark plugin:
1sympkg i ExtensityAI/benchmark
Then you can go to the \$HOME/.symai/packages/ExtensityAI/benchmark directory and run the benchmark through:
1python test.py
To compile models optimized for your machine, follow the instructions from the ExtensityAI/benchmark repository.
Appendix E Implementation Details
Let us now define some Symbol objects and perform some basic manipulations.
E.1 Fuzzy Comparison
For instance, let’s consider the use of fuzzy555 Not related to fuzzy logic, which is a topic under active consideration. comparisons. Within SymbolicAI, it enables more adaptable and context-aware evaluations, accommodating the inherent uncertainties and variances often encountered in real-world data.
1import numpy
2
3s = symai.Symbol(’3.1415...’)
4s == numpy.pi
1:[Output]:
2True
E.2 Dynamic Casting
By enabling sentence subtraction and dynamic casting within SymbolicAI, we utilize the generalization capability of NeSy engines to manipulate and refine text-based data, creating more meaningful and contextually relevant outcomes. The integration of dynamic casting with Symbol objects in our API allows the users to perform operations between Symbol objects and various data types, such as strings, integers, floats, lists, etc. without compromising on readability or simplicity.
1s = symai.Symbol(’Hello my enemy’)
2s - ’enemy’ + ’friend’
1:[Output]:
2<class ’symai.expressions.Symbol’>(value=Hello my friend)
E.3 Translation
In today’s increasingly interconnected world, translation between languages is fundamental, making it an essential feature.
1s = symai.Symbol("Welcome to our tutorial.")
2s.translate(’German’)
1:[Output]:
2<class ’symai.expressions.Symbol’>(value=Willkommen zu unserem Tutorial.)
E.4 Filtering, Ranking, Extraction
Incorporating data-agnostic operations like filtering, ranking, and pattern extraction into our API allow the users to easily manipulate and analyze diverse data sets.
1s = symai.Symbol(numpy.array([1, 2, 3, 4, 5, 6, 7]))
2s.rank(measure=’numerical’, order=’descending’)
1:[Output]:
2<class ’symai.expressions.Symbol’>(value=[’7’, ’6’, ’5’, ’4’, ’3’, ’2’, ’1’])
E.5 Implications
One of the main objectives behind developing SymbolicAI was to facilitate reasoning capabilities in conjunction with the statistical inference inherent in LLMs. Consequently, we can carry out deductive reasoning operations utilizing the Symbol objects. For instance, it is feasible to establish a series of operations with rules delineating the causal relationship between two symbols. The subsequent example illustrates the utilization of the & operator to compute the logical implication derived from the interaction of two distinct symbols.
1s1 = symai.Symbol(’The horn only sounds on Sundays.’)
2s2 = symai.Symbol(’I hear the horn.’)
3s1 & s2
1:[Output]:
2<class ’symai.expressions.Symbol’>(value=It is Sunday.)
In the above example, the & operator overloads the and logical operator and extends its functionality. Furthermore, we can establish more sophisticated logical operators for and, or, and xor that can be grounded in formal proofs, as well as utilize the NeSy engine to parse data structures before evaluating the expressions. This enables the definition of bespoke operations for executing intricate and robust logical operations, incorporating constraints to validate outcomes and guide the computation towards the desired behavior.
E.6 Custom operations
The following example demonstrates how to define a custom == operation by overriding the __eq__ method and providing a custom prompt object with a list of examples:
1import symai
2
3class Demo(symai.Symbol):
4  def __eq__(self, other) -> bool:
5    # define nested function
6    @symai.core.equals(examples=symai.Prompt([
7      "1 == ’ONE’ =>True",
8      "’six’ == 7 =>False",
9      "’Acht’ == ’eight’ =>True",
10      ...
11    ]))
12    def _func(_, other) -> bool: # [optional] cast return type           (1. below)
13      return False             # [optional] default behavior on failure  (2. below)
14    return _func(self, other)
As illustrated in the example, this is also the method we used to implement basic operations in Symbol, namely by defining local functions that are then decorated with the respective operation decorator from the symai.core.py file. The symai.core.py is a collection of pre-defined operation decorators that can be quickly applied to any function. We use locally defined functions instead of directly decorating the main methods for two reasons:
  1. 1.
    We want to cast return types of the operation outcome to symbols or other derived classes thereof.
  2. 2.
    We do not necessarily want all of our operations to be sent to the NeSy engine and might need to implement a default behavior.
This is achieved using the _sym_return_type method, which can provide contextualized behavior based on the defined return type. More details can be found in the actual Symbol class.
In the context of LLMs, zero- and few-shot learning domains have emerged as essential techniques (Yao et al., 2023b; Shinn et al., 2023; Kim et al., 2023; Wei et al., 2022b; Lyu et al., 2023; Pitis et al., 2023; Madaan et al., 2023; Wang et al., 2022; Ye et al., 2023)666 This is by no means an exhaustive list, we only point the reader to some very influential and recent research. to enable models to generalize from limited training data and adapt to new tasks without requiring extensive retraining. This capability to learn and perform tasks with minimal examples is highly desirable, as it reduces the need for large labeled data sets and allows for faster deployment in new applications. In this section, we demonstrate how our Symbolic API incorporates Python decorators to define custom operations in the zero- and few-shot domains.
Consider the following example, where we define a custom operation to generate a random integer between 0 and 10 using the Symbolic API and Python decorators:
1import symai
2
3class Demo(symai.Symbol):
4  def __init__(self, value = ’’) -> None:
5    super().__init__(value)
6
7  @symai.core.zero_shot(prompt="Generate a random integer between 0 and 10.",
8                        constraints=[
9                          lambda x: x >= 0,
10                          lambda x: x <= 10
11                        ])
12  def get_random_int(self) -> int:
13    pass
In this example, the @symai.core.zero_shot decorator is used to define a custom operation that does not require any examples, as the prompt is expressive enough. The zero_shot decorator takes in two arguments: prompt and constraints. The prompt defines the conditioning for our desired operation behavior, while the constraints are used to validate the computed outcome, ensuring it meets our expectations. If the constraints are not fulfilled, the implementation would resort to the specified default implementation or the default value. If neither is provided, the Symbolic API raises a ConstraintViolationException. The return type in the example is defined as int. The resulting value from the wrapped function must be of type int because of the specific implementation of the auto-casting realized through ->. If the cast fails, the Symbolic API raises a ValueError. If no return type is specified, the return type defaults to Any.
The @symai.core.few_shot decorator is a generalized version of @symai.core.zero_shot and is used to define custom operations requiring examples. The function signature of the few_shot decorator is as follows:
1def few_shot(prompt: str,
2             examples: Prompt,
3             constraints: List[Callable] = [],
4             default: Any = None,
5             limit: int = 1,
6             pre_processor: Optional[List[PreProcessor]] = None,
7             post_processor: Optional[List[PostProcessor]] = None,
8             **wrp_kwargs):
The behavior of the prompt and constraints attributes is similar to the zero_shot decorator. The examples and limit arguments are new, with examples defining a list of instructions conditioning the NeSy engine, and limit specifying the maximum number of examples returned. The pre_processor and post_processor arguments accept lists of PreProcessor and PostProcessor objects, respectively, which are used to process the input before being fed into the NeSy engine and the output before being returned to the user. The wrp_kwargs argument passes additional arguments to the wrapped method, streamlining them towards the NeSy engine, or other engines.
E.7 Prompting
In this section, we discuss the design of prompts and their role in shaping the behavior of operations. Prompts serve as containers for information necessary to define specific operations, and the Prompt class serves as the base class for all the other Prompt classes. Consider the following example, where we define a Prompt for comparing two values using the NeSy engine. In it, when the <= operation on two Symbol objects will be resolved, the NeSy engine evaluates them in the context of the CompareValues prompt.
1class CompareValues(symai.Prompt):
2  def __init__(self) -> symai.Prompt:
3    super().__init__([
4      "4 > 88 =>False",
5      "-inf < 0 =>True",
6      "inf > 0 =>True",
7      "4 > 3 =>True",
8      "1 < ’four’ =>True",
9      ...
10    ])
1res = symai.Symbol(1) <= symai.Symbol(’one’)
1:[Output]:
2True
This evaluation returns True, as the fuzzy comparison operation conditions the engine to compare the two Symbol objects based on their semantic meaning. More generally, the semantics of Symbol operations may vary depending on the context hierarchy of the Expression class and the operations used. We used three main prompt designs: Context-based Prompts, Operational Prompts, and Templates. Prompts can be curated either through inheritance or composition. For instance, the static context can be defined by inheriting from the Expression class and overwriting the static_context property, while an Operation and Template prompt can be created by providing a PreProcessor to modify the input data.
We will now provide a more detailed explanation for each prompt design:
  1. 1.
    Context-based Prompts are considered optional and can be defined in a static manner, either by sub-classing the Expression class and overwriting the static_context property, or at runtime by updating the dynamic_context property or passing a payload kwargs to a method. Below is an example of using the payload kwargs through the method signature:
    1# creating a query to ask if an issue was resolve or not
    2s = symai.Symbol("<some-community-conversation>")
    3q = s.query("Was the issue resolved?")
    4# write manual condition to check if the issue was resolved
    5if ’not resolved’ in q:
    6  # do a new query but payload the previous query answer to the new query
    7  s.query("What was the resolution?", payload=q)
    8  ...
    9else:
    10  pass # all good
    Regardless of how the context is set, the contextualized prompt defines the desired behavior of Expression operations. For instance, if we want to operate in the context of a DSL without having to overwrite each base class method, we can use this approach777 See more details in this notebook..
  2. 2.
    Operational Prompts define the behavior of an atomic operation and are therefore mandatory to express the nature of such an operation. For example, the + operation is used to add two Symbol objects together, and the + operation prompt explains its behavior. The examples kwargs provide another optional structure that conditions the NeSy engine with a set of instructions.
  3. 3.
    Template Prompts are optional and encapsulate the resulting prediction to enforce a specific format. For example, to generate HTML tags, we can use a curated <html>...</html> template. This template enforces the NeSy engine to begin the generation process already in the context of an HTML tags format and not produce irrelevant descriptions about its task.
E.8 Complex expressions
We will now attempt to obtain logical answers based on questions of the kind:
  • A line parallel to $y=4x+6$ passes through $(5,10)$. What is the $y$-coordinate of the intercept?
  • Bob has two sons, John and Jay. Jay has one brother and father. The father has two sons. Jay’s brother has a brother and a father. Who is Jay’s brother?
  • Is 1000 bigger than 1063.472?
To solve these tasks, we would initially employ a series of operations to identify the most suitable engine for handling the specific requirements. Subsequently, we would prepare the input tailored to the selected engine.
1val = "<one of the examples above>"
2
[…]
4class ComplexExpression(symai.Expression):
5  # write a method that returns the causal evaluation
6  def causal_expression(self):
7    pass
8
9# instantiate an object of the class
10expr = ComplexExpression(val)
11# set WolframAlpha as the main expression engine to use
12expr.command(engines=[’symbolic’], expression_engine=’wolframalpha’)
13# evaluate the expression
14res = expr.causal_expression()
The implementation of causal_expression could in principle look like this:
1def causal_expression(self):
2  # very which case to use ‘self.value‘ contains the input
3  if self.isinstanceof(’mathematics’):
4    # get the mathematical formula
5    formula = self.extract(’mathematical formula’)
6    # verify which problem type we have
7    if formula.isinstanceof(’linear function’):
8      # prepare for wolframalpha
9      question = self.extract(’question sentence’)
10      req = question.extract(’what is requested?’)
11      # get coordinate point / could also ask for other points
12      x   = self.extract(’coordinate point (.,.)’)
13      # concatenate to the question and formula
14      query = formula | f’, point x = {x}’ | f’, solve {req}’
15      res = self.expression(query) # send prepared query to wolframalpha
16
17    elif formula.isinstanceof(’number comparison’):
18      res = formula.expression() # send directly to wolframalpha
19
20      ... # more cases
21
22  elif self.isinstanceof(’graph construction’):
23    sentences = self / ’.’ # first split into sentences
24    graph = {} # define graph
25    for s in sentences:
26      sym = symai.Symbol(s)
27      relations = sym.extract(
28        # extract and split by pipe
29        ’connected entities (e.g. A has three B => A | A: three B)’) / ’|’
30      for r in relations:
31        ... # add relations and populate graph; reading suggestion
32
33  ... # more cases
34
35  return res
The aforementioned example demonstrates the utilization of the causal_expression method, which allows us to extract information that can be resolved either manually or through external solvers, say WolframAlpha. Initially, when utilizing the GPT-3 backend, we anticipated a significant engineering effort to develop such a complex expression, as the GPT-3 backend frequently struggled with accurate information extraction and comparison resolution. However, we remained confident in the field’s progress, specifically with fine-tuned models like ChatGPT utilizing RLHF. We were delighted to witness these challenges being further tackled through the latest GPT-4 model (OpenAI, 2023).
Furthermore, it is worth highlighting that, given sufficient data, we could refine methods for information extraction or knowledge graph construction from natural language, enabling more intricate reasoning tasks, such as those previously mentioned. We also direct readers to recent publications on Text-to-Graph translations, especially the very influential CycleGT (Guo et al., 2020). This approach allows us to answer queries by simply traversing the graph and extracting the required information.
Lastly, recent research (Kıcıman et al., 2023; Ellis, 2023) has demonstrated that algorithms based on GPT-3.5 and GPT-4 establish new state-of-the-art accuracy on multiple causal benchmarks, while also exhibiting unique capabilities previously considered exclusive to humans, such as generating causal graphs and identifying background causal context from natural language. This points to the potential for LLMs to be used alongside existing causal methods as proxies for human domain knowledge, reducing human effort in setting up causal analyses and ultimately accelerating the widespread adoption of causal methods. Moreover, recent advances in LLMs have opened new frontiers for research, practice, and adoption of causal reasoning, transforming the way causal analysis is conducted and broadening the scope of applications for our framework.
One of the most prominent illustrations of this concept is exhibited by Word2Vec (Mikolov et al., 2013). Word2Vec generates dense representations of words by training a shallow neural network to predict a word based on its neighboring words within a text corpus. These resulting vectors are extensively utilized in various natural language processing applications, including sentiment analysis, text classification, and clustering.
Drawing parallels with Word2Vec, our objective is to execute contextualized operations on different symbols. However, the key distinction lies in the fact that we operate within the natural language domain, as opposed to a vector space. Consequently, this grants us the capability to conduct arithmetic on words, sentences, paragraphs, and the like, while simultaneously validating the outcomes in a human-readable format.
The following example, we illustrate the methodology for evaluating such an expression through a string representation:
1s = symai.Symbol(’King - Man + Woman’)
2s.expression()
1:[Output]:
2<class ’symai.expressions.Symbol’>(value=Queen)
In contrast to the Symbol object, the Expression represents a non-terminal symbol. It allows for further evaluation and extends the Symbol class by overwriting the __call__ method. It serves as the foundation for all other expressions and possesses additional capabilities, namely to fetch data from URLs, search the internet, or open files. These operations are intentionally separated from Symbol, as they do not utilize the value attribute of the Symbol class.
E.9 Composition
E.10 Sequences
Sequences offer a multitude of advantages in the realm of Expression objects, as they facilitate the creation of more sophisticated structural configurations. By embodying the Sequence expression, multiple expressions can be effectively evaluated at runtime, thus enhancing the flexibility, modularity, and adaptability of the framework.
1# first import all expressions
2from symai.components import *
3# define a sequence of expressions
4Sequence(
5  Clean(),
6  Translate(),
7  Outline(),
8  Compose(),
9)
E.11 Streams
As demonstrated earlier, creating contextualized prompts refines the behavior of operations in the NeSy engine. However, this also consumes a considerable portion of the available context size. Given a limited context size, this constraint may pose challenges. Fortunately, the Stream processing expression offers a solution by opening a data stream and performing chunk-based operations on the input stream. Stream expressions can encapsulate other expressions. For instance, chunks can be managed using a Sequence expression, which permits multiple compositional operations in a sequential manner. The example below illustrates the definition of a Stream expression:
1Stream(
2  Sequence(
3    Clean(),
4    Translate(),
5    Outline(),
6    Embed()
7  )
8)
In this case, a stream is opened and a Sequence expression is passed, which cleans, translates, outlines, and embeds the input. Internally, the stream operation estimates the available model context size and segments the lengthy input text into smaller chunks transmitted to the inner expression. The returned object type is a generator.
The limitation of this approach is that the resulting chunks are processed independently, lacking shared context or information among them. To address this, the Cluster expression can be employed, merging the independent chunks based on their similarity, as it illustrated in Figure 12.
Figure 12: Stream processing expression in NeSy engine, illustrating data stream segmentation into chunks, each undergoing operations like cleaning, outlining, and embedding. The Cluster expression then merges chunks based on similarity, allowing contextually related information to be consolidated meaningfully. Node summaries are generated by extracting key labels from each cluster’s content, overcoming context size limitations and maintaining shared information among processed chunks.
By merging individual chunks by clustering their contents, contextually related information can be consolidated in a meaningful manner. Additionally, the clustered information can be labeled by streaming through each cluster’s content and extracting the most pertinent labels, yielding interpretable node summaries.
The complete example is depicted as follows:
1stream = Stream(
2  Sequence(
3    Clean(),
4    Translate(),
5    Outline(),
6  )
7)
8
9s    = symai.Symbol(’<some long text>’)
10res  = symai.Symbol(list(stream(s)))
11expr = Cluster()
12expr(res)
Subsequently, this process can be recursively repeated on each summary node to construct a hierarchical clustering structure. As each node represents a summarized subset of the original information, the summary can function as an index. The resulting tree can be utilized to navigate and retrieve the original information, transforming the large data stream problem into a search problem. Alternatively, vector-based similarity searches can be employed to identify similar nodes. For searching within a vector space, dedicated libraries such as Annoy (Spotify, 2017), Faiss (Johnson et al., 2019), or Milvus (Wang et al., 2021a) can be used.
In summary, Stream expressions offer the advantage of processing large data streams in a more efficient and organized manner, while also enabling the integration with other expressions like Sequence and Cluster expressions. These combinations allow for a more effective approach to handling context limitations, promoting the extraction of meaningful information and improving the overall performance of the framework.
E.12 Error handling, debugging, and explainability
Effective error handling and debugging are essential for ensuring the robustness and reliability of any software system, while explainability is essential for understanding the underlying behavior of the system, especially in the context of AI-driven applications. By developing a system that is both transparent and interpretable, we can gain valuable insights into the performance of the NeSy engines and identify potential areas for improvement.
E.13 Error handling
One of the fundamental aspects of the SymbolicAI API is being able to generate code. Consequently, errors may arise, and handling them contextually becomes vital. In pursuit of a self-evolving API, we introduce the Try expression, which includes built-in fallback statements and automatically retries execution after performing dedicated error analysis and correction. This expression analyzes both the input and the error, conditioning itself to resolve the error by modifying the original code888 This is similar to the recently released Auto-GPT application.. If the fallback expression succeeds, the result is returned; otherwise, the process is repeated for the specified number of retries. If the maximum number of retries is reached without resolving the issue, the error is raised again.
Consider the example of executing previously generated code that contains a syntax error. By employing the Execute expression, we can evaluate the generated code, which takes a symbol and proceeds with the execution. Despite the initial failure, the Try expression resolves the syntactic error, returning the corrected and evaluated code:
1expr = Try(expr=Execute())
2s    = symai.Symbol(’a = int("3,")’) # some code with a syntax error
3expr(s)
1:Output:
2a = 3
While not all errors can be resolved as easily as the demonstrated syntax error, we continue to explore more sophisticated error handling mechanisms, including the use of streams and clustering to address errors in a hierarchical and contextual manner.

### v2

@@
 are linked below.
-Figure 1: Our neuro-symbolic framework enables a seamless transition between symbolic and differentiable programming, each with distinct dynamics and strengths.
+Figure 1: Our neuro-symbolic framework enables a seamless transition between classical and differentiable programming, each with distinct dynamics and strengths.
 Differentiable programming provides access to foundational and specialist models.
@@
 The recent surge in generative AI, particularly involving large language models (LLMs), has demonstrated their wide-ranging applicability across various domains (Badita, 2022; Degrave, 2022).
-These models have enhanced the functionality of tools for search-based interactions (Microsoft, 2023; YouWrite, 2022; Writesonic, 2022), program synthesis (Jain et al., 2021; Romera-Paredes et al., 2023; Key et al., 2023), chat-based interactions (ReplikaAI, 2016; OpenAI, 2022; Google, 2023), and many more.
+These models have enhanced the functionality of tools for search-based interactions (YouWrite, 2022; Writesonic, 2022; Microsoft, 2023), program synthesis (Jain et al., 2021; Romera-Paredes et al., 2023; Key et al., 2023), chat-based interactions (ReplikaAI, 2016; OpenAI, 2022; Google, 2023), and many more.
 Moreover, language-based approaches have facilitated connections between different modalities, enabling text-to-image (Ramesh et al., 2021; Saharia et al., 2022), text-to-video (Singer et al., 2022), text-to-3D (Poole et al., 2022), text-to-audio (Oord et al., 2016; Wang et al., 2017), and text-to-code (Wang et al., 2021b; Lu et al., 2021; Li et al., 2022b) transformations, to name a few.
@@
 However, these efforts only partially capture the vast potential inherent in leveraging LLMs as semantic parsers.
-In light of these considerations, we introduce SymbolicAI, a compositional neuro-symbolic (NeSy) framework able to represent and manipulate compositional, multi-modal, and self-referential structures (Schmidhuber, 2007; Fernando et al., 2023).
+In light of these considerations, we introduce SymbolicAI, a compositional neuro-symbolic (NeSy) framework able to represent and manipulate multi-modal and self-referential structures (Schmidhuber, 2007; Fernando et al., 2023).
 SymbolicAI augments the generative process of LLMs with functional zero- and few-shot learning operations and enables the creation of versatile applications through in-context learning (Wei et al., 2022a).
@@
 It exposes these solvers as building blocks for constructing compositional functions as computational graphs, and facilitates the development of an extensible toolkit that bridges classical and differentiable programming paradigms, aiming to create domain-invariant problem solvers.
-In designing the architecture of SymbolicAI, we drew inspiration from the body of evidence that suggests the human brain possesses a selective language processing module (Deniz et al., 2019; Hu et al., 2022; Menenti et al., 2011; Scott et al., 2016; Macsweeney, 2002; Fedorenko et al., 2010; Regev et al., 2013), prior research on cognitive architectures (Newell & Simon, 1956; Newell et al., 1957; Newell & Simon, 1972; Newell, 1990; Laird, 2022), and the significance of language on the structure of semantic maps in the human brain (Huth et al., 2016).
+In designing the architecture of SymbolicAI, we drew inspiration from the body of evidence that suggests the human brain possesses a selective language processing module (Macsweeney, 2002; Fedorenko et al., 2010; Menenti et al., 2011; Regev et al., 2013; Scott et al., 2016; Deniz et al., 2019; Hu et al., 2022), prior research on cognitive architectures (Newell & Simon, 1956; Newell et al., 1957; Newell & Simon, 1972; Newell, 1990; Laird, 2022), and the significance of language on the structure of semantic maps in the human brain (Huth et al., 2016).
 We consider language as a central processing module, distinct from other cognitive processes such as reasoning or memory (Paischer et al., 2022, 2023), that defines a stepping stone towards broad AI systems (see Section B).
@@
 various classical algorithms), offloading computational complexity and bridging various modalities.
-Secondly, it enables sub-symbolic generalization to focus on evidence-based decision-making.
+Secondly, it enables sub-symbolic generalization to focus on evidence-based decision-making (e.g.
+selecting the respective tool based on in-context classification).
 Thirdly, it provides an interpretable language-based control layer for explainable, autonomous systems.
@@
 4 Design Principles
-In the following subsections, we elaborate on the key design principles underlying SymbolicAI and how we guide the generative processes of NeSy engines.
+In the following subsections, we elaborate on the key design principles underlying SymbolicAI and how we guide the generative processes of NeSy engines. For further technical details, see Section 5.
 Symbols and Expressions
@@
 Recently, several in-context learning methodologies evolved to enable tool usage through LLMs (Schick et al., 2023), or refine the generative outcome of LLMs (Yang et al., 2023).
-This includes chain-of-thought (CoT) prompting, a method that conditions the model to reveal its step-by-step reasoning process (Singhal et al., 2023; Wei et al., 2022b).
+This includes chain-of-thought (CoT) prompting, a method that conditions the model to reveal its step-by-step reasoning process (Wei et al., 2022b; Singhal et al., 2023).
 CoT prompting breaks down complex tasks into simpler, sequential steps, and helps with interpreting LLM’s output.
@@
 The expression showcases the left shift operator $\ll$ and how the information of the symbol B is included in the symbol AC.
-The violet placeholder in the DSL Prompt represents an instruction, such as ”Insert the right-hand side value into the left-hand value in a chronological order.” The positions below that represent task-specific few-shot examples.
+The violet placeholder in the DSL Prompt represents an instruction, such as ”Insert the right-hand side value into the left-hand value in a chronological order.” The positions below represent task-specific few-shot examples.
 Domain-Invariant Associations
@@
 Furthermore, SubGD reflects the potential of in-context learning when combined with task-specific fine-tuning by showing that fine-tuning within a learned subspace significantly outperforms traditional fine-tuning methods.
-We believe that the extent of in-context learning is not yet exhausted, holding considerable promise when use alongside with task-specific fine-tuning and solvers.
+We believe that the extent of in-context learning is not yet exhausted, holding considerable promise when used alongside with task-specific fine-tuning and solvers.
 To develop learning and reasoning systems capable of general problem-solving, we adopt a hybrid methodology.
@@
 In Figure 3 we give an overview of the entire prompt composition based on the user input, the Symbol object structure, and the NeSy engine evaluation pipeline.
-Figure 3: The illustrated pipeline utilizes GPT-4 as a NeSy engine and maps all user input and object instances to prompts.
-Description from left to right; yellow marks user data, blue marks templating and transformations, red marks constraints, and violet the model prediction placeholder.
-The user arguments and Symbol instance values are evaluated and transformed according to the pre- and post-processing steps.
-A custom method is constructed by using a reference implementation of decorators and passes its user arguments and instance attributes to the respective engine.
-A pre-processor manipulates the representation of user arguments and Symbol instance attributes before forwarding them to the engine.
-For instance, in case the engine is an LLM (i.e.
-NeSy engine), the static and dynamic context of a Symbol instance are mapped to the system and user prompts for that respective LLM using the prepare method of the engine.
+Figure 3: We illustrate our NeSy engine pipeline and how we map user input and object instance attributes to a prompt instruction.
+Yellow marks user input, blue marks templating and transformations, red marks constraints, and violet the model prediction placeholder.
+The user input and Symbol instance attribute values are evaluated and transformed according to the pre- and post-processing steps.
+A custom method is constructed by using our reference implementation of decorators and passes the user input and instance attributes to the respective engine.
+A pre-processor manipulates the string representation of user input and Symbol instance attributes before forwarding them to the engine.
+For instance, in case the NeSy engine is an LLM (i.e.
+GPT-4), the static and dynamic context of a Symbol instance are mapped to the system and user prompts for that respective LLM using the prepare method of the engine.
 Other attributes, such as examples or templates, can help pre-define structures for the desired LLM output format.
@@
 Benchmarks GPT-4 GPT-3.5 Gemini-Pro LlaMA 2 Mistral Zephyr Random Associations 0.93 0.61 1.00 0.41 0.49 0.28 0.01 Modality 0.83 0.77 0.75 0.47 0.39 0.43 0.07 Code 0.78 0.58 0.53 0.10 0.19 0.13 0.00 Logic 0.41 0.44 0.39 0.44 0.46 0.43 0.00 Graphs 0.36 0.31 0.21 0.05 0.06 0.08 0.00 Total 0.66 0.54 0.58 0.30 0.32 0.27 0.02
-Figure 5: We evaluate GPT-3.5, GPT4, Gemini-Pro, LLaMA2-Chat 13B, Mistral 7B and Zephyr 7B on five benchmark categories: 1) Associative Prediction (Association) 2) Multi-modal Binding (Modality) 3) Program Synthesis (Code) 4) Functional Logic Components (Logic) and 5) Hierarchical Computational Graphs (Graphs).
+Figure 5: We evaluate GPT-3.5 Turbo (GPT-3.5), GPT-4 Turbo revision 1106 (GPT-4), Gemini-Pro, LLaMA2-Chat 13B, Mistral 7B and Zephyr 7B on five benchmark categories: 1) Associative Prediction (Association) 2) Multi-modal Binding (Modality) 3) Program Synthesis (Code) 4) Functional Logic Components (Logic) and 5) Hierarchical Computational Graphs (Graphs).
 We denote the similarity score such that higher is better.
@@
 However, our objective is to ultimately assess a non-sequential task execution model, allowing for dynamic insertion and out-of-sequence task execution.
+In addition, we are interested into exploring similarities of our work with Generative Flow Networks (Bengio et al., 2021a, b; Lahlou et al., 2023).
 7.3 Broader Impact
@@
 We believe this will enable the creation of autonomously self-evolving cognitive architectures (Langley et al., 2009; Dawid & LeCun, 2023; Sumers et al., 2023).
-We therefore see an inherent connection to generative design as an analogy for creating coherent and stable ”thought” computational graphs, and believe this paves the path toward broad AI systems (see Section B) and is a requirement for developing artificial general intelligent agents.
+We therefore see an inherent connection to generative design as an analogy for creating coherent and stable ”thought” computational graphs, and believe this paves the path toward broad AI systems (see Section B) and is a requirement for developing General AI agents.
 Finally, we also wish to express our concern about recent economic trends in the deep-tech industry, where we observe AI-related concentration of data and resources, coupled with a tendency towards closed-source practices.
@@
 Additionally, our thanks are due to Fabian Paischer and Kajetan Schweighofer, whose preliminary work and assistance have been of enormous benefit.
-We are also grateful to Tim Scarfe, a friend whose community has been a hub for exhilarating discussions.
-His online presence and engagement have enriched the AI research landscape and broadened our perspectives.
+We are also grateful to our friends John Chong Min Tan and Tim Scarfe, whose communities have been a hub for exhilarating discussions.
+Their online presence and engagement have enriched the AI research landscape and broadened our perspectives.
 Moreover, we wish to honor the memories of the cherished family members we lost in 2023.
@@
 In broad AI, systems are engineered to handle a range of tasks with a high degree of autonomy, utilizing sensory input, accumulated experiences, and previously developed skills.
-Figure 6: Integration of symbolic and sub-symbolic AI in broad AI Systems, showcasing central concepts of broad AI systems, including knowledge, adaptability, abstraction, interaction, interpretability, efficiency, robustness, and reasoning (Hochreiter, 2022).
-Our primary objective is to combine the strengths of symbolic and sub-symbolic approaches to overcome individual limitations.
-Symbolic AI is characterized by its emphasis on knowledge representation, the ability to abstract and formulate mathematical concepts, and the capacity for interactions with users or other systems in a human-understandable manner.
-These attributes ensure that we develop reasoning-based, interpretable AI systems with innate robustness and trustworthiness (Winter et al., 2021).
-Both methodologies form the basis for developing new cognitive architectures, as illustrated in Figure 7.
+Figure 6: Hierarchical model of ”cognitive” abilities of AI systems (Chollet, 2019; Hochreiter, 2022).
+The figure contrasts the emergent paradigm of Broad AI with current Narrow AI systems, showcasing Broad AI’s wider range of capabilities, such as knowledge transfer, interaction, adaptability, robustness, abstraction, advanced reasoning, and efficiency.
+Broad AI aims to mimic human cognitive adaptability and robustness through advanced methodologies like few-shot learning, self-supervised contrastive learning, and context-sensitive sensory processing.
+Notably, Broad AI applies principles such as conceptual short-term memory and modern Hopfield networks (Ramsauer et al., 2020) to better integrate context and memory, thus avoiding pitfalls like explaining away and short-cut learning.
+We acknowledge the potential of NeSy systems as a significant step towards AI systems capable of performing any cognitive task with human-like proficiency.
+NeSy methods form the basis for developing new cognitive architectures (Newell & Simon, 1956; Newell et al., 1957; Newell & Simon, 1972; Newell, 1990; Langley et al., 2009; Laird, 2022; Dawid & LeCun, 2023; Sumers et al., 2023; LeCun, 2022; Assran et al., 2023), as illustrated in Figure 7.
 This hybridization produces computational graphs capable of context-aware learning and reasoning, allowing AI to execute complex tasks with human-like flexibility.

### v3

@@
 We leverage probabilistic programming principles to tackle complex tasks, and utilize differentiable and classical programming paradigms with their respective strengths.
-The framework introduces a set of polymorphic, compositional, and self-referential operations for data stream manipulation, aligning LLM outputs with user objectives.
-As a result, we can transition between the capabilities of various foundation models endowed with zero- and few-shot learning capabilities and specialized, fine-tuned models or solvers proficient in addressing specific problems.
-In turn, the framework facilitates the creation and evaluation of explainable computational graphs.
-We conclude by introducing a quality measure and its empirical score for evaluating these computational graphs, and propose a benchmark that compares various state-of-the-art LLMs across a set of complex workflows.
+The framework introduces a set of polymorphic, compositional, and self-referential operations for multi-modal data that connects multi-step generative processes and aligns their outputs with user objectives in complex workflows.
+As a result, we can transition between the capabilities of various foundation models with in-context learning capabilities and specialized, fine-tuned models or solvers proficient in addressing specific problems.
+Through these operations based on in-context learning our framework enables the creation and evaluation of explainable computational graphs.
+Finally, we introduce a quality measure and its empirical score for evaluating these computational graphs, and propose a benchmark that compares various state-of-the-art LLMs across a set of complex workflows.
 We refer to the empirical score as the ”Vector Embedding for Relational Trajectory Evaluation through Cross-similarity”, or VERTEX score for short.
-The framework codebase111 SymbolicAI framework released on January 20th, 2023, on GitHub: https://github.com/ExtensityAI/symbolicai.
-and benchmark222 Evaluation benchmark released on February 1st, 2024, on GitHub: https://github.com/ExtensityAI/benchmark.
+The framework codebase111 SymbolicAI framework: https://github.com/ExtensityAI/symbolicai and benchmark222 Evaluation benchmark: https://github.com/ExtensityAI/benchmark
+     * Correspondence to: dinu@ml.jku.at, {marius, leo}@extensity.ai
+Work done outside of Amazon.
 are linked below.
-Figure 1: Our neuro-symbolic framework enables a seamless transition between classical and differentiable programming, each with distinct dynamics and strengths.
+Figure 1: Our neuro-symbolic framework enables a seamless transition between symbolic and differentiable programming, each with distinct dynamics and strengths.
 Differentiable programming provides access to foundational and specialist models.
@@
 Moreover, language-based approaches have facilitated connections between different modalities, enabling text-to-image (Ramesh et al., 2021; Saharia et al., 2022), text-to-video (Singer et al., 2022), text-to-3D (Poole et al., 2022), text-to-audio (Oord et al., 2016; Wang et al., 2017), and text-to-code (Wang et al., 2021b; Lu et al., 2021; Li et al., 2022b) transformations, to name a few.
-Therefore, by training on vast quantities of unlabelled textual data, LLMs have been shown to not only store factual knowledge (Petroni et al., 2019; Kassner et al., 2020) and approximate users’ intentions to some extent (Andreas, 2022), but also to unlock deep specialist capabilities through innovative prompting techniques (Nori et al., 2023).
-Yet, these applications merely scratch the surface of the transformation that language-based interactions are expected to bring to human-computer interactions in both the near and distant future.
-In part, instruction-based fine-tuning of LLMs through reinforcement learning from human feedback (Ouyang et al., 2022; Li et al., 2023) or direct preference optimization (Rafailov et al., 2023) has shown promising results dealing with value misalignment issues (Bradley Knox & Stone, 2008; MacGlashan et al., 2017; Christiano et al., 2017; Ibarz et al., 2018; Goyal et al., 2022), unlocking new possibilities for chain of thoughts (Wei et al., 2022b), tree of thoughts (Yao et al., 2023a), and graph of thoughts interactions (Besta et al., 2023).
-However, recent research also highlights the limitations of LLMs in functional linguistic competence despite their proficiency in formal linguistic competence (Mahowald et al., 2023).
-Whereas formal linguistic competence encompasses the ability to understand and generate language, functional linguistic competence pertains to the application of language in real-world contexts, such as conveying sensory input or recalling information from memory.
-Examples of functional linguistic competence include implicatures (Ruis et al., 2022) and contextual language comprehension beyond the statistical manifestation of data distributions (Bransford & Johnson, 1972).
-Consequently, operating LLMs through a purely inference-based approach confines their capabilities within their provided context window, severely limiting their horizon.
-This results in deficiencies for situational modeling, non-adaptability through contextual changes, and short-term problem-solving, amongst other capabilities.
-However, simply increasing the context length may not yield greater capabilities, as demonstrated by the observed U-shaped performance curve (Liu et al., 2023) where LLMs excel when using information at the beginning or end of the input context, but struggle with information located in the middle, especially as context increases.
-These challenges are actively being researched, with novel approaches such as Hyena (Poli et al., 2023), RWKV (Bo, 2021), GateLoop (Katsch, 2023), and Mamba (Gu & Dao, 2023) surfacing.
-Meanwhile, the re-emergence of interest in retrieval-augmented generative approaches (Li et al., 2022a) offers an alternative by circumventing the autoregressive nature of the widely-used Transformer architecture (Vaswani et al., 2017), enabling context enrichment with lateral information.
-In parallel, efforts have focused on developing tool-based approaches (Schick et al., 2023) or template frameworks (Chase, 2023) to extend large LLMs’ capabilities and enable a broader spectrum of applications.
-However, these efforts only partially capture the vast potential inherent in leveraging LLMs as semantic parsers.
-In light of these considerations, we introduce SymbolicAI, a compositional neuro-symbolic (NeSy) framework able to represent and manipulate multi-modal and self-referential structures (Schmidhuber, 2007; Fernando et al., 2023).
-SymbolicAI augments the generative process of LLMs with functional zero- and few-shot learning operations and enables the creation of versatile applications through in-context learning (Wei et al., 2022a).
-These operations guide the generative process and facilitate a modular design with a wide range of existing solvers, including formal language engines for mathematical expression evaluation, theorem provers, knowledge bases, and search engines for information retrieval.
-It exposes these solvers as building blocks for constructing compositional functions as computational graphs, and facilitates the development of an extensible toolkit that bridges classical and differentiable programming paradigms, aiming to create domain-invariant problem solvers.
-In designing the architecture of SymbolicAI, we drew inspiration from the body of evidence that suggests the human brain possesses a selective language processing module (Macsweeney, 2002; Fedorenko et al., 2010; Menenti et al., 2011; Regev et al., 2013; Scott et al., 2016; Deniz et al., 2019; Hu et al., 2022), prior research on cognitive architectures (Newell & Simon, 1956; Newell et al., 1957; Newell & Simon, 1972; Newell, 1990; Laird, 2022), and the significance of language on the structure of semantic maps in the human brain (Huth et al., 2016).
-We consider language as a central processing module, distinct from other cognitive processes such as reasoning or memory (Paischer et al., 2022, 2023), that defines a stepping stone towards broad AI systems (see Section B).
-Lastly, alongside the framework, we introduce a benchmark and derive a quality measure and its empirical score to address the evaluation of multi-step NeSy generative processes.
-See Section 6 for more details.
-In summary, we highlight the following key contributions of this work:
+Consequently, by training on vast quantities of unlabelled textual data, LLMs have been shown to not only store factual knowledge (Petroni et al., 2019; Kassner et al., 2020) and approximate users’ intentions to some extent (Andreas, 2022), but also to unlock deep specialist capabilities through innovative prompting techniques (Nori et al., 2023).
+Despite their versatility, current LLMs face challenges such as fallacious reasoning and the generation of erroneous content, commonly referred to as hallucinations (Jones & Steinhardt, 2022).
+These limitations highlight the importance of integrating complementary symbolic methods to validate and guide the generative processes of LLMs, ensuring more accurate and reliable outputs.
+In parallel, efforts have focused on developing tool-based approaches (Schick et al., 2023) or template frameworks (Chase, 2023) to extend LLMs’ capabilities and enable a broader spectrum of applications.
+However, these efforts only partially capture the potential inherent in leveraging LLMs as semantic parsers.
+In contrast to parsers for structured languages a semantic parser is able to break down unstructured human language into semantically meaningful components and transform those into a structured form.
+While traditionally semantic parsing has been a role filled by specialized algorithms and models, we posit that LLMs, through their training on diverse linguistic data, have developed the ability to perform semantic parsing as part of their broader natural language processing capabilities.
+In turn, we identify LLMs as a central component in creating sophisticated neuro-symbolic (NeSy) AI systems.
+These systems integrate symbolic and sub-symbolic concepts and utilize the capabilities of semantic parsing to develop symbolic expressions that enable new probabilistic programming paradigms.
+We introduce SymbolicAI, a compositional NeSy framework able to represent and manipulate multi-modal and self-referential structures (Schmidhuber, 2007; Fernando et al., 2023).
+SymbolicAI augments the generative process of LLMs with in-context learning operations, realized through functional primitives, and enables the creation of versatile applications through in-context learning (Wei et al., 2022a).
+These operations enable logic-based components that guide the generative process and enable a modular NeSy system, including a wide range of existing solvers, formal language engines for mathematical expression evaluation, theorem provers, knowledge bases, and search engines for information retrieval.
+SymbolicAI exposes these solvers as building blocks for constructing compositional functions as computational graphs, making it possible to bridge classical and differentiable programming paradigms with the aim to create domain-invariant problem solvers.
+In designing the architecture of SymbolicAI, we drew inspiration from a body of evidence that suggests the human brain possesses a selective language processing module (Macsweeney, 2002; Fedorenko et al., 2010; Menenti et al., 2011; Regev et al., 2013; Scott et al., 2016; Deniz et al., 2019; Hu et al., 2022), prior research on cognitive architectures (Newell & Simon, 1956; Newell et al., 1957; Newell & Simon, 1972; Newell, 1990; Laird, 2022), and the significance of language on the structure of semantic maps in the human brain (Huth et al., 2016).
+We consider language as a central processing module, distinct from other cognitive processes such as reasoning or memory (Paischer et al., 2022; 2023).
+We hypothesize that such a central processing module based in language is a core component of broad AI systems (see Appendix Section A) and enables the development of fully autonomous AI systems for decision-making.
+A significant challenge encountered in the development of our framework pertained to the evaluation of LLMs when used as semantic parsers in a NeSy workflow.
+Current evaluation of generated content relies on metrics for single-step generative processes, such as the BLEU score (Papineni et al., 2002).
+These metrics are not suitable for evaluating multi-step generative processes.
+BLEU has limitations, as it measures $n$-gram-based overlap of generated output with a reference that does not consider the semantic meaning.
+As a result, BLEU fails to capture semantic equivalence, especially in more complex tasks.
+More recent metrics such as CIDEr (Vedantam et al., 2014) or SPICE (Anderson et al., 2016) are also not suitable for our problem, either because they are built upon BLEU (in case of CIDEr) or designed with inductive biases specifically for image captioning.
+Therefore, alongside our framework we introduce a quality measure (and its empirical score) for multi-step generative processes based on semantic meaning.
+We refer to our score as ”Vector Embedding for Relational Trajectory Evaluation through Cross-similarity”, or VERTEX score for short.
+Our VERTEX score uses embeddings to compare node distributions within a computational graph.
+It measures the semantic meaning across the distributional path by computing at each node the cross-similarity between the generated embeddings and embeddings sampled from a reference distribution.
+Furthermore, the VERTEX score is designed such that it can be used as a reward signal in a reinforcement learning setting (Sutton, 1984).
+Finally, we propose a benchmark for evaluating complex workflows.
+We define a set of basic evaluations, particularly associative predictions based on in-context learning, multi-modal bindings for tool utilization, and program synthesis for subroutine execution.
+Furthermore, we introduce complex evaluations for logic-based components and hierarchical computational graphs.
+In summary, the key contributions presented in this work are as follows:
   • We introduce SymbolicAI, a logic-based framework for concept learning and flow management in generative processes, enabling seamless integration with a wide range of foundation models and solvers.
-  • We combine the broad applicability of LLMs as semantic parsers with symbolic expressions by leveraging a modular probabilistic programming paradigm, facilitating the creation of complex computational graphs.
-  • We introduce a quality measure and its empirical score alongside a benchmark for comparing state-of-the-art LLMs across a wide range of tasks.
+  • We leverage LLMs as semantic parsers to enable the creation of complex computational graphs by combining symbolic expressions with probabilistic programming paradigms.
+  • We introduce a quality measure and its empirical score alongside a benchmark designed for multi-step generative processes for comparing LLMs across a wide range of complex tasks.
 2 Related Work
@@
 The field of symbolic AI has its foundations in the works of the Logic Theorist (LT) (Newell & Simon, 1956) and the General Problem Solver (GPS) (Newell et al., 1957).
-These programs represented the first steps towards automated reasoning and problem-solving using symbolic representations.
+These programs represented the first steps towards automated reasoning and problem-solving utilizing symbolic representations.
 Despite their advancements, both faced challenges in dealing with the complexity of real-world problems, particularly due to the combinatorial nature of the solution space.
 To address these limitations, the Soar (Laird et al., 1987) cognitive architecture was developed, advancing the notion that intelligent behavior results from goal-oriented search through a problem space (Newell & Simon, 1972; McCarthy et al., 2006), with each step consisting of selecting and applying operators.
-Soar introduced components like reinforcement learning, impasses, substates, and chunking to enhance its problem-solving capabilities.
+Soar introduced components like reinforcement learning, impasses, sub-states, and chunking to enhance its problem-solving capabilities.
 It also demonstrated the importance of learning from experiences to adapt and improve performance over time.
@@
 While in-context learning bypasses the need for explicit retraining, it demands meticulous prompt design to steer models towards desired behaviors.
-Despite their versatility, current LLMs face challenges such as fallacious reasoning and the generation of erroneous content, commonly referred to as hallucinations (Jones & Steinhardt, 2022).
-These limitations highlight the importance of integrating complementary symbolic methods to validate and guide the generative processes of LLMs, ensuring more accurate and reliable outputs.
 Neuro-Symbolic Methods
-To overcome the limitations of each individual method, NeSy approaches meld the statistical inference strengths of deep neural architectures with the generalization and explainability of symbolic systems (Besold et al., 2017; Yu et al., 2023; Hamilton et al., 2022; Garcez et al., 2015; d’Avila Garcez et al., 2019; d’Avila Garcez & Lamb, 2020; Lamb et al., 2020).
+To overcome the limitations of each individual method, NeSy approaches meld the statistical inference strengths of deep neural architectures with the generalization and explainability of symbolic systems (Garcez et al., 2015; Besold et al., 2017; d’Avila Garcez et al., 2019; d’Avila Garcez & Lamb, 2020; Lamb et al., 2020; Hamilton et al., 2022; Yu et al., 2023).
 Some approaches focus on different strategies for integrating learning and reasoning processes (Yu et al., 2023; Fang et al., 2024).
-Firstly, learning for reasoning methods treat the learning aspect as an accelerator for reasoning, in which deep neural networks are employed to reduce the search space for symbolic systems (Qu & Tang, 2019; Silver et al., 2016, 2017b, 2017a; Schrittwieser et al., 2020).
+Firstly, learning for reasoning methods treat the learning aspect as an accelerator for reasoning, in which deep neural networks are employed to reduce the search space for symbolic systems (Silver et al., 2016; 2017b; 2017a; Qu & Tang, 2019; Schrittwieser et al., 2020).
 Secondly, reasoning for learning views reasoning as a way to regularize learning, in which symbolic knowledge acts as a guiding constraint that oversees machine learning tasks (Hu et al., 2016; Xu et al., 2018).
@@
 This synergy further extends when considering graph-based methods, which closely align with the objectives of our proposed framework.
-Research in this area, such as CycleGT (Guo et al., 2020) and Paper2vec (Ganguly & Pudi, 2017), explored unsupervised techniques for bridging graph and text representations.
+Research in this area, such as CycleGT (Guo et al., 2020) and Paper2vec (Ganguly & Pudi, 2017) explores unsupervised techniques for bridging graph and text representations, GPTSwarm (Zhuge et al., 2024) explores graph optimizers to refine node-level prompts and edge optimization.
 Subsequently, graph embeddings, when utilized within symbolic frameworks, can enhance knowledge graph reasoning tasks (Zhang et al., 2021), or more generally, provide the bedrock for learning domain-invariant representations (Park et al., 2023).
@@
 While previous work has effectively identified opportunities and challenges, we have taken a more ambitious approach by developing a comprehensive framework from the ground up to facilitate a wide range of NeSy integrations.
-3 Problem Definition
-Conventional approaches employing foundation models for inference, such as LLMs, are predominantly confined to single-step or few-step executions and primarily reliant on hand-crafted in-context learning prompt instructions.
-This restricted scope limits the utilization to single modalities, lacks refinement or verification, and exhibits limited tool proficiency.
-We posit that the integration of NeSy engines as core computation units, realized through logic-based methodologies coupled with sub-symbolic foundation models, offers a more general, robust, and verifiable perspective.
-This approach has several advantages.
-Firstly, it facilitates the integration of pre-existing engineered solutions (e.g.
-various classical algorithms), offloading computational complexity and bridging various modalities.
-Secondly, it enables sub-symbolic generalization to focus on evidence-based decision-making (e.g.
-selecting the respective tool based on in-context classification).
-Thirdly, it provides an interpretable language-based control layer for explainable, autonomous systems.
-Central to our solution is a method to define and measure the orchestration of interactions between symbolic and sub-symbolic systems, and the level at which instructions are formulated for effective control and task execution.
-4 Design Principles
-In the following subsections, we elaborate on the key design principles underlying SymbolicAI and how we guide the generative processes of NeSy engines. For further technical details, see Section 5.
-Symbols and Expressions
-As posited by Newell & Simon (1976), symbols are elemental carriers of meaning within a computational context333 We base our framework’s name on the aspirational work of Newell and Simon..
-These symbols define physical patterns capable of composing complex structures, and are central to the design and interpretation of logic and knowledge representations (Augusto, 2022).
-Thus, SymbolicAI conceptualizes the notion that symbols, and the expressions they form, are reflections of the information inherent in a system, and serve as surrogate for the interaction between the system and the problem space.
-Moreover, we argue that real patterns, as Dennett (1991) speaks of, can be effectively realized through the use of symbols because these symbols act as versatile abstractions that capture and represent the underlying structures and dynamics of these patterns, facilitating their interpretation and manipulation in computational models.
-Furthermore, we attribute task-specific mappings to a language-centric strategy, leveraging their inherent semantics and abstraction to describe the states and properties of the problem at hand.
-These mappings are universal and may be used to define scene descriptions, long-horizon planning, acoustic properties, emotional states, physical conditions, etc.
-Therefore, we adhere to the analogy of language representing the convex hull of the knowledge of our society, utilizing it as a fundamental tool to define symbols.
-This approach allows us to map the complexities of the world onto language, where language itself serves as a comprehensive, yet abstract, framework encapsulating the diversity of these symbols and their meanings.
-This perspective resonates with our innate human tendencies to attribute existing physical objects with abstract concepts, as exemplified by our natural inclination to link tangible objects to colors and emotions, such as blending the color ”red” with ”heart”, ”warm”, and ”passion”.
-However, this language-centric model does not inherently encompass all forms of representation, such as sensory inputs and non-discrete elements, requiring the establishment of additional mappings to fully capture the breadth of the world.
-This limitation is manageable, since we care to engage in operations within this abstract conceptual space, and then define corresponding mappings back to the original problem space.
-These are typically applied through function approximation, as in typical modality-to-language and language-to-modality use cases, where modality is a placeholder for various skill sets such as text, image, video, audio, motion, etc.
-Ultimately, this approach also anchors our work in the field of formal language theory, as we require a structured method to construct mappings from the world to language.
-This grounding sets the foundation for employing formal language structures, such as grammars, to systematically define our language-centric approach to problem-solving and the associated translation of real-world complexities into linguistic terms.
-Formal Languages
-In formal language theory and linguistics, languages are structured following the Chomsky hierarchy, which classifies languages by the complexity of their grammatical structure (Chomsky, 1956).
-This hierarchy, comprising four types of grammars (Type-3 to Type-0), delineates formal languages by their grammatical complexity.
-A grammar in this context consists of terminal and non-terminal symbols, production rules, and a designated start symbol, facilitating the generation of valid strings within a language.
-In developing SymbolicAI, we posit that all symbols can be represented as strings, augmented with conditional instructions and types derived from a domain-specific language (DSL) tailored for directing NeSy computation engines, like LLMs (see Figure 2).
-A key advancement of LLMs over previous systems lies in their ability to generalize from formal languages (Wang et al., 2023a) and knowledge systems, primarily due to their world-knowledge and proficiency in understanding context-based analogies.
-While there is currently no universal consensus among experts regarding the precise classification of natural language within the Chomsky hierarchy, we have seen in our empirical evaluations promising results utilizing LLMs as semantic parsers.
-This approach can be viewed as employing a form of flexible, context-sensitive grammar, which enables the processing of instructions and analogies with a nuanced understanding of language’s inherent variability and complexity.
-The intersection between formal and natural languages becomes evident when considering how language patterns, through prompts like ”You are a helpful assistant…”, elicit structured responses, indicating a potential underlying formal mechanism at play.
-This observation underlines the utility of such a grammar in our framework, particularly within in-context learning, where it serves as an explicit schema guiding the structure of examples used in few-shot learning scenarios.
-For instance, equating ”3.1415…” with ”$\pi$” or ”August 4, 1961” with ”1961-08-04” in a given context demonstrates this probabilistic, context-dependent interpretation of symbols.
-Such a system doesn’t rigidly adhere to standard grammatical rules but instead adjusts and interprets based on the presented context, effectively creating a dynamic and situation-specific grammar.
+Large Language Models
+In part, instruction-based fine-tuning of LLMs through reinforcement learning from human feedback (Ouyang et al., 2022; Li et al., 2023) or direct preference optimization (Rafailov et al., 2023) has shown promising results dealing with value misalignment issues (Bradley Knox & Stone, 2008; MacGlashan et al., 2017; Christiano et al., 2017; Ibarz et al., 2018; Goyal et al., 2022), unlocking new possibilities for chain of thoughts (Wei et al., 2022b), tree of thoughts (Yao et al., 2023a), and graph of thoughts interactions (Besta et al., 2023).
+However, recent research also highlights the limitations of LLMs in functional linguistic competence despite their proficiency in formal linguistic competence (Mahowald et al., 2023).
+Whereas formal linguistic competence encompasses the ability to understand and generate language, functional linguistic competence pertains to the application of language in real-world contexts, such as conveying sensory input or recalling information from memory.
+Examples of functional linguistic competence include implicatures (Ruis et al., 2022) and contextual language comprehension beyond the statistical manifestation of data distributions (Bransford & Johnson, 1972; Mikolov et al., 2013b).
+Consequently, operating LLMs through a purely inference-based approach confines their capabilities within their provided context window, severely limiting their horizon.
+This results in deficiencies for situational modeling, non-adaptability through contextual changes, and short-term problem-solving, amongst other capabilities.
+However, simply increasing the context length may not yield greater capabilities, as demonstrated by the observed U-shaped performance curve (Liu et al., 2023) where LLMs excel when utilizing information at the beginning or end of the input context, but struggle with information located in the middle, especially as context increases.
+These challenges are actively being researched, with novel approaches such as Hyena (Poli et al., 2023), RWKV (Bo, 2021), GateLoop (Katsch, 2023), Mamba (Gu & Dao, 2023) and xLSTM (Beck et al., 2024) surfacing.
+Meanwhile, the re-emergence of interest in retrieval-augmented generative approaches (Li et al., 2022a) offers an alternative by circumventing the autoregressive nature of the widely-utilized Transformer architecture (Vaswani et al., 2017), enabling context enrichment with lateral information.
 In-Context Learning
@@
 Ensemble techniques further enhance the robustness and accuracy of model predictions by combining several strategies to establish a consensus (Nori et al., 2023).
-Conceptually, we designed our framework to enable all these techniques and combine them into dedicated components and sub-processes.
-Figure 2 outlines a typical expression evaluation in our NeSy pipeline.
-Figure 2: Illustration for NeSy pipeline, showcasing conceptual usage of in-context learning methodologies, domain-specific language (DSL) structures, and the expression evaluations through NeSy engines.
-The expression showcases the left shift operator $\ll$ and how the information of the symbol B is included in the symbol AC.
-The violet placeholder in the DSL Prompt represents an instruction, such as ”Insert the right-hand side value into the left-hand value in a chronological order.” The positions below represent task-specific few-shot examples.
+3 Problem Definition
+Conventional approaches employing foundation models, such as LLMs, are predominantly confined to single-step or few-step executions and primarily reliant on hand-crafted prompt instructions, often referred to as in-context learning.
+This restricted scope limits the utilization of different modalities, lacks verification, and exhibits limited tool proficiency.
+We posit that the use of NeSy engines as core computation units, realized through logic-based methodologies coupled with sub-symbolic foundation models, offers a more general, robust, and verifiable perspective.
+This approach has several advantages.
+Firstly, it enables the integration of pre-existing solutions (e.g.
+various classical algorithms), offloading computational complexity and bridging different modalities.
+Secondly, it allows sub-symbolic components to focus on decision-making (e.g.
+selecting the respective tool based on in-context classification).
+Thirdly, it provides an interpretable language-based control layer for explainable, autonomous systems.
+In the following section, we elaborate on the key design principles underlying SymbolicAI and how we guide the generative processes of NeSy engines.
+For further technical details, see Appendix Section 5.
+4 Design Principles
+Symbols and Expressions
+[…]
+Furthermore, we utilize language as a tool for mapping complex concepts, leveraging its inherent semantics and abstractions to describe states and properties of a problem at hand.
+These mappings are universal, e.g.
+they may be utilized to define scene descriptions, long-horizon planning, acoustic properties, emotional states, physical conditions, etc.
+Therefore, language serves as a comprehensive, yet abstract framework to encapsulate meanings, and refer to it as the convex hull of the knowledge of our society.
+Subsequently, it is common to attribute existing physical objects with abstract concepts, as exemplified by our natural tendency to link tangible objects to colors and emotions, such as blending the color ”red” with ”heart”, ”warm”, and ”passion”.
+This approach also anchors our work in the field of formal language theory, as we require a structured method to construct mappings from the world to language.
+Consequently, we use formal language structures, such as grammars, to systematically define our language-centric approach to problem-solving and the associated translation of real-world complexities into linguistic terms.
+Formal Languages
+In formal language theory and linguistics, languages are structured following the Chomsky hierarchy, which classifies languages by the complexity of their grammatical structure (Chomsky, 1956).
+This hierarchy defines four types of grammars (Type-3 to Type-0) and separates formal languages by their grammatical complexity.
+A grammar in this context consists of terminal and non-terminal symbols, production rules, and a designated start symbol, enabling the generation of valid strings within a language.
+We define a NeSy engine as a mapping $\mathcal{V}_{\mathcal{S}^{*}}:\Omega\times N\times\mathcal{T}\rightarrow\Sigma$, where $N\subset\Sigma$ is a set of non-terminal symbols, $\mathcal{T}\subset\Sigma$ is a set of terminal symbols and $N\cap\mathcal{T}=\emptyset$, and $\mathcal{S}^{*}\in\Sigma$ is a starting symbol.
+We further formalize a grammar $G=(N,\mathcal{T},P,\mathcal{S}^{*})$ with production rules defined as a $P:=\mathcal{V}_{\mathcal{S}^{*}}(\omega,N,\mathcal{T})$.
+This grammar describes the generation of symbols through expressions $\omega$.
+For simplicity, we will drop the subscript of $\mathcal{V}_{\mathcal{S}^{*}}$ and use it as $\mathcal{V}$.
+Figure 2: Illustration for NeSy pipeline, showcasing conceptual usage of in-context learning methodologies, domain-specific language (DSL) structures, and the expression evaluations through a NeSy engine based on an LLM and constraint verification.
+The expression showcases the sorted insert operator $\ll$ and how the information of the symbol B is included in the symbol AC.
+The violet placeholder in the DSL Prompt represents an instruction, such as ”Insert the right-hand side value into the left-hand value in ascending order.” The positions below represent task-specific few-shot examples.
+The DSL Prompt receives the expression $\omega_{<<}$ and maps it to $\hat{\omega}_{<<}$ that can be processed by the LLM-based NeSy function $\mathcal{V}_{\mathcal{S}^{*}}$ and outputs a new symbol.
+We identify LLMs as promising candidates for functioning as part of NeSy engines.
+In SymbolicAI, a symbol $\mathcal{S}$ is augmented with conditional instructions and types derived from DSLs, custom defined or not (e.g.
+HTML, SQL, etc.), tailored for directing the LLMs.
+The key advantage of LLMs over previous systems lies in their ability to generalize across formal languages (Wang et al., 2023a) and knowledge systems.
+Although there is currently no universal consensus regarding the precise classification of natural language within the Chomsky hierarchy, our approach can be understood as employing a situation-specific, context-sensitive grammar, which enables the processing of instructions and analogies with a nuanced understanding of language.
+The intersection between formal and natural languages becomes evident when considering how language patterns, through prompts like ”You are a helpful assistant…”, elicit structured responses, indicating a potential underlying formal mechanism at play.
+This observation underlines the utility of such a grammar in our framework, where it serves as an explicit schema guiding the structure of examples for in-context learning.
+For instance, equating ”3.1415…” with ”$\pi$” or ”August 4, 1961” with ”1961-08-04” in a given context demonstrates context-dependent interpretation of symbols.
+Such a system doesn’t rigidly adhere to standard grammatical rules but instead adjusts and interprets based on the context, effectively creating a situation-specific grammar, capable of forming Domain-Invariant Associations through in-context learning.
+We further address this in a later paragraph.
+Function Composition
+In SymbolicAI, we use function composition to construct complex hierarchies and behaviors from fundamental elements.
+Therefore, our framework enables modeling of interconnected processes, where the output of one function is used as input for another, thus creating a sequence of operations.
+Through function composition, we construct computational graphs, in which intermediate symbols represent the nodes or states within these graphs.
+Formally, function composition is denoted by $\circ$, where combining functions $f$ and $g$ yields a new function $h=g\circ f$, defined as $h(x)=g(f(x))$ For functions $f:X\rightarrow Y$ and $g:Y\rightarrow Z$, their composition results in a function mapping elements from domain $X$ to codomain $Z$ through $g(f(x))$.
+Although traditionally the codomain of the inner function $f$ aligns with the domain of the outer function $g$, SymbolicAI relaxes this constraint by allowing for any subset relationship between these domains and codomains, which is particularly beneficial for in-context learning.
+When using LLMs for NeSy production rules $\mathcal{V}$, we can derive a multi-step generative process by composing a computational graph as a sequence of zero- and few-shot function compositions:
+$\mathcal{V}(\omega_{j},N,\mathcal{T})=\mathcal{V}(\omega_{j-1},\cdot)\circ\mathcal{V}(\omega_{j-2},\cdot)\circ\dots\circ\mathcal{V}(\omega_{0},\cdot),$ (1)
+where $\omega_{0}$ is the initial instruction and $j$ defines the index variable for a multi-step generative process.
+By leveraging functional in-context learning, where zero- and few-shot examples act as dynamic elements of the function’s domain, SymbolicAI has the ability to interpret and respond to diverse input contexts.
+For instance, a function can classify a user request and select an appropriate interface (e.g.
+WolframAlpha) to process the request.
+The output modality may even vary based on the respective engine.
+This enables SymbolicAI to handle operations over multi-modal data that connects multi-step generative processes and establishes function composition as a central tenet in bridging multiple modalities and coordinating a variety of tasks.
 Domain-Invariant Associations
@@
 We attribute this to associations formed within the input space and the capacity of Transformer architectures for defining domain-invariant feature sub-spaces.
-This phenomenon has striking parallels with advancements in few-shot learning approaches such as SubGD (Gauch et al., 2022), a method based on identifying and utilizing a low-dimensional subspace, learned from various tasks, that effectively acts to regularize the learning process, representing features that are invariant across different learning tasks.
-Furthermore, SubGD reflects the potential of in-context learning when combined with task-specific fine-tuning by showing that fine-tuning within a learned subspace significantly outperforms traditional fine-tuning methods.
-We believe that the extent of in-context learning is not yet exhausted, holding considerable promise when used alongside with task-specific fine-tuning and solvers.
-To develop learning and reasoning systems capable of general problem-solving, we adopt a hybrid methodology.
-This approach leverages the in-context generalization capability of LLMs, constructing symbolic associations that aim to preserve and propagate situational context, and validating solutions with established solvers.
-Function Composition
-In SymbolicAI, function composition is relevant for constructing complex hierarchies and behaviors from more basic, fundamental elements.
-It enables our framework to model interconnected processes, where the output of one function seamlessly transitions into the input of another, thus creating a cohesive sequence of operations.
-Through function composition, we construct computational graphs, in which intermediate symbols represent the nodes or states within these graphs.
-Formally, function composition is denoted by $\circ$, where combining functions $f$ and $g$ yields a new function $h=g\circ f$, defined as $h(x)=g(f(x))$ For functions $f:X\rightarrow Y$ and $g:Y\rightarrow Z$, their composition results in a function mapping elements from domain $X$ to codomain $Z$ through $g(f(x))$.
-Although traditionally the codomain of the inner function $f$ aligns with the domain of the outer function $g$, SymbolicAI relaxes this constraint by allowing for any subset relationship between these domains and codomains, enhancing data flow flexibility.
-For example, this relaxed constraint in domain and codomain alignment is particularly beneficial for in-context learning.
-By leveraging functional few-shot learning, where few-shot examples act as dynamic elements of the function’s domain, SymbolicAI enhances its ability to interpret and respond to diverse input contexts.
-For instance, a function can classify a user request and select an appropriate engine to process the request.
-The target modality may vary based on the respective engine.
-Therefore, in SymbolicAI, the use of function composition leads to the creation of richer and more nuanced symbol-based generative flows, where each functional unit is capable of representing either logical or data-driven behaviors.
-Importantly, function composition is not confined to strictly symbolic representations; it also connects with sub-symbolic processes.
-This enables SymbolicAI to handle complex data streams, establishing function composition as a central tenet in bridging multiple modalities and coordinating a variety of tasks.
-5 Framework
-In this section, we discuss the design choices pertaining to implementation and the specifics of the framework we’ve employed. For an extended overview, see our supplementary materials Section C.
+This phenomenon has parallels with few-shot learning approaches such as SubGD (Gauch et al., 2022), a method based on identifying and utilizing a low-dimensional subspace, learned from various tasks that effectively regularize the learning process.
+Since LLMs have been trained on different domains and tasks, which also include formulations of mathematical expressions, we posit that specific tokens, such as the equality sign, can be leveraged to associate meanings between different symbolic objects.
+Unlike domain-invariant representations that create invariant features across different learning tasks, our approach leverages the in-context generalization capability of LLMs to construct invariant symbolic associations that aim to preserve, manipulate and propagate situational context.
+We can use these properties to build operations that apply transformations on objects that are substitutes to the semantically aligned few-shot learning examples.
+5 SymbolicAI Framework
+In this section, we discuss the specifics of the proposed SymbolicAI framework.
+For more details about the framework structure, see Appendix Section C.
+For installation and usage of our framework, see Appendix Section D.
+For more technical details and code snippets, see Appendix Section E.
 Types and Representations
-Analogous to the Python object type, the base type of SymbolicAI is a symbol represented through its name equivalent base type Symbol.
-A Symbol object marks a non-reducible atomic unit.
-All other subtypes, such as Expression and its derivatives, are analogous to their mathematical namesakes, representing expressions or units that can be further evaluated and simplified.
+Analogous to the type object in Python, the base type of SymbolicAI is a symbol represented by the base type Symbol.
+All other subtypes, such as Expression, represent their mathematical namesake and can be evaluated and simplified.
 These subtypes inherit from Symbol the base attributes, primitive operators, and helper methods.
-Furthermore, each Symbol object contains valued and vector-valued representations, obtained through value and embedding attributes.
-The latter, in particular, serve as a means to attribute a symbol’s current context, akin to embedding text and storing it as a PyTorch tensor (Paszke et al., 2019) or NumPy array (Harris et al., 2020).
-While for an LLM, the numerical tensors may lack inherent meaning, vector-valued representations play a strategic role when 1) composite symbols coalesce into more complex expressions, and 2) these embedded tensors become amenable to updates through gradient-based optimization.
-This design choice is critical in the dynamic landscape of symbolic interactions, particularly for promising applications, such as developing automated and self-evolving NeSy systems.
-To enable the processing of symbols by LLMs, we assume that each Symbol object is represented through Python’s native string functionality, where the __str__ method acts an interpretable string representation.
-Worth noting is that encoding a complex object into a string sometimes precludes the object reconstitution.
-However, this concern does not substantially impede our methodology: we can employ approximations or proxy representations stored by the vector-valued property to effectively re-map objects.
-These representations are obtained through respective embedding models.
-Therefore, we can theoretically assert that any Python object is parsable by an LLM by design.
-For more details, see our supplementary materials Section E.
+Although SymbolicAI uses a language-centric design, modeling and manipulating every interaction into symbolic representations is not inherently efficient.
+Therefore, we establish mappings between symbolic and sub-symbolic representations for sensory inputs and non-discrete elements.
+Such mappings are typically realized through function approximation.
+This allows us to map between modality-to-language and language-to-modality use cases.
+Here, modality serves as a placeholder for various types such as text, image, video, audio, motion, etc.
+In turn, each Symbol object contains valued and vector-valued representations, obtained through value and embedding attributes.
+The latter represents a symbol’s current value, akin to embedding text and storing it as a PyTorch tensor (Paszke et al., 2019) or NumPy array (Harris et al., 2020).
+While for an LLM, the numerical tensors may lack inherent meaning, vector-valued representations play an important role when 1) composite symbols are combined into more complex expressions, and 2) these embedded tensors are updated through gradient-based optimization.
+To enable the processing of symbols by LLMs, we assume that each Symbol object implements Python’s native string functionality, where the __str__ method returns an interpretable string representation.
+Therefore, we can assert that any Python object is parsable by an LLM, however, the user must ensure a meaningful representation.
+For more details, see Appendix Section E.
 Polymorphic Context
-Polymorphism is a central concept in programming language theory and prominently featured in SymbolicAI, with significant implications for the design and functionality of our NeSy architecture.
+Polymorphism is a central concept in programming language theory and prominently featured in SymbolicAI.
 Polymorphism refers to the ability of different objects to be accessed through the same interface, or of a single identifier to represent different types based on the context of execution.
-The provision of a single interface to entities of different types allows operations to be performed in ways specific to their derived types.
-We employ a polymorphic structure for instruction composition.
-We designed the Symbol object to contain a global context, which is composed of static and dynamic context parts.
+Providing a single interface for entities of different types allows operations to be performed in ways specific to their derived types.
+We therefore designed the Symbol object to contain a global context, which is composed of static and dynamic context parts, and enables this polymorphic behavior.
 The static context is class dependent and defined at design time.
 The dynamic context is runtime adaptable and can be changed to adhere to runtime specific logic and changes.
-Moreover, Symbol associated operations resolve in a polymorphic manner before being evaluated by the NeSy engines.
+Moreover, Symbol associated operations are resolved following polymorphic design before being evaluated by the NeSy engine.
 SymbolicAI’s engine implementation contains a prepare method to resolve and compile the engine specific representation by evaluating the Symbol-specific operations and context.
-For example, when utilizing GPT-4 vision as a NeSy engine backend, we compose the system and user level prompts in the prepare statement and resolve image or video related URL queries by parsing meta tags such as the <<vision: ...
-:>> tag using regular expressions.
+For an example on polymorphic context see part a) in Figure 3.
 Operators and Methods
 In SymbolicAI, operators are overloaded to facilitate transformations of Symbol objects.
-These operator primitives employ dynamic casting to assure type compatibility, simplifying declarations.
+These operator primitives employ dynamic casting to ensure type compatibility.
 Consequently, Symbol objects can be easily manipulated through type specific attributions or symbolically evaluated by the NeSy engine.
@@
 If the evaluation was unsuccessful, we then consider semantic equality through the NeSy engine.
-SymbolicAI leverages decorators for compose operators and custom class methods.
-For more details, see supplementary materials Section C.
-Upon invoking an operator or method, the respective primitive function evaluates the symbol’s specific type and its respective attributes, and if further necessary, resolves a nested decorated function that then references the NeSy engine for an evaluation.
+SymbolicAI leverages decorators for composing operators and custom class methods.
+For more details, see Appendix Section C.
+Upon invoking an operator or method, the respective primitive function evaluates the symbol’s specific type and its respective attributes, and if necessary, resolves a nested decorated function that then uses the NeSy engine for evaluation.
 Should the evaluation fail, a predefined fallback implementation executes.
@@
 JSON formatting through grammar-based verification).
-In Figure 3 we give an overview of the entire prompt composition based on the user input, the Symbol object structure, and the NeSy engine evaluation pipeline.
-Figure 3: We illustrate our NeSy engine pipeline and how we map user input and object instance attributes to a prompt instruction.
-Yellow marks user input, blue marks templating and transformations, red marks constraints, and violet the model prediction placeholder.
-The user input and Symbol instance attribute values are evaluated and transformed according to the pre- and post-processing steps.
-A custom method is constructed by using our reference implementation of decorators and passes the user input and instance attributes to the respective engine.
-A pre-processor manipulates the string representation of user input and Symbol instance attributes before forwarding them to the engine.
-For instance, in case the NeSy engine is an LLM (i.e.
-GPT-4), the static and dynamic context of a Symbol instance are mapped to the system and user prompts for that respective LLM using the prepare method of the engine.
-Other attributes, such as examples or templates, can help pre-define structures for the desired LLM output format.
-After the result is computed by the engine, it is passed to the post-processors and validated against the constraints.
+In Figure 3 b) we give an overview of the entire prompt composition based on the user input, the Symbol object structure, and in part c) the Symbol evaluation pipeline.
+Figure 3: a) Illustration of polymorphic context on the example of a SQLExpression type for the add-operator.
+Without a polymorphic context a regular Expression evaluation concatenates two Symbol objects together.
+The polymorphic context in SQLExpression overwrites the base behavior such that two added SQL-expressions get semantically combined, not concatenated.
+b) Illustration of the translation of a Symbol object to a prompt statement to be processed by an LLM in the NeSy engine.
+The User Input Args can be attached with a Payload from previous executions and gets applied to the Custom Method.
+The user input with the polymorphic context of the Symbol Object attributes (Static Context and Dynamic Context) are translated to a prompt statement according to the schema of the Prompt Design.
+The fields Operation, Examples and Template mark operation description, DSL-based prompt examples and template structures respectively.
+These translations are processed according to PreProcessor and engine-specific formatting.
+c) Illustrates the evaluation pipeline from user input to output, with multiple translation processes before and after the Engine invocation.
+The Input gets passed to the Custom Method and reformatted according to a PreProcessor to adhere to DSL-specific structure.
+The engine then takes the output of the PreProcessor and composes the final prompt according to the engine-specific Prompt Design and resolves polymorphic context and auxiliary fields.
+The output of the Engine then can be restructured by a PostProcessor to match DSL-requirements of the desired Output and gets applied Constraints to verify the outcome.
 Self-Referential Structures
@@
 We derive subtypes from Expression and enclose their functionalities in task-specific components, which we then expose again through templating and the model-driven design of the NeSy engine.
-This design choice allows a system to create and utilize its own sub-process definitions, analogous to concepts discussed in Schmidhuber (2007, 2009).
-Concretely, we utilize generalization properties from LLMs to interpret and formulate a set of operations that incorporate self-instructions (Wang et al., 2022).
+This design choice allows a system to create and utilize its own sub-process definitions, analogous to concepts discussed in Schmidhuber (2007; 2009).
+Concretely, we utilize generalization properties of LLMs to interpret and formulate a set of operations that incorporate self-instructions (Wang et al., 2022).
 Consequently, the operations hold the flexibility to adapt to the context, and derive sub-processes that self-instruct LLMs to engage in situational modeling and context-sensitive problem-solving.
 Ultimately, this enables the construction of hierarchical computational graphs for self-referential meta-reasoning systems without the need to explicitly training a meta-learner (Kirsch & Schmidhuber, 2022).
-6 Evaluation
-In this section, we conclude our paper with an empirical evaluation and potential areas of application using state-of-the-art LLMs as NeSy engine backends.
-For our evaluation we focus on the GPT family of models GPT-3.5 Turbo (revision 1106) and GPT-4 Turbo (revision 1106) (Brown et al., 2020) as they are the most proficient models to this date; Gemini-Pro as the best performing model available through API from Google; LlaMA 2 13B as it defines a good reference implementation for available open-source LLMs from Meta; Mistral 7B and Zephyr 7B as good baselines for revised and fine-tuned open-source contestants respectively.
-The selected open-source models Mistral, Zephyr, and LlaMA 2 are expected to have roughly equivalent parameter counts compared to GPT-3.5 Turbo and Gemini-Pro.
-All our experiments are expected to require a context size smaller or equal to $4096$ to enable the comparisons among the in-context capabilities across model architectures.
-For LlaMA 2 we use the chat version since it better follows instructions.
-Our evaluation focuses on three main parts.
-First, we evaluate basic capabilities of the models to verify their consistency and proficiency with in-context learning tasks.
-Second, we evaluate the logic capability of models by defining logical expressions in mathematical and natural language form and measure the capability of the models to translate and evaluate logical statements across domains.
-Lastly, we evaluate the capability of models in planning, constructing, maintaining, and executing hierarchical computational graphs.
-The models are instructed to extract an execution plan, which they then must schedule and execute.
-A significant challenge is the evaluation of the semantics of a multi-step generative process, since the naïve assessment of task succession would score all models to zero and render them as unusable.
-Therefore, we derive a quality measure to address these issues.
-Performance Measure
-One of the challenges in model evaluation is handling irrelevant predictions from models.
+In Figure 4 we illustrate a step-wise evaluation of a contextual computational graph, in which the NeSy engine is processing conditioned on the current execution context and producing a next symbol prediction.
+Figure 4: We showcase a multi-step hierarchical computational graph, with each node in the graph represented by a symbol.
+The edges are relations between symbols.
+The left-hand side illustrates how a new node (Symbol 3) is obtained by evaluating an operation with its respective context on a NeSy engine.
+The right-hand side illustrates the context information window (yellow rectangle) and relationship of the resulting graph with its respective nodes.
+6 Performance Measure
+One of the challenges when creating multi-step generative processes with LLMs as part of NeSy engines relies on model evaluation and handling irrelevant predictions.
+The naïve assessment that measures only task succession would score all models to zero and render them as unusable.
 Even if models follow instructions and produce parts of the expected solution, we regularly observe that they — especially open-source models — append a continuation of task irrelevant predictions.
@@
 We therefore adopt a cumulative measure capable of taking into account the entire generative trajectory.
-In theory, this process would entail calculating the path integral over the latent space representations for models, cumulating the Fréchet distances (Dowson & Landau, 1982) traversed along these distributional trajectories:
+In theory, this process would entail calculating the path integral over the latent space representations for models, cumulating the Fréchet distances (Dowson & Landau, 1982) traversed along these trajectories:
 […]
@@
 However, this approach is computationally intractable for large-scale problems, and requires access to latent representations, which — especially in the context of LLMs — is not always given.
-For computational feasibility, we therefore introduce an approximation that measures the embedding distances over the path trajectories utilizing an auxiliary embedding model, based on prior work on distribution regression (Szabó et al., 2016).
+For computational feasibility, we introduce an approximation that measures the embedding distances over the path trajectories through an auxiliary embedding model, based on prior work on distribution regression (Szabó et al., 2016).
 The embedding model maps the symbolic representations into a RKHS, such that we can apply a kernel mean embedding function to measure their respective distances (You et al., 2019; Dinu et al., 2023).
@@
 We therefore analyze the properties of the MMD and derive a similarity score, which follows the same statistical principles as the MMD, and is bound between $[0,1]$.
-We concluded that we can use only the MMD cross terms to evaluate the similarities.
-See our supplementary materials Section A for more details.
+We concluded that we can utilize only the MMD cross terms to evaluate the similarities.
+See Appendix Section B for more details.
 For our comparisons as referenced in Figure 6 we therefore denote the similarities rather than distances.
@@
 to scores obtained from comparing related solutions $z$.
-The use of the min-max scaling ensures the final measure is bounded between $[0,1]$.
-This process reflects properties such as Hölder continuity that bounds the kernel function within certain limits, ensuring that the similarity measure is appropriately scaled.
+Min-max scaling ensures the final measure is bounded between $[0,1]$.
+This process reflects properties such as Hölder continuity that bounds the kernel function within certain limits.
 To compute the embeddings, we utilize the embedding model all-mpnet-base-v2 (Song et al., 2020), due to its widespread availability, and its balance between speed and quality.
-As a similarity measure, we select a Gaussian kernel following our derivation from the supplementary materials Section A.
-In our supplementary implementations, we also explore other kernels, including some preliminary experiments with cosine similarity.
+As a similarity measure, we select a Gaussian kernel following our derivation from the Appendix Section B.
+In our implementations, we also explore other kernels, including preliminary experiments with cosine similarity.
 We also note that one can integrate Bernoulli distributed trials into our score, with 00 values representing failure modes and values of $1$ being successes.
@@
 including a similarity measure tailored towards capturing the nuances between two sub-structures of abstract syntax tree.
-In our evaluation of Figure 5 we conclude with the cumulative score, for the following base performance criteria.
-Basic Capabilities
-We analyze the consistency and stability of models and establish baselines by addressing the following performance criteria:
-  • Associative Prediction: We evaluate the success rate of models to follow simple and complex instructions and associations with zero- and few-shot examples.
-We therefore address the proficient use of our operators between Symbol types.
-  • Multi-modal Binding: We perform data transformations between multiple modalities by binding through language-based representations, and evaluate their proficiency in tool utilization, classification and routing of requests to relevant modules.
-  • Program Synthesis: We evaluate executable code with and without including concepts from retrieval augmented generation, model-driven development, such as templating to direct the generative flow, and experiment with self-generated instructions by creating self-referential expressions.
-We not only address similarity between produced solutions, but also include the measurement of abstract syntax tree equivalence and the success of executions.
+7 Evaluation
+We introduce a benchmark that evaluates multi-step generative processes as complex workflows.
+Our benchmark consists of five different evaluation categories, and uses the VERTEX score to measure the capabilities of an LLM to solve tasks from each category.
+The five categories of our benchmark are divided into three basic evaluations and two advanced categories that combine different basic capabilities.
+The three basic categories are (i) associative prediction which measures a models proficiency in understanding associations between symbols, (ii) multi-modal binding where we evaluate the capability to employ tools and operate on different modalities, and (iii) program synthesis for measuring a models proficiency in generating consistent code and executing subroutines.
+The two advanced benchmark categories are (iv) logic, for evaluating logic-based components and (v) computational graphs where complex workflows need to be processed, evaluating all aforementioned capabilities.
+For our evaluation we focus on the GPT family (Brown et al., 2020) of models, specifically GPT-3.5 Turbo (revision 1106) and GPT-4 Turbo (revision 1106) as they are the most proficient models to date; Gemini-Pro (Google, 2023) as the best performing model available through API from Google; LLaMA2-Chat 13B (Touvron et al., 2023), LLaMA3-Chat 8B and LLaMA3-Chat 70B from Meta represent open-source LLMs.
+Finally, Mistral 7B (Jiang et al., 2023) and Zephyr 7B (Tunstall et al., 2023) serve as baselines for revised and fine-tuned open-source models.
+The open-source models Mistral, Zephyr, and smaller LLaMA variants are estimated to have roughly equivalent parameter counts compared to GPT-3.5 Turbo and Gemini-Pro.
+All our experiments require a context size smaller or equal to $4096$ to enable the comparisons among the in-context capabilities across model architectures.
+For the LLaMA models, we utilize the chat versions since they are specifically tuned to follow instructions.
+Associative Prediction
+We evaluate a model’s proficiency to follow simple and complex instructions and associations with zero- and few-shot examples.
+Therefore, we evaluate the proficiency in applying our operators between Symbol types.
+We defined a total of 15 tasks involving in-context associations between two Symbol instances.
+SymbolicAI’s overloaded operators rely on predefined pseudo-grammars, as described in Section 4, that augment the operators with few-shot examples.
+For instance, the overloaded operator + utilized between two Symbol instances provides few-shot examples how to resolve additions with various data types.
+Consequently, we can now test if the models can solve the addition between Symbol("two hundred and thirty four") and Symbol(7000).
+See Appendix Section F.1 for more details.
+Multi-modal Binding
+We perform transformations between multiple modalities through language-based representations.
+Therefore, we need to evaluate the model’s proficiency in tool utilization, classification and routing of requests to relevant modules.
+We define a multi-modal Expression to detect the category of a task based on its content and to forward the task to the appropriate tool.
+The expression creates interfaces to tools like WolframAlpha for mathematical expressions, Selenium for website content scraping, SerpApi for search queries, and APILayer for optical character recognition.
+Each of the five tests aims to evaluate the appropriate handling of a specific type of input by the multi-modal Expression type, such as processing a website URL for scraping, interpreting a search engine query, testing if two vectors are linearly independent, comparing large numbers, and extracting text from an image.
+See Appendix Section F.2 for more details.
+Program Synthesis
+We evaluate executable code with and without concepts from retrieval augmented generation, model-driven development, and experiment with self-generated instructions by creating self-referential expressions.
+We designed three separate tests related to program synthesis, where each task assesses the ability of the models to generate and execute code based on natural language instructions or provided templates:
+1) The first task involves reading a LaTeX table template and data, then generating a function to populate the table with the given data.
+2) The second task tests the automatic code generation for API calls by fetching data from a specified URL and extracting specific information from the retrieved content.
+3) The third task evaluates the ability to construct a custom Expression that processes a Symbol through a specific Function component from the SymbolicAI package.
+Each of the three tests follows a similar pattern, where the generated code is scored based on its similarity to valid references and normalized with random samples.
+See Appendix Section F.3 for more details.
 Logical Components
-To evaluation the logic capabilities of models, we conditioned them to create a sequence of expressions as self-contained components, and refer to higher-order predicate logic for their assessment.
-Based on the underlying type theory originating from Whitehead & Russell (1925–1927), we evaluate a model’s capability in the form of there exists x such that x fulfills y.
-Such quantifiers define the standard semantics over expressions, where the meaning is given by a semantic function.
+To evaluation the capabilities for logical reasoning of models, we condition them to create a sequence of expressions as self-contained components, and refer to higher-order logic for their assessment.
+Based on the underlying type theory originating from Whitehead & Russell (1925–1927), we evaluate a models’ capability to resolve statements in the form of there exists x such that x fulfills y.
+Such quantifiers define the standard semantics of expressions, where their meaning is given by a semantic function.
 A semantic function maps a term from an abstract definition to a point in a domain, which is an interpretation of the term’s type and value.
 Therefore, these functions operate on types and values of expressions, and relations thereof.
-Subsequently, NeSy engines can formulate and evaluate at inference time logic-based instructions through Lisp, Prolog, or Mathematica (McCarthy, 1959; Colmerauer & Roussel, 1993; Chen et al., 1993; Inc., 2022), or leverage SAT solvers such as Z3 (Moura & Bjørner, 2008).
-Therefore, the evaluation of a natural language statement can be interpreted by any expert system which defines the corresponding semantic functions and process them either in a symbolic (Feigenbaum et al., 1965; Gamble et al., 1994), differentiable (Veličković & Blundell, 2021; Ibarz et al., 2022), or hybrid manner (Kuncicky et al., 1991).
-We evaluate how proficient models interpret custom DSLs and define expression statements.
-The DSL is designed to express logical relations and operations in a structured format, and supports human-readable and machine-interpretable formulations.
+Subsequently, NeSy engines can formulate and evaluate at inference time logic-based instructions through Lisp, Prolog, or Mathematica (McCarthy, 1959; Colmerauer & Roussel, 1993; Chen et al., 1993; Inc., 2022), or leverage solvers such as Z3 (Moura & Bjørner, 2008).
+Therefore, the result of a natural language statement when evaluated by a NeSy engine can be interpreted by any expert system which defines the corresponding semantic functions and process them either in a symbolic (Feigenbaum et al., 1965; Gamble et al., 1994), differentiable (Veličković & Blundell, 2021; Ibarz et al., 2022), or hybrid manner (Kuncicky et al., 1991).
+We evaluate how proficient models are at interpreting custom DSLs and define expression statements.
+DSLs are designed to express logical relations and operations in a structured format, and supports human-readable and machine-interpretable formulations.
 The following example illustrates such relationships by translating a natural language statement into an expression statement, as follows:
 Marvins has four paws and likes to meow when I pet its fur. Is Marvins a cat?
-A DSL may enforce the usage of $\text{HAS}(\cdot)$, $\text{IS}(\cdot)$, etc. and may condition a LLM to produce the following expressions:
+A DSL may enforce the usage of $\text{HAS}(\cdot)$, $\text{IS}(\cdot)$, etc. and may condition an LLM to produce the following expressions:
   • $\text{HasFourPaws}(x)$: $x$ has four paws.
@@
   • $\text{IsCat}(x)$: $x$ is a cat.
-These are then used to define the following logical expression:
+These are then utilized to define the following logical expression:
 $\forall x\big{(}\text{HasFourPaws}(x)\land\text{LikesToMeowWhenPetted}(x)\Rightarrow\text{IsCat}(x)\big{)}.$
-An automated theorem prover can now evaluate this statement for all $x$ values and assess the validity of the original query.
+An automated theorem prover can now evaluate this statement for all values of $x$ and assess the validity of the original query.
 Lastly, our evaluation uses symbolic mathematics to manipulate algebraic expressions.
@@
 The symbols are placeholders for any value, enabling the definition of general expressions without specifying their values upfront.
+We designed six tests to assess the logical capabilities of the candidate models and group them as follows. See Appendix Section F.4 for more details.
+1) We utilize the Python library SymPy for symbolic mathematics to create the mathematical expression $ax+bx-cx-ay-by+cy+d$.
+The task for the model is then to factorize the expression and extract all unique symbols as a list.
+2) Three tasks evaluate a models’ capability to resolve the logical operations AND, OR, and XOR.
+For instance, the test for logical AND combines the symbols Symbol("The horn only sounds on Sundays") and Symbol("I hear the horn") and compares the answer against the human-generated references ”The horn only sounds on Sundays and I hear the horn.” and ”Since I hear the horn it is Sunday.” Since there is a large number of possible solutions, there is high variability in the solution space.
+Each model might prefer a different solution.
+3) For another task we use a custom Expression that defines a DSL syntax and semantic structure.
+We use this Expression to extract higher-order logic expressions from a natural language statement, namely the puzzle ’Who is Jay’s brother?’444 Bob has two sons, John and Jay.
+Jay has one brother and father.
+The father has two sons.
+Jay’s brother has a brother and a father.
+Who is Jay’s brother?, that preserves the original relationships.
+4) For the final task, we again use the puzzle ’Who is Jay’s brother?’ to evaluate a models’ capability for complex conversions.
+We use the Z3 theorem prover (Moura & Bjørner, 2008) to solve the ’Who is Jay’s brother’ puzzle conditioned on the Z3 solvers’ solution to Einsteins’ famous puzzle ’Who owns the fish?’.
+The task involves an indirect translation from natural language to executable code by the Z3 solver; the solution to Einstein’s puzzle acts as a form of self-contained ”documentation” for how the Z3 solver should be utilized.
+The test constructs a template, which includes the task instructions, puzzle statement, and reference to the Einstein’s puzzle solution.
+The models are then asked to analyze the given problem and solution format and create a Python function with Z3 syntax that can solve the ’Who is Jay’s brother?’ puzzle.
+The dynamically generated code is executed within the test environment utilizing Python’s exec function.
+We check the access to the Z3 solver and run the generated solve_puzzle function supposed to contain the logic to solve the puzzle.
+Once executed, the assembled Z3 logical clauses are processed by the solver, which verifies that the set of constraints is satisfiable.
+If so, the model generated by the solver is queried for the puzzle’s solution and scored using our VERTEX score.
 Hierarchical Computational Graphs
 We evaluate the capabilities of models to orchestrate a multi-step generative process and evaluate a set of tasks.
-Models need to direct sub-processes and associate computational results from and to Symbol nodes, and maintain relationships between these nodes.
-Given that the field is currently at an early stage in developing even sequential schedulers for LLM-based planning systems, our initial evaluations will be confined to sequential execution only.
-In Algorithm 1, we detail our evaluation protocol, which is designed not only to analyze and score a series of instructions but also to provide a structured basis for recording these processes.
-We also note that our evaluation protocol is generally formulated, which allows the application of non-sequential plan schedulers.
-In Figure 4 we illustrate a step-wise evaluation of the contextual computational graph, in which the NeSy engine is processing conditioned on the current execution context and producing the next symbol prediction.
-Figure 4: We showcase a multi-step hierarchical computational graph, with each node in the graph represented by a symbol.
-The edges are relations between symbols.
-The left-hand side illustrates how a new node (Symbol 3) is obtained by evaluating an operation with its respective context on a NeSy engine.
-The right-hand side illustrates the context information window (yellow rectangle) and relationship of the resulting graph with its respective nodes.
+Models need to direct sub-processes and associate computational results from and to Symbol nodes, and maintain relationships between these nodes, which we refer to as a computational graph as shown in Figure 5.
+In a computational graph, the VERTEX score compares the results produced by a generative model at each node against samples obtained from a reference distribution, usually modeled by sampling from multiple valid references.
+We also account for randomness through predefined random samples for normalizing the result.
+Our reference to hierarchical computational graphs stems from the fact that we operate on multiple levels.
+On a higher level of abstraction we are able to perform planning, sub-task scheduling, and define operational instructions.
+On a lower level of abstraction, we execute these plans based on the defined instructions and data, which can also span generative processes that produce new information.
+Figure 5: We illustrate the hierarchical computational graph for the Paper expression.
+Each node represents an instance of an expression with distinct properties and behaviors, such as file sourcing, generative process, tool utilization, or transformation operation.
+The edges denote the reference relationships between expressions and indicate the flow of information.
+The blue highlighted nodes mark the main sequence nodes of expressions utilized to create parts of the paper draft, such as Method section, RelatedWork section, Abstract section, and so on.
+Each generative node is used for evaluating the VERTEX score.
+None-generative nodes such as search engine results are not evaluated, and we assume to obtain ground-truth values.
+Given that the field is currently at an early stage in developing even sequential schedulers for LLM-based planning systems, our evaluations will be confined to sequential execution only.
+We introduce two tests designed to evaluate multi-step generative processes:
+1) We simulate and evaluate the process of writing a research paper draft based on a predefined hierarchical computational graph that focuses on the content output of the computational graph rather than planning and scheduling functionality.
+See Appendix Section F.5 for more details.
+2) We test the VERTEX Protocol as defined in Algorithm 1, which represents our general method for evaluating multi-step generative processes.
+We create a self-contained test scenario to illustrate an end-to-end evaluation and as a go-to reference for how our protocol can be deployed in a realistic environment.
+Our evaluation protocol is not only designed to analyze and score a series of instructions, but also to provide a structured basis for recording these processes.
+Furthermore, we note that our evaluation protocol is generally formulated, which allows the application of non-sequential planning and scheduling.
 Algorithm 1 VERTEX Protocol
-1:NeSy engine: $\mathcal{V}:\mathcal{S}\rightarrow\mathcal{S}$, where $\mathcal{S}=\bigcup_{n\geq 0}\mathbb{L}^{n}$ is the set of all possible strings formed by concatenating characters from a character set $\mathbb{L}$, embedding engine $\mathcal{E}:\mathcal{S}\rightarrow\mathcal{H}\subset\mathbb{R}^{d}$, symbols $\{x_{0},x^{*},y^{*}\}\subset\mathcal{S}$, with $x_{0}$ as the initial instruction, $x^{*}$ as the payload resulted from executing $\mathcal{V}$, $y^{*}$ as the reference, and *** acting as a placeholder for $\mathcal{P},\mathcal{T},\mathcal{C}$, capabilities $\mathcal{C}=\{\mathcal{F}_{1},\mathcal{F}_{2},\mathcal{F}_{3},\ldots\}$, where each $\mathcal{F}_{i}$ represents a specific functional role within the system, plan $\mathcal{P}\subset\mathcal{S}$, task $\mathcal{T}\in\mathcal{P}$, memory buffer $\mathcal{M}\subset\mathcal{S}$, a scoring function $\tilde{s}:\mathcal{H}\times\mathcal{H}\rightarrow[0,1]$, a scheduler $\mathcal{Q}$, an aggregator $\mathcal{A}$, and score variables $\{s\}\in[0,1]$.
+1:NeSy engine $\mathcal{V}$ as an LLM, embedding engine $\mathcal{E}:\Sigma\rightarrow\mathcal{H}\subset\mathbb{R}^{d}$, symbols $\{x_{0},x^{*},y^{*}\}\subset\Sigma$, with $x_{0}$ as the initial instruction, $x^{*}$ as the payload resulted from executing $\mathcal{V}$, $y^{*}$ as the reference, and * acting as a placeholder for $\mathcal{P},\mathcal{T},\mathcal{C}$, capabilities $\mathcal{C}=\{\mathcal{F}_{1},\mathcal{F}_{2},\mathcal{F}_{3},\ldots\}$, where each $\mathcal{F}_{i}$ represents a specific functional role within the system, plan $\mathcal{P}\subset\Sigma$, task $\mathcal{T}\in\mathcal{P}$, memory buffer $\mathcal{M}\subset\Sigma$, a scoring function $\tilde{s}:\mathcal{H}\times\mathcal{H}\rightarrow[0,1]$, a scheduler $\mathcal{Q}$, an aggregator $\mathcal{A}$, and score variables $\{s\}\in[0,1]$.
 2:
-3:$\mathcal{V},\mathcal{E},\mathcal{Q},\mathcal{C},y^{\mathcal{P}}\leftarrow$ Init($\cdot$) $\triangleright$ Initialize the engines, the scheduler, the capabilities, and expected plan.
-4:$\mathcal{M}\leftarrow\emptyset,\mathcal{A}\leftarrow\emptyset$triangleright Initialize the memory buffer and aggregator.
-5:$x^{\mathcal{P}}\leftarrow$ GeneratePlan($x_{0},\mathcal{V}$) $\triangleright$ Generate a plan starting from the initial instruction.
+3:$\mathcal{V},\mathcal{E},\mathcal{Q},\mathcal{C},y^{\mathcal{P}}\leftarrow$ Init($\cdot$) $\triangleright$ Initialize engines, scheduler, capabilities, expected plan.
+4:$\mathcal{M}\leftarrow\emptyset,\mathcal{A}\leftarrow\emptyset$triangleright Initialize memory buffer and aggregator.
+5:$x^{\mathcal{P}}\leftarrow$ GeneratePlan($x_{0},\mathcal{V}$) $\triangleright$ $\mathcal{V}$ generates plan based on initial instruction.
 6:Evaluate($x^{\mathcal{P}},y^{\mathcal{P}},\mathcal{E},\mathcal{A},\tilde{s}$) $\triangleright$ Embed, score, and aggregate plan similarity.
 […]
-8:while $\mathcal{P}\neq\emptyset$ do $\triangleright$ Run until all tasks have been dequed.
-9:     $\mathcal{T},y^{\mathcal{C}},y^{\mathcal{T}}\leftarrow$ Schedule($\mathcal{M},\mathcal{Q}$) $\triangleright$ Select the next task according to the scheduler and progression.
-10:     $\mathcal{F}_{i}\leftarrow$ Identify($\mathcal{T},\mathcal{C},\mathcal{V}$) $\triangleright$ Identify the task-related capability $\mathcal{F}_{i}$ using $\mathcal{V}$.
-11:     $x^{\mathcal{C}},x^{\mathcal{T}}\leftarrow$ Execute($\mathcal{T},\mathcal{F}_{i}$) $\triangleright$ Execute $\mathcal{T}$ using capability $\mathcal{F}_{i}$ and assign the results $x^{\mathcal{C}},x^{\mathcal{T}}$.
+8:while $\mathcal{P}\neq\emptyset$ do $\triangleright$ Run until list of tasks is exhausted.
+9:     $\mathcal{T},y^{\mathcal{C}},y^{\mathcal{T}}\leftarrow$ Select($\mathcal{M},\mathcal{V}$) $\triangleright$ $\mathcal{V}$ selects next task based on task progression.
+10:     $\mathcal{F}_{i}\leftarrow$ Identify($\mathcal{T},\mathcal{C},\mathcal{V}$) $\triangleright$ $\mathcal{V}$ identifies task-related capability $\mathcal{F}_{i}$.
+11:     $x^{\mathcal{C}},x^{\mathcal{T}}\leftarrow$ Execute($\mathcal{T},\mathcal{F}_{i},\mathcal{Q}$) $\triangleright$ $\mathcal{Q}$ executes $\mathcal{T}$ with capability $\mathcal{F}_{i}$ and assign results $x^{\mathcal{C}},x^{\mathcal{T}}$.
 12:     Evaluate($x^{\mathcal{C}},y^{\mathcal{C}},x^{\mathcal{T}},y^{\mathcal{T}},\mathcal{E},\mathcal{A},\tilde{s}$) $\triangleright$ Embed, score, and aggregate capability similarity.
-13:     $\mathcal{P},\mathcal{M}\leftarrow$ Update($\mathcal{T},\mathcal{P},\mathcal{M},\mathcal{A}$) $\triangleright$ Update the plan and task progression.
+13:     $\mathcal{P},\mathcal{M}\leftarrow$ Update($\mathcal{T},\mathcal{P},\mathcal{M},\mathcal{Q}$) $\triangleright$ $\mathcal{Q}$ updates plan and progression.
 14:end while
 15:$s$leftarrow Finalize($\mathcal{A}$) $\triangleright$ Finalize aggregation of scores.
-16:return $s$triangleright Return the aggregated score of the plan execution.
-Algorithm 1 This algorithm demonstrates the implementation of our Vector Embedding for Relational Trajectory Evaluation through Cross-similarity (VERTEX) protocol.
+16:return $s$triangleright Return aggregated score of plan execution.
+Algorithm 1 This algorithm defines the pseudocode of our VERTEX protocol with our respective VERTEX score as a scoring criteria.
 We start by initializing the NeSy engine $\mathcal{V}$, the embedding engine $\mathcal{E}$, the scheduler $\mathcal{Q}$, and a set of capabilities $\mathcal{C}$.
-The initial instruction $x_{0}$ is used to generate a plan $x^{\mathcal{P}}$ through $\mathcal{V}$.
-The plan and its expected outcome $y^{\mathcal{P}}$ are embedded, and their similarity is scored and aggregated.
+The initial instruction $x_{0}$ is utilized to generate a plan $x^{\mathcal{P}}$ through $\mathcal{V}$.
+The plan and its expected outcome $y^{\mathcal{P}}$ are embedded, and their similarity is scored according to our VERTEX score and aggregated.
 The plan is then unfolded into actionable tasks.
-Each task $\mathcal{T}$ is selected and executed using the appropriate capability $\mathcal{C}$, resulting in the capability and task results $x^{\mathcal{C}},x^{\mathcal{T}}$, and expected outcomes $y^{\mathcal{C}},y^{\mathcal{T}}$ updated in the memory buffer $\mathcal{M}$.
+Each task $\mathcal{T}$ is selected and executed with the appropriate capability $\mathcal{C}$, resulting in the capability and task results $x^{\mathcal{C}},x^{\mathcal{T}}$, and expected outcomes $y^{\mathcal{C}},y^{\mathcal{T}}$ updated in the memory buffer $\mathcal{M}$.
 The process continues, with each task’s result being embedded, scored, and aggregated until the plan is complete.
 The final aggregated score $s$ is returned, reflecting the overall effectiveness of the plan execution.
-Benchmarks GPT-4 GPT-3.5 Gemini-Pro LlaMA 2 Mistral Zephyr Random Associations 0.93 0.61 1.00 0.41 0.49 0.28 0.01 Modality 0.83 0.77 0.75 0.47 0.39 0.43 0.07 Code 0.78 0.58 0.53 0.10 0.19 0.13 0.00 Logic 0.41 0.44 0.39 0.44 0.46 0.43 0.00 Graphs 0.36 0.31 0.21 0.05 0.06 0.08 0.00 Total 0.66 0.54 0.58 0.30 0.32 0.27 0.02
-Figure 5: We evaluate GPT-3.5 Turbo (GPT-3.5), GPT-4 Turbo revision 1106 (GPT-4), Gemini-Pro, LLaMA2-Chat 13B, Mistral 7B and Zephyr 7B on five benchmark categories: 1) Associative Prediction (Association) 2) Multi-modal Binding (Modality) 3) Program Synthesis (Code) 4) Functional Logic Components (Logic) and 5) Hierarchical Computational Graphs (Graphs).
-We denote the similarity score such that higher is better.
-The similarity score is measured according to a reference baseline and normalized to exclude noise.
+We start with a high-level workflow description which consists of a list of tasks and optionally their respective sub-tasks; we refer to this as the plan $\mathcal{P}$.
+To perform the experiment, we utilize an expected plan $y^{\mathcal{P}}$ which was handcrafted for this evaluation.
+The expected plan is a queue of predefined tasks (in a particular order) that the system should follow to achieve the goal.
+The goal statement defines the end objective that the workflow aims to accomplish.
+We also have a set of plans similar to the expected plan, which are trajectories in the solution space, as well as the plan $x^{\mathcal{P}}$ that the LLM generates utilizing the GeneratePlan call for a specific seed.
+We score the predicted plan against the expected plan and the trajectories, then we continue to the next phase in which we utilize the expected plan to execute the tasks.
+At each step, the LLM will receive in its context the goal, the tasks, the current progress, and a query asking for the next task to execute; we refer to this as the memory buffer $\mathcal{M}$.
+If the LLM is not able to predict the next task, it will return a failure, and the expected plan will be utilized to execute the next task.
+The LLM has access to a predefined set of capabilities $\mathcal{C}$, specifically WolframAlpha, SerpApi, Selenium, and the LLM itself, which also represents our self-referential structure.
+We keep executing tasks until the queue is exhausted, and at each step, we utilize the Evaluate call to measure the performance of the LLM with our VERTEX score.
+The scheduler class $\mathcal{Q}$ oversees the execution of the test workflow.
+It takes the setup configuration and orchestrates the linear execution of tasks, utilizing the expected plan as a reference.
+It maintains a pool of tasks to be executed and updates progress as tasks are completed.
+The UnfoldPlan call is a method of the scheduler class $\mathcal{Q}$.
+The method calls itself recursively until there are no tasks left.
+The Select call is responsible for determining which task to execute next from a pool of remaining tasks.
+It utilizes the LLM through self-reflection (Shinn et al., 2023) to choose the most suitable next task based on a template that gets progressively updated in the memory buffer $\mathcal{M}$ by the Update call.
+The Identify call uses self-reflection and similarity scoring to determine the best interface based on the task at hand, then passes the interface to the Execute call to execute the task.
+Lastly, the test ends with the Finalize call, which provides an aggregated assessment of the model’s ability to manage and execute the workflow.
+In Figure 6 we conclude with our evaluation and compute the cumulative score for all described evaluation categories and in the next section we discuss how to interpret the results of our framework.
+Figure 6: We evaluate GPT-4 Turbo, GPT-3.5 Turbo, Gemini-1.0 Pro, LLaMA2-Chat 13B, LLaMA3-Chat 8B, LLaMA3-Chat 70B, Mistral 7B and Zephyr 7B on five benchmark categories: 1) Associative Prediction (Association) 2) Multi-modal Binding (Modality) 3) Program Synthesis (Code) 4) Functional Logic Components (Logic) and 5) Hierarchical Computational Graphs (Graphs).
+We denote the VERTEX scores for each category as a normalized value between 00 and $1$, where higher values are better.
+The VERTEX score is measured according to a reference baseline and normalized by random sequences to exclude noise and similarities among references distributions to rescale solutions.
+The shown scores are an average over all tests per category and across 8 different seeds per test.
+Benchmarks GPT-4 Turbo GPT-3.5 Turbo Gemini 1.0 Pro LLaMA 2 13B LLaMA 3 8B LLaMA 3 70B Mistral 7B Zephyr 7B Random Associations 0.94 0.51 0.93 0.40 0.46 0.57 0.46 0.27 0.08 Modality 0.83 0.65 0.81 0.50 0.43 0.63 0.39 0.43 0.07 Code 0.75 0.63 0.61 0.13 0.25 0.30 0.19 0.13 0.00 Logic 0.48 0.46 0.28 0.46 0.21 0.11 0.53 0.47 0.00 Graphs 0.39 0.34 0.26 0.06 0.05 0.15 0.03 0.03 0.00 Total 0.68 0.52 0.58 0.31 0.28 0.35 0.32 0.27 0.03
 8 Discussion
-In this section, we address the limitations of SymbolicAI and the future directions we are focusing on.
+In this section, we address the evaluation results, auxiliary findings and limitations of SymbolicAI and the future directions we are focusing on.
 Some of the limitations stem from the inherent constraints of current technologies and dependencies on third-party systems.
 Additionally, the nuanced complexities of working with generative models presents further challenges.
-7.1 Limitation
+8.1 Results
+In Figure 6 we show the VERTEX score for all five evaluation categories on 8 different state-of-the-art models.
+We show the aggregated results per category, meaning the average score among all tests averaged per category and average across 8 different seeds per test.
+The VERTEX score is normalized between 00 and $1$, where higher values are better.
+Our score is non-linear due to its nature of using non-linear kernels, and captures semantic, ordinal and relative structures among the data samples.
+However, since our score is highly dependent on the quality of the underlying embedding model, it may omit to capture fine-grained syntactic differences such as ‘Hello‘ vs ‘hello‘.
+In our experiments, we have noticed that for associative predictions and multi-modal bindings, GPT-4 Turbo is on par with Gemini-1.0 Pro.
+Furthermore, there is still a large gap between open-source contestants such as LLaMA 3 even with 70B parameters compared to the closed-source alternatives from OpenAI and Google.
+For the rest of the experiments, we see that GPT-4 almost always outperforms all other models, except for the functional logic components category.
+Here, we analyzed results and found that the larger models sometimes take shortcuts by automatically returning the solution and answering that the task instructions are too complex for such a straight-forward puzzle query.
+However, we would rather state in general that for logic-based, planning and scheduling tasks all models act unreliably, even if slight performance differences between the models are seen in the plot.
+We believe this is in part due to lack of training data specifically for workflows, planning and scheduling tasks, and to imprecision in generating reliably structured output formats, such as custom DSLs or other custom in-context instructed formats.
+This also stems from their instruction fine-tuning, since most models are chat-based models and offer verbose responses which need to be suppressed or post-processed.
+We see similar performance between GPT-3.5 Turbo and LLaMA 3 70B except for the logical and graphs evaluations.
+We found that LLaMA 3 70B has a tendency to ask questions back if it does not understand the request instead of following the specified instructions provided.
+We assume this also stems from the chat-based instruction fine-tuning.
+Zephyr 7B and Mistral 7B have shown on par capabilities in functional logic components with larger models, however fail in program synthesis and hierarchical computational graphs experiments.
+We observe that they perform well when resolving the overloaded logic operators such as OR, AND and XOR, and show decent performance for text generation, but fail to resolve more complex instructions.
+8.2 Limitations
+Framework
+Since the framework interfaces with many tools and API services, it requires a substantial engineering feat to integrate all available functionalities and keep the API-based services up-to-date.
+For us this means, that although we support a variety of tools and frameworks like Selenium, WolframAlpha, or Z3, we only scratch the surface of these tools.
+Moreover, the utilization of grammar-based constraints validations is still experimental and limited in functionality for specific formats such as JSON and HTML.
+Finally, we encounter also challenges related to engineering parallelization and multiprocessing of prompts, since the concurrent execution is non-trivial, especially with intricacies of Python process management.
 Embedding Measure
@@
 Since we were very limited in the availability of development resources, and some presented models are only addressable through costly API walls.
-We are actively seeking investors and supporters to scale our solution and offer a more compelling benchmark suite in the future.
+We are actively seeking sponsors to scale our solution and offer a more compelling benchmark suite in the future.
 Model Capabilities
@@
 Furthermore, our framework’s reliance on the model’s ability to grasp the semantics of operations can be influenced by the training data and the model’s innate biases and conceptual understanding (Mahowald et al., 2023).
-We also point out that the initial development of SymbolicAI started with the GPT family of models, and we may encounter innate biases in prompt design and expressiveness when using other reference models.
+We also point out that the initial development of SymbolicAI started with the GPT family of models, and we may encounter innate biases in prompt design and expressiveness when utilizing other reference models.
 However, we also point out that prompt engineering instruction-based statements is not a reliable direction for improvement.
@@
 However, our objective is to ultimately assess a non-sequential task execution model, allowing for dynamic insertion and out-of-sequence task execution.
-In addition, we are interested into exploring similarities of our work with Generative Flow Networks (Bengio et al., 2021a, b; Lahlou et al., 2023).
-7.3 Broader Impact
-With LLMs becoming more and more accessible, progress recently made possible by the vast open source contributions from Köpf et al.
-(2023); Touvron et al.
-(2023); Taori et al.
-(2023); Xu et al.
-(2023); Geng et al.
-(2023); Biderman et al.
-(2023), embedded accelerators for LLMs – or more generally NeSY engines – will be ubiquitous in future computation platforms, such as wearables, smartphones, tablets, consoles, or notebooks.
-Although current execution cycles are slow and error-prone, we expect to see further performance gains through improved operating system level optimizations, dedicated GPU-centric hardware refinement, and improved software interoperability.
-We believe that modern programming paradigms should natively support probabilistic concepts and provide a boilerplate-free set of features for constructing and evaluating generative computational graphs.
-This includes but is not limited to compositional, parallelizable, and simulation-based executions with polymorphic and self-referential structures.
-Current programming languages often have disjointed or makeshift solutions for these concepts in the context of generative processes.
-We believe integral probabilistic support for these concepts into modern software and hardware will unlock new programming paradigms that can fully take advantage of generative architectures.
-We hope the community will consider these ideas as essential components of contemporary computing.
-We also expect to see significant progress by processing central language concepts through system-on-a-chip (SoC) solutions of pre-trained models, with linear probing layers for hot-swappable weight exchange of task-specific projections and executions.
-A wide range of functionalities can be then offloaded to probabilistic programming languages to operate on dedicated symbols and streamline the vector-valued mappings between the concept space and underlying problem space, avoiding defining boilerplate code to load and unload network weights.
-Furthermore, we believe that many gains in representational stability and consistency may be obtained through multi-modal data training and improved alignment based on operator learning oriented functionalities and workflow-related scoring functionalities, analogous to our introduced quality measure.
-Gains in representational stability also benefit self-instruction and self-referential sub-process evaluations, which enable the dynamic creation and evaluation of complex hierarchical computational graphs.
-This will enable online learning models to perform, in real-time, skill acquisition of complex concepts with only one or few examples at inference time.
-We believe this will enable the creation of autonomously self-evolving cognitive architectures (Langley et al., 2009; Dawid & LeCun, 2023; Sumers et al., 2023).
-We therefore see an inherent connection to generative design as an analogy for creating coherent and stable ”thought” computational graphs, and believe this paves the path toward broad AI systems (see Section B) and is a requirement for developing General AI agents.
-Finally, we also wish to express our concern about recent economic trends in the deep-tech industry, where we observe AI-related concentration of data and resources, coupled with a tendency towards closed-source practices.
-We strongly advocate for increased transparency and exchange of ideas to ensure a diverse and collective growth in our socio-economic landscape.
-Therefore, we push towards a democratic and open-source initiative.
-7.4 Conclusion
+In addition, we are interested into exploring similarities of our work with Generative Flow Networks (Bengio et al., 2021a; b; Lahlou et al., 2023).
+Lastly, we also discuss limitations in Appendix Section 8.2 with further opportunities for future improvements.
+9 Conclusion
 In this work, we introduced SymbolicAI, a framework that unifies generative models with an array of solvers, blending the strengths of symbolic and sub-symbolic AI paradigms within a cohesive NeSy framework.
@@
 SymbolicAI provides a basis for further research in advanced program synthesis, hierarchical computational graphs, the development of self-referential systems, and the integration of probabilistic methods with AI design for creating autonomous agents.
-Furthermore, our commitment to open-source ideals attempts to promote a culture of collaborative growth and innovation.
 Acknowledgement
@@
 It is with great respect and affection that we acknowledge the indelible impact they have made, enabling us to persist in our scientific pursuits with determination and integrity.
-Appendix
-Appendix A Connection between Fréchet Distance and Maximum Mean Discrepancy
+Appendix A Broad AI and Neuro-Symbolic Systems
+Our work focuses on broad AI (Hochreiter, 2022) (see Figure 7) through the integration of symbolic and sub-symbolic AI methodologies.
+Broad AI extends beyond restricted focus on single-task performance of narrow AI.
+In broad AI, systems are engineered to handle a range of tasks with a high degree of autonomy, utilizing sensory input, accumulated experiences, and previously developed skills.
+Figure 7: Hierarchical model of ”cognitive” abilities of AI systems (Chollet, 2019; Hochreiter, 2022).
+The figure contrasts the emergent paradigm of Broad AI with current Narrow AI systems, showcasing Broad AI’s wider range of capabilities, such as knowledge transfer, interaction, adaptability, robustness, abstraction, advanced reasoning, and efficiency.
+Broad AI aims to mimic human cognitive adaptability and robustness through advanced methodologies like few-shot learning, self-supervised contrastive learning, and context-sensitive sensory processing.
+Notably, Broad AI applies principles such as conceptual short-term memory and modern Hopfield networks (Ramsauer et al., 2020) to better integrate context and memory, thus avoiding pitfalls like explaining away and short-cut learning.
+We acknowledge the potential of NeSy systems as a significant step towards AI systems capable of performing any cognitive task with human-like proficiency.
+NeSy methods form the basis for developing new cognitive architectures (Newell & Simon, 1956; Newell et al., 1957; Newell & Simon, 1972; Newell, 1990; Langley et al., 2009; Laird, 2022; Dawid & LeCun, 2023; Sumers et al., 2023; LeCun, 2022; Assran et al., 2023).
+This hybridization produces computational graphs capable of context-aware learning and reasoning, allowing AI to execute complex tasks with human-like flexibility.
+Figure 8: Evolution of software paradigms: From Software 1.0’s rigid specification in classical programming to Software 2.0’s data-driven and objective function-focused differentiable programming, leading to Software 3.0’s NeSy systems that emphasize human-centric, interaction-based programming with computational graphs.
+This progression represents a shift from explicit task-specific programming to abstract, adaptive systems that cater to dynamic user preferences.
+Borrowing nomenclature from Karpathy (2017); Dilhara et al.
+(2021), we refer to the next generation of software as Software 3.0, which consists of applications that are neither pre-determined at design time, nor learned through statistical inference, but triggered by an interaction which stimulates the realization of a computational graph analogous to neuromorphic circuits (Indiveri et al., 2011), however, purely established at inference time in the ”thought” process of a NeSy system.
+To enable such systems, we require a more native integration (see illustration in Figure 9) of probabilistic programming paradigms into our contemporary programming stack, and make their utilization a commodity for practitioners and researchers alike.
+Figure 9: The illustration demonstrates the integration of Neuro-Symbolic computation within the contemporary programming stack.
+Probabilistic programming paradigms are embedded natively alongside traditional languages and environments, facilitated by interfaces to various data sources, compute engines, and human interaction tools, streamlining their adoption in practical and research applications.
+A.1 Broader Impact
+With LLMs becoming more and more accessible, progress recently made possible by the vast open source contributions from Köpf et al.
+(2023); Touvron et al.
+(2023); Taori et al.
+(2023); Xu et al.
+(2023); Geng et al.
+(2023); Biderman et al.
+(2023), embedded accelerators for LLMs — or more generally NeSY engines — will be ubiquitous in future computation platforms, such as wearables, smartphones, tablets, consoles, or notebooks.
+Although current execution cycles are slow and error-prone, we expect to see further performance gains through improved operating system level optimizations, dedicated GPU-centric hardware refinement, and improved software interoperability.
+We believe that modern programming paradigms should natively support probabilistic concepts and provide a boilerplate-free set of features for constructing and evaluating generative computational graphs.
+This includes but is not limited to compositional, parallelizable, and simulation-based executions with polymorphic operations and self-referential structures.
+Current programming languages often have disjointed or makeshift solutions for these concepts in the context of generative processes.
+We believe integral probabilistic support for these concepts into modern software and hardware will unlock new programming paradigms that can fully take advantage of generative architectures.
+We hope the community will consider these ideas as essential components of contemporary computing.
+We also expect to see significant progress by processing central language concepts through system-on-a-chip (SoC) solutions of pre-trained models, with linear probing layers for hot-swappable weight exchange of task-specific projections and executions.
+A wide range of functionalities can be then offloaded to probabilistic programming languages to operate on dedicated symbols and streamline the vector-valued mappings between the concept space and underlying problem space, avoiding defining boilerplate code to load and unload network weights.
+Furthermore, we believe that many gains in representational stability and consistency may be obtained through multi-modal data training and improved alignment based on operator learning oriented functionalities and workflow-related scoring functionalities, analogous to our introduced quality measure.
+Gains in representational stability also benefit self-instruction and self-referential sub-process evaluations, which enable the dynamic creation and evaluation of complex hierarchical computational graphs.
+This will enable online learning models to perform, in real-time, skill acquisition of complex concepts with only one or few examples at inference time.
+We believe this will enable the creation of autonomously self-evolving cognitive architectures (Langley et al., 2009; Dawid & LeCun, 2023; Sumers et al., 2023).
+We therefore see an inherent connection to generative design as an analogy for creating coherent and stable ”thought” computational graphs, and believe this paves the path toward broad AI systems (see Section A) and is a requirement for developing artificial general intelligent agents.
+Finally, we also wish to express our concern about recent economic trends in the deep-tech industry, where we observe AI-related concentration of data and resources, coupled with a tendency towards closed-source practices.
+We strongly advocate for increased transparency and exchange of ideas to ensure a diverse and collective growth in our socio-economic landscape.
+Therefore, we push towards a democratic and open-source initiative.
+Appendix B Connection between Fréchet Distance and Maximum Mean Discrepancy
 Let us consider a Gaussian kernel defined by the expression
@@
 Heuristically, at least for small $\|\mu_{1}-\mu_{2}\|$, also the associated kernel evaluations $K(X_{1},X_{2})$ tend to be small (see also Hochreiter & Schmidhuber (1997)), which leads to a small MMD, if we ignore the terms associated to $K(X_{1},X_{1})$, $K(X_{2},X_{2})$ (which cancel out due to same covariance structure).
-In the next section we want to further elaobarate on the MMD and a possible score, that can be derived from it.
-A.1 Extended Simplification of the MMD Calculation
+In the next section, we want to further elaborate on the MMD and a possible score, that can be derived from it.
+B.1 Extended Simplification of the MMD Calculation
 To understand the simplification of the MMD, we are formally expressing the MMD in terms of kernel sums over pairs of samples within and across two samples $X$ and $Y$:
@@
 $\widetilde{\text{MMD}^{2}}(X,Y)\approx\frac{2}{mn}\sum_{i=1}^{m}\sum_{j=1}^{n}k(x_{i},y_{j}).$ (11)
-Appendix B Broad AI and Neuro-Symbolic Systems
-Our work focuses on broad artificial intelligence (AI) (Hochreiter, 2022) (see Figure 6) through the integration of symbolic and sub-symbolic AI methodologies.
-Broad AI extends beyond restricted focus on single-task performance of narrow AI.
-In broad AI, systems are engineered to handle a range of tasks with a high degree of autonomy, utilizing sensory input, accumulated experiences, and previously developed skills.
-Figure 6: Hierarchical model of ”cognitive” abilities of AI systems (Chollet, 2019; Hochreiter, 2022).
-The figure contrasts the emergent paradigm of Broad AI with current Narrow AI systems, showcasing Broad AI’s wider range of capabilities, such as knowledge transfer, interaction, adaptability, robustness, abstraction, advanced reasoning, and efficiency.
-Broad AI aims to mimic human cognitive adaptability and robustness through advanced methodologies like few-shot learning, self-supervised contrastive learning, and context-sensitive sensory processing.
-Notably, Broad AI applies principles such as conceptual short-term memory and modern Hopfield networks (Ramsauer et al., 2020) to better integrate context and memory, thus avoiding pitfalls like explaining away and short-cut learning.
-We acknowledge the potential of NeSy systems as a significant step towards AI systems capable of performing any cognitive task with human-like proficiency.
-NeSy methods form the basis for developing new cognitive architectures (Newell & Simon, 1956; Newell et al., 1957; Newell & Simon, 1972; Newell, 1990; Langley et al., 2009; Laird, 2022; Dawid & LeCun, 2023; Sumers et al., 2023; LeCun, 2022; Assran et al., 2023), as illustrated in Figure 7.
-This hybridization produces computational graphs capable of context-aware learning and reasoning, allowing AI to execute complex tasks with human-like flexibility.
-Figure 7: Evolution of software paradigms: From Software 1.0’s rigid specification in classical programming to Software 2.0’s data-driven and objective function-focused differentiable programming, leading to Software 3.0’s NeSy systems that emphasize human-centric, interaction-based programming with computational graphs.
-This progression represents a shift from explicit task-specific programming to abstract, adaptive systems that cater to dynamic user preferences.
-Borrowing nomenclature from Karpathy (2017); Dilhara et al.
-(2021), we refer to the next generation of software as Software 3.0, which consists of applications that are neither pre-determined at design time, nor learned through statistical inference, but triggered by an interaction which stimulates the realization of a computational graph analogous to neuromorphic circuits (Indiveri et al., 2011), however, purely established at inference time in the ”thought” process of a NeSy system.
-To enable such systems, we require a more native integration (see illustration in Figure 8) of probabilistic programming paradigms into our contemporary programming stack, and make their utilization a commodity for practitioners and researchers alike.
-Figure 8: The illustration demonstrates the integration of Neuro-Symbolic computation within the contemporary programming stack.
-Probabilistic programming paradigms are embedded natively alongside traditional languages and environments, facilitated by interfaces to various data sources, compute engines, and human interaction tools, streamlining their adoption in practical and research applications.
-Appendix C Framework Structure
+Appendix C Structure
 Primitives
@@
 Expressions extend the capabilities of Symbol objects by providing a structured way to define complex functionalities that can later be evaluated to produce new Symbol objects or modify existing ones.
-In SymbolicAI, an Expression object inherits the properties of Symbol objects while also being able to define its own unique behavior through the use of a forward method, which is analogous to a computational graph node’s evaluation function.
+In SymbolicAI, an Expression object inherits the properties of Symbol objects while also being able to define its own unique behavior through a forward method, which is analogous to a computational graph node’s evaluation function.
 Figure 10 gives an overview of an exemplary inheritance branch.
@@
 For instance, SymbolicAI integrates the zero- and few-shot learning with default fallback functionalities of pre-existing code.
-The use of decorators brings several advantages (Beazley, 2009; Martelli et al., 2005; Summerfield, 2010; Lutz, 2013):
+Decorators brings several advantages (Beazley, 2009; Martelli et al., 2005; Summerfield, 2010; Lutz, 2013):
   • Reusability: Decorators promote code modularity, enhancing code reusability and contributing to software maintainability.
@@
   • Composition: Decorators support function composition, allowing developers to construct complex functionalities from pre-existing code blocks without the need to expand the codebase or rely on complex inheritance hierarchies.
-  • Adaptability: Using decorators we can easily modify or extend the behavior of operations without changing their core implementation.
+  • Adaptability: Through decorators we can easily modify or extend the behavior of operations without changing their core implementation.
 This flexibility facilitates the generation of adaptive workflows and reliable fallback mechanisms when experimental implementations do not fulfill required constraints.
@@
 Expression of a Computational Graph
-In practice, consider the Expression class, which extends the functionality of ‘Symbol‘.
+In practice, consider the Expression class, which extends the functionality of the Symbol class.
 When composing a Sequence of Expression objects, we are effectively composing operations in a predetermined order.
-For instance, an expression such as:
+For instance, an expression like:
 1Sequence(
@@
 When this sequence is executed, the operations unfold in the exact order specified, with each step receiving the output of its predecessor as input and passing its result to the successor.
-Generating entire workflows
-SymbolicAI allows the creation of entire workflows, such as writing scientific papers.
-The following example defines a Paper expression that takes in a sequence of expressions which are executed in sequentially.
-The Method expression contains a Source expression, which addresses the code base of the actual method of the paper.
-The RelatedWork expression contains a sequence of Cite expressions which are executed in parallel and are used to define the context of the related work section.
-Each expression in the sequence of expressions from Paper takes in the context of its predecessors.
-All expressions also link their results to a global linker object, which is used after the execution to retrieve individual results from distinct expression nodes.
-In Figure 10 we show the resulting computational graph of the Paper expression.
-1# define the computational graph
-2expression = Paper(
-3  Method(
-4    # link to original code base where the main method is defined
-5    Source(file_link=’/path/to/.../file’),
-6  ),
-7  # gather resources and write the related work
-8  RelatedWork(
-9    Cite(file_link=’/path/to/.../file’),
-10    Cite(file_link=’/path/to/.../file’),
-11    ...
-12  ),
-13  # write the abstract and title
-14  Abstract(),
-15  Title(),
-16)
-17# run the graph
-18paper_result = expression(’Write a scientific paper’)
-19# access linker to retreive the results from the method expression
-20method = expr.linker.find(’Method’)
-21# print result of the method expression
-22print(method)
-Figure 10: We illustrate the computational graph for the Paper expression.
-Each node represents an instance of an expression with distinct properties and behaviors.
-The edges denote the reference relationship between expressions.
-The blue highlighted nodes mark the main sequence of expressions, used to create the paper, such as Method, RelatedWork, Abstract, etc.
 Appendix D Installation
@@
 This command will install the latest version of SymbolicAI along with its core dependencies, enabling the integration of the framework into Python applications.
-If you intend to use the framework with local engines444 The local engines are experimental and are run on your local machine.
+If you intend to utilize the framework with local engines555 The local engines are experimental and are run on your local machine.
 For more details, refer to the ”Local Neuro-Symbolic Engine” section in the documentation., or with engines powered by external APIs such as OpenAI’s API, additional installation steps are required.
@@
 2export OPENAI_API_KEY="<OPENAI_API_KEY>"
-After setting up the API keys, the SymbolicAI library is imported in Python using:
+After setting up the API keys, the SymbolicAI library is imported in Python as following:
 1import symai
@@
 This documentation will give detailed steps on installing optional dependencies and configuring additional API keys.
-D.3 Benchmark
-To run the benchmark evaluation you first need to install the ExtensityAI/benchmark plugin:
-1sympkg i ExtensityAI/benchmark
-Then you can go to the \$HOME/.symai/packages/ExtensityAI/benchmark directory and run the benchmark through:
-1python test.py
-To compile models optimized for your machine, follow the instructions from the ExtensityAI/benchmark repository.
 Appendix E Implementation Details
@@
 E.1 Fuzzy Comparison
-For instance, let’s consider the use of fuzzy555 Not related to fuzzy logic, which is a topic under active consideration.
+For instance, let’s consider fuzzy666 Not related to fuzzy logic, which is a topic under active consideration.
 comparisons.
@@
 2<class ’symai.expressions.Symbol’>(value=Hello my friend)
-E.3 Translation
+E.3 Symbols and Embeddings
+It is worth noting that encoding a complex object into a string sometimes precludes the object reconstitution.
+However, this concern does not substantially impede our methodology as we can employ approximations or proxy representations stored by the vector-valued property to effectively re-map objects.
+These representations are obtained through respective embedding models.
+E.4 Translation
 In today’s increasingly interconnected world, translation between languages is fundamental, making it an essential feature.
@@
 2<class ’symai.expressions.Symbol’>(value=Willkommen zu unserem Tutorial.)
-E.4 Filtering, Ranking, Extraction
+E.5 Filtering, Ranking, Extraction
 Incorporating data-agnostic operations like filtering, ranking, and pattern extraction into our API allow the users to easily manipulate and analyze diverse data sets.
@@
 2<class ’symai.expressions.Symbol’>(value=[’7’, ’6’, ’5’, ’4’, ’3’, ’2’, ’1’])
-E.5 Implications
+E.6 Implications
 One of the main objectives behind developing SymbolicAI was to facilitate reasoning capabilities in conjunction with the statistical inference inherent in LLMs.
-Consequently, we can carry out deductive reasoning operations utilizing the Symbol objects.
+Consequently, we can carry out deductive reasoning operations through the Symbol objects.
 For instance, it is feasible to establish a series of operations with rules delineating the causal relationship between two symbols.
@@
 In the above example, the & operator overloads the and logical operator and extends its functionality.
-Furthermore, we can establish more sophisticated logical operators for and, or, and xor that can be grounded in formal proofs, as well as utilize the NeSy engine to parse data structures before evaluating the expressions.
+Furthermore, we can establish more sophisticated logical operators for and, or, and xor that can be grounded in formal proofs and utilize the NeSy engine to parse data structures before evaluating the expressions.
 This enables the definition of bespoke operations for executing intricate and robust logical operations, incorporating constraints to validate outcomes and guide the computation towards the desired behavior.
-E.6 Custom operations
+E.7 Custom operations
 The following example demonstrates how to define a custom == operation by overriding the __eq__ method and providing a custom prompt object with a list of examples:
@@
     We do not necessarily want all of our operations to be sent to the NeSy engine and might need to implement a default behavior.
-This is achieved using the _sym_return_type method, which can provide contextualized behavior based on the defined return type. More details can be found in the actual Symbol class.
-In the context of LLMs, zero- and few-shot learning domains have emerged as essential techniques (Yao et al., 2023b; Shinn et al., 2023; Kim et al., 2023; Wei et al., 2022b; Lyu et al., 2023; Pitis et al., 2023; Madaan et al., 2023; Wang et al., 2022; Ye et al., 2023)666 This is by no means an exhaustive list, we only point the reader to some very influential and recent research.
+This is achieved through the _sym_return_type method, which can provide contextualized behavior based on the defined return type. More details can be found in the actual Symbol class.
+In the context of LLMs, zero- and few-shot learning domains have emerged as essential techniques (Yao et al., 2023b; Shinn et al., 2023; Kim et al., 2023; Wei et al., 2022b; Lyu et al., 2023; Pitis et al., 2023; Madaan et al., 2023; Wang et al., 2022; Ye et al., 2023)777 This is by no means an exhaustive list, we only point the reader to some very influential and recent research.
 to enable models to generalize from limited training data and adapt to new tasks without requiring extensive retraining.
@@
 In this section, we demonstrate how our Symbolic API incorporates Python decorators to define custom operations in the zero- and few-shot domains.
-Consider the following example, where we define a custom operation to generate a random integer between 0 and 10 using the Symbolic API and Python decorators:
+Consider the following example, where we define a custom operation to generate a random integer between 00 and $10$ using the Symbolic API and Python decorators:
 1import symai
@@
 The examples and limit arguments are new, with examples defining a list of instructions conditioning the NeSy engine, and limit specifying the maximum number of examples returned.
-The pre_processor and post_processor arguments accept lists of PreProcessor and PostProcessor objects, respectively, which are used to process the input before being fed into the NeSy engine and the output before being returned to the user.
+The pre_processor and post_processor arguments accept lists of PreProcessor and PostProcessor objects, respectively, which are utilized to process the input before being fed into the NeSy engine and the output before being returned to the user.
 The wrp_kwargs argument passes additional arguments to the wrapped method, streamlining them towards the NeSy engine, or other engines.
-E.7 Prompting
+E.8 Prompting
 In this section, we discuss the design of prompts and their role in shaping the behavior of operations.
 Prompts serve as containers for information necessary to define specific operations, and the Prompt class serves as the base class for all the other Prompt classes.
-Consider the following example, where we define a Prompt for comparing two values using the NeSy engine.
+Consider the following example, where we define a Prompt for comparing two values through the NeSy engine.
 In it, when the <= operation on two Symbol objects will be resolved, the NeSy engine evaluates them in the context of the CompareValues prompt.
@@
 10 ])
+Evaluating a fuzzy comparison statement:
 1res = symai.Symbol(1) <= symai.Symbol(’one’)
+Output of the evaluation:
 1:[Output]:
@@
     Context-based Prompts are considered optional and can be defined in a static manner, either by sub-classing the Expression class and overwriting the static_context property, or at runtime by updating the dynamic_context property or passing a payload kwargs to a method.
-Below is an example of using the payload kwargs through the method signature:
+Below is an example of payload kwargs through the method signature:
     1# creating a query to ask if an issue was resolve or not
@@
     Regardless of how the context is set, the contextualized prompt defines the desired behavior of Expression operations.
-For instance, if we want to operate in the context of a DSL without having to overwrite each base class method, we can use this approach777 See more details in this notebook..
+For instance, if we want to operate in the context of a DSL without having to overwrite each base class method, we can utilize this approach888 See more details in this notebook..
   2. 2.
@@
     Template Prompts are optional and encapsulate the resulting prediction to enforce a specific format.
-For example, to generate HTML tags, we can use a curated <html>...</html> template.
+For example, to generate HTML tags, we can utilize a curated <html>...</html> template.
 This template enforces the NeSy engine to begin the generation process already in the context of an HTML tags format and not produce irrelevant descriptions about its task.
-E.8 Complex expressions
+E.9 Complex expressions
 We will now attempt to obtain logical answers based on questions of the kind:
@@
 1def causal_expression(self):
-2  # very which case to use ‘self.value‘ contains the input
 2 if self.isinstanceof(’mathematics’):
@@
 14 res = self.expression(query) # send prepared query to wolframalpha
-16
+15
 16 elif formula.isinstanceof(’number comparison’):
 17 res = formula.expression() # send directly to wolframalpha
-19
+18
 19 ... # more cases
-21
+20
 21 elif self.isinstanceof(’graph construction’):
@@
 30 ... # add relations and populate graph; reading suggestion
-32
+31
 32 ... # more cases
-34
+33
 34 return res
@@
 Initially, when utilizing the GPT-3 backend, we anticipated a significant engineering effort to develop such a complex expression, as the GPT-3 backend frequently struggled with accurate information extraction and comparison resolution.
-However, we remained confident in the field’s progress, specifically with fine-tuned models like ChatGPT utilizing RLHF.
+However, we remained confident in the field’s progress, specifically with fine-tuned models like RLHF ChatGPT.
 We were delighted to witness these challenges being further tackled through the latest GPT-4 model (OpenAI, 2023).
@@
 Moreover, recent advances in LLMs have opened new frontiers for research, practice, and adoption of causal reasoning, transforming the way causal analysis is conducted and broadening the scope of applications for our framework.
-One of the most prominent illustrations of this concept is exhibited by Word2Vec (Mikolov et al., 2013).
+One of the most prominent illustrations of this concept is exhibited by Word2Vec (Mikolov et al., 2013a).
 Word2Vec generates dense representations of words by training a shallow neural network to predict a word based on its neighboring words within a text corpus.
@@
 These operations are intentionally separated from Symbol, as they do not utilize the value attribute of the Symbol class.
-E.9 Composition
-E.10 Sequences
+E.10 Composition
+E.11 Sequences
 Sequences offer a multitude of advantages in the realm of Expression objects, as they facilitate the creation of more sophisticated structural configurations.
@@
 9)
-E.11 Streams
+E.12 Streams
 As demonstrated earlier, creating contextualized prompts refines the behavior of operations in the NeSy engine.
@@
 Stream expressions can encapsulate other expressions.
-For instance, chunks can be managed using a Sequence expression, which permits multiple compositional operations in a sequential manner.
+For instance, chunks can be managed through a Sequence expression, which permits multiple compositional operations sequentially.
 The example below illustrates the definition of a Stream expression:
@@
 These combinations allow for a more effective approach to handling context limitations, promoting the extraction of meaningful information and improving the overall performance of the framework.
-E.12 Error handling, debugging, and explainability
+E.13 Error handling, debugging, and explainability
 Effective error handling and debugging are essential for ensuring the robustness and reliability of any software system, while explainability is essential for understanding the underlying behavior of the system, especially in the context of AI-driven applications.
 By developing a system that is both transparent and interpretable, we can gain valuable insights into the performance of the NeSy engines and identify potential areas for improvement.
-E.13 Error handling
+E.14 Error handling
 One of the fundamental aspects of the SymbolicAI API is being able to generate code.
@@
 In pursuit of a self-evolving API, we introduce the Try expression, which includes built-in fallback statements and automatically retries execution after performing dedicated error analysis and correction.
-This expression analyzes both the input and the error, conditioning itself to resolve the error by modifying the original code888 This is similar to the recently released Auto-GPT application..
+This expression analyzes both the input and the error, conditioning itself to resolve the error by modifying the original code999 This is similar to the recently released Auto-GPT application..
 If the fallback expression succeeds, the result is returned; otherwise, the process is repeated for the specified number of retries.
@@
 2a = 3
-While not all errors can be resolved as easily as the demonstrated syntax error, we continue to explore more sophisticated error handling mechanisms, including the use of streams and clustering to address errors in a hierarchical and contextual manner.
+While not all errors can be resolved as easily as the demonstrated syntax error, we continue to explore more sophisticated error handling mechanisms, including streams and clustering to address errors in a hierarchical and contextual manner.
+Appendix F Evaluation Details
+In a computational graph, the VERTEX score compares the distribution of the generated model answer at each node against a reference distribution by sampling multiple valid trajectories at each node for the reference distribution and accounting for randomness through some predefined random trajectories.
+For instance, one of the predefined random trajectories in our benchmark was the string of ASCII characters which are considered printable, namely 0123456789abcdefghijklmnopqrstuvwxyzABCDEFGHIJKLMNOPQRSTUVWXYZ!"#\$%&’()*+, -./:;<=>?@[]_̂‘{|}~.
+Moreover, the VERTEX score is particularly well suited for the evaluation of multi-step workflows and in contexts where the solution space is or is expected to be diverse.
+We will now proceed by describing in detail the tasks that we defined in our benchmark.
+F.1 Associative Prediction
+We defined a total of 15 tasks involving in-context associations between two Symbol instances.
+SymbolicAI’s overloaded operators rely on predefined pseudo-grammars, as described in Section 4, that augment the operators with few-shot examples.
+For instance, the overloaded operator + utilized between two Symbol instances provides few-shot examples how to resolve additions with various data types:
+1"’1’ + 2 =>3",
+2"17 + pi =>20.1415926535...",
+3"7.2 + five =>12.2",
+4"True + 0 => False",
+5"False + True =>False",
+6"[’a’, b’] + [’c’, d’] =>[’a’, b’, c’, d’]",
+7"[’apple’] + banana =>[’apple’, banana’]",
+8"’Zero + 1 =>1",
+9"’One + Two =>3",
+10"’Three + 4 =>7",
+11"’a + b + c + d =>a + b + c + d",
+12...
+Consequently, we can now test if the models can solve the addition between Symbol("two hundred and thirty four") and Symbol(7000).
+F.2 Multi-modal Binding
+We perform transformations between multiple modalities through language-based representations.
+Therefore, we need to evaluate the model’s proficiency in tool utilization, classification and routing of requests to relevant modules.
+We define a multi-modal Expression to detect the category of a task based on its content and to forward the task to the appropriate tool.
+The expression creates interfaces to tools like WolframAlpha for mathematical expressions, Selenium for website content scraping, SerpApi for search queries, and APILayer for optical character recognition.
+Each of the five tests aims to evaluate the appropriate handling of a specific type of input by the multi-modal Expression type, such as processing a website URL for scraping, interpreting a search engine query, testing if two vectors are linearly independent, comparing large numbers, and extracting text from an image.
+The following example shows the MultiModalExpression implementation of the forward function that uses isinstanceof operator on its own context to determine its current expression value and select the sub-routine that can evaluate the request.
+1class MultiModalExpression(Expression):
+2 def forward(self, ...):
+3 formula = self.extract(’mathematical formula’)
+4 ...
+5 if self.isinstanceof(LINEAR_ALGEBRA):
+6 ...
+7 res = self.solver(formula)
+8 res = res.query(’write a one sentence summary of the answer’)
+9 ...
+10 elif self.isinstanceof(NUMBER_COMPARISON):
+11 res = self.solver(formula) # send directly to wolframalpha
+12 else:
+13 ...
+14
+15query = Symbol("is 100044347 bigger than 129981063.472?")
+16expr = MultiModalExpression(query)
+17res = expr(...)
+F.3 Program Synthesis
+We designed three separate tests related to program synthesis, where each task assesses the ability of the models to generate and execute code based on natural language instructions or provided templates:
+1) The first task involves reading a LaTeX table template and data, then generating a function to populate the table with the given data.
+2) The second task tests the automatic code generation for API calls by fetching data from a specified URL and extracting specific information from the retrieved content.
+3) The third task evaluates the ability to construct a custom Expression that processes a Symbol through a specific Function component from the SymbolicAI package.
+Each of the three tests follows a similar pattern, where the generated code is scored based on its similarity to valid references and normalized with random samples.
+Figure 13 shows possible samples from the third task category.
+Figure 13: Samples for constructing custom expressions.
+The expected references are human-generated and are distributed according to the solution space.
+The set of valid references can be as well human-generated.
+The set of invalid (or undesirable) samples represent the subset we are the least interested in.
+Lastly, note that in general, the references need not be human-generated; any combination of human- and machine-generated references is possible.
+For instance, if synthetic data is utilized to distill knowledge from a larger model to a smaller model, the valid references and the expected references can be sampled from the larger model.
+F.4 Logical Components
+We designed six tests to assess the logical capabilities of the candidate models and group them as follows.
+1) We utilize the Python library SymPy for symbolic mathematics to create the mathematical expression $ax+bx-cx-ay-by+cy+d$.
+The task for the model is then to factorize the expression and extract all unique symbols as a list.
+1import sympy as sym
+2
+3...
+4a, b, c, d, x, y = sym.symbols(’a, b, c, d, x, y’)
+5expr = a * x + b * x - c * x - a * y - b * y + c * y + d
+6# validate with sympy
+7fact = sym.collect(expr, d, func=sym.factor)
+8# model based factorization
+9func = Factorization(’Factorize d from the expression such that your final start with: ‘d + (...‘:’)
+10res = func(expr)
+11# compare res with fact
+12...
+2) Three tasks evaluate a models’ capability to resolve the logical operations AND, OR, and XOR.
+For instance, the test for logical AND combines the symbols Symbol("The horn only sounds on Sundays") and Symbol("I hear the horn") and compares the answer against the human-generated references ”The horn only sounds on Sundays and I hear the horn.” and ”Since I hear the horn it is Sunday.” Since there is a large number of possible solutions, there is high variability in the solution space.
+Each model might prefer a different solution.
+The following snippet shows how one can define a custom primitive class (CustomLogicPrimitive) for logic operators.
+The __or__ function gets overloaded and uses the built-in logic decorator from the core package to create a local function that evaluates two Symbol instances.
+1from symai import core
+2from symai.ops.primitives import Primitive
+3
+4class CustomLogicPrimitive(Primitive):
+5 def __or__(self, other: Any) -> Any:
+6 @core.logic(operator=’or’)
+7 def _func(_, a: str, b: str):
+8 pass # could impl. a fallback behavior here
+9 return self._to_symbol(_func(self, other))
+10 ...
+11
+12subject = ’cat’
+13res = (Symbol(f’The {subject} has whiskers.’, primitives=CustomLogicPrimitive) | \
+14 Symbol(f’The {subject} has a tail.’, primitives=CustomLogicPrimitive))
+3) For another task we use a custom Expression that defines a DSL syntax and semantic structure.
+We use this Expression to extract higher-order logic expressions from a natural language statement, namely the puzzle ’Who is Jay’s brother?’101010 Bob has two sons, John and Jay.
+Jay has one brother and father.
+The father has two sons.
+Jay’s brother has a brother and a father.
+Who is Jay’s brother?, that preserves the original relationships.
+The following is a DSL snippet of the ’Who is Jay’s brother?’ puzzle:
+1// Query
+2IsBrotherOf(jay, john, bob) <- BrotherOf(jay, john) AND FatherOf(bob, jay) AND FatherOf(bob, john);
+3
+4// Facts
+5BrotherOf(x, y) <- HAS(x, brother) AND HAS(y, brother) AND Sibling(x, y);
+6FatherOf(x, y) <- HAS(x, son) AND ParentOf(x, y);
+7ParentOf(x, y) <- IS(x, parent) AND IS(y, child);
+8Sibling(x, y) <- IS(x, father) AND IS(y, father) OR IS(x, mother) AND IS(y, mother);
+9
+10...
+4) For the final task, we again use the puzzle ’Who is Jay’s brother?’ to evaluate a models’ capability for complex conversions.
+We use the Z3 theorem prover (Moura & Bjørner, 2008) to solve the ’Who is Jay’s brother’ puzzle conditioned on the Z3 solvers’ solution to Einsteins’ famous puzzle ’Who owns the fish?’.
+The task involves an indirect translation from natural language to executable code by the Z3 solver; the solution to Einstein’s puzzle acts as a form of self-contained ”documentation” for how the Z3 solver should be utilized.
+The test constructs a template, which includes the task instructions, puzzle statement, and reference to the Einstein’s puzzle solution.
+The models are then asked to analyze the given problem and solution format and create a Python function with Z3 syntax that can solve the ’Who is Jay’s brother?’ puzzle.
+The dynamically generated code is executed within the test environment utilizing Python’s exec function.
+We check the access to the Z3 solver and run the generated solve_puzzle function supposed to contain the logic to solve the puzzle.
+Once executed, the assembled Z3 logical clauses are processed by the solver, which verifies that the set of constraints is satisfiable.
+If so, the model generated by the solver is queried for the puzzle’s solution and scored using our VERTEX score.
+The following is an example output from the Z3 representation of the solution to ’Who is Jay’s brother?’ puzzle:
+1from z3 import Solver, Bool, And, Not, Const, BoolSort, EnumSort, Function, IntSort
+2
+3def solve_puzzle(S: Solver) -> Const:
+4 # Define the enumeration sort for the individuals
+5 Person, (BobE, JohnE, JayE, JaysBrotherE, FatherE) = EnumSort(’Person’, [’Bob’, ’John’, ’Jay’, ’JaysBrother’, ’Father’])
+6
+7 # Define a function from boolean to persons (for brother status)
+8 is_brother = Function(’is_brother’, Person, BoolSort())
+9
+10 # Define the relationships
+11 S.add(is_brother(JohnE) == True) # John is a brother
+12 S.add(is_brother(JayE) == True) # Jay is a brother
+13
+14 ...
+15
+16 return query
+F.5 Hierarchical Computational Graphs
+In this section we extend on the hierarchical computational graphs section.
+Research Paper Draft
+The following example defines a Paper expression that takes in a sequence of expressions which are sequentially executed.
+The Method expression contains a Source expression which points to the actual human-written method.
+The Method expression acts as the root node that bootstraps the generation process.
+The RelatedWork expression contains a sequence of Cite expressions which are executed in parallel and utilized to define the context of the related work section.
+The Abstract and Title expressions get executed last because they require all the previous information to be available in their respective contexts.
+Each expression in the sequence of expressions from Paper takes in the context of its predecessors.
+All expressions also link their results to a global linker object which is utilized after the execution to retrieve individual results from the nodes of the expression’s computational graph.
+Each node was evaluated against its corresponding reference, all references representing actual sections from this research paper.
+The samples for each node were generated with a separate model (Claude 2) that was not part of this evaluation.
+In Figure 5 we show the resulting computational graph of the Paper expression.
+1# define the computational graph
+2expression = Paper(
+3 Method(
+4 # link to original code base where the main method is defined
+5 Source(file_link=’/path/to/.../file’),
+6 ),
+7 # gather resources and write the related work
+8 RelatedWork(
+9 Cite(file_link=’/path/to/.../file’),
+10 Cite(file_link=’/path/to/.../file’),
+11 ...
+12 ),
+13 # write the abstract and title
+14 Abstract(),
+15 Title(),
+16)
+17# run the graph
+18paper_result = expression(’Write a scientific paper’)
+19# access linker to retreive the results from the method expression
+20method = expr.linker.find(’Method’)
+21# print result of the method expression
+22print(method)
+F.6 Caveats
+One may have noticed that there are cases in which there is no need to sample multiple samples as there is only one expected answer, for instance, if we need to extract a specific number from a string and cast it to an integer.
+If the extraction process appends additional characters other than the number, the casting will fail.
+In such cases, the VERTEX score simply defaults to the chosen similarity measure, registering and penalizing any deviation from the expected answer.